### v1

# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

Abstract
We present Connection-Aware Motif Sequencing (CamS), a graph-to-sequence representation that enables decoder-only Transformers to learn molecular graphs via standard next-token prediction (NTP). For molecular property prediction, SMILES-based NTP scales well but lacks explicit topology, whereas graph-native masked modeling captures connectivity but risks disrupting the pivotal chemical details (e.g., activity cliffs). CamS bridges this gap by serializing molecular graphs into structure-rich causal sequences. CamS first mines data-driven connection-aware motifs. It then serializes motifs via scaffold-rooted breadth-first search (BFS) to establish a stable core-to-periphery order. Crucially, CamS enables hierarchical modeling by concatenating sequences from fine to coarse motif scales, allowing the model to condition global scaffolds on dense, uncorrupted local structural evidence. We instantiate CamS-LLaMA by pre-training a vanilla LLaMA backbone on CamS sequences. It achieves state-of-the-art performance on MoleculeNet and the activity-cliff benchmark MoleculeACE, outperforming both SMILES-based language models and strong graph baselines. Interpretability analysis confirms that our multi-scale causal serialization effectively drives attention toward cliff-determining differences.
Graph Representation Learning, Autoregressive Modeling, Self-Supervised Learning, Science Foundation Models, Molecular Property Prediction
1 Introduction
Molecular property prediction is a core challenge in drug discovery (wu2018moleculenet) and has increasingly become a primary benchmark for “Foundation Models (FM) for Science.” (khan2025comprehensive) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (wang2024emu3; xia2025nature; zhang2025unigenx). Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (touvron2023llama). However, a critical gap remains: current NTP-based molecular models, which largely rely on 1D SMILES strings, fail to explicitly capture graph topology and consequently trail behind specialist graph models in predictive accuracy (xia2025nature).
Conversely, graph-native methods explicitly model topology but typically depart from pure decoder-only NTP, instead adopting hybrid architectures or objectives that involve, to varying degrees, input corruption (e.g., masked node prediction) (li2023knowledge; shehzad2024graph; lu2025uni; kong2025unimomo). While sophisticated strategies—such as incorporating global knowledge nodes or avoiding random masking—have been developed to mitigate the information loss caused by masking (li2023knowledge; rong2020self; liu2024mask; you2020graph; xu2021self), these methods remain fundamentally bound by the corruption-coverage trade-off (wettig2023should). This limitation is particularly acute in activity-cliff scenarios, where a single atomic change triggers a drastic property shift; masking the local neighborhood of a critical functional group effectively removes the precise evidence needed to discern such subtle differences. Standard NTP, by contrast, offers a dense supervision signal without corrupting the visible prefix context (clark2020electra), yet it lacks a representation interface that exposes graph connectivity as effectively as text.
To bridge this gap, we propose Connection-Aware Motif Sequencing (CamS), a tokenizer-level interface that makes molecular graphs directly learnable by standard decoder-only NTP through a three-stage design. First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (shibata1999byte; geng2023novo; shen2024graphbpe). This produces motif vocabulary with frequent merge operation and introduces a tunable molecular scale that allows direct granularity control. To ensure faithful encoding of fine-grained chemical detail, we further integrate a Single-Atom Vocabulary Closure (SAVC) mechanism. Second, to make the graph autoregressive-ready, we partition the molecule into a BPEGraph at a specific scale and serialize it via scaffold-rooted breadth-first search (BFS)  (bundy1984breadth). This produces a CamS subsequence with a stable Intra-scale Order, moving from the global core to peripheral functional groups. Third, and most critically, to construct a hierarchical context, we concatenate subsequences from fine to coarse. This strategy circumvents the trade-off between scales and establishes an Inter-scale Order in the final CamS sequence, enabling high-level motifs to be predicted by conditioning on dense, uncorrupted fine-scale evidence. Collectively, these steps unify substructure compression and serialization through a dual-causal ordering—sequencing motifs within each scale and stacking scales hierarchically—to enable the model to comprehend global scaffolds based on explicit, uncorrupted local structural evidence.
Built on CamS, we instantiate CamS-LLaMA by pre-training a native decoder-only backbone with standard NTP on CamS sequences, keeping the architecture and objective unchanged (touvron2023llama). To keep attribution controlled, we inject only a lightweight molecular fingerprint (cereto2015molecular) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven. Across MoleculeNet (wu2018moleculenet) and the activity-cliff stress test MoleculeACE (van2022exposing), CamS-LLaMA achieves state-of-the-art performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (li2023knowledge). Beyond accuracy, we provide mechanism-level evidence. Targeted ablation studies confirm that multi-scale concatenation is the key driver of performance, while interpretability analysis reveals that the causal serialization effectively drives attention toward cliff-determining structural differences.
Contributions. (1) Methodological Framework: We introduce CamS, a unified graph-to-causal-sequence interface that resolves the conflict between graph topology preservation and scalable autoregressive training, enabling NTP to function as a structure-native objective. (2) Empirical Validation: We demonstrate that CamS-LLaMA achieves state-of-the-art performance on MoleculeNet and MoleculeACE benchmarks. Mechanism analysis reveals that the multi-scale causal context explicitly improves attention focus on subtle, activity-determining structural edits. (3) Implementation Recipe: We provide a reproducible pipeline—from the CamS-tokenizer to the CamS-LLaMA—establishing a robust baseline for applying standard foundation model architectures to molecular science.
2 Related Work
2.1 Molecular Property Prediction within Foundation Models Paradigm
Molecular property prediction is increasingly adopting a foundation model approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (awais2025foundation; xia2025nature). One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (xia2025nature; cai2025chemfm). This leverages mature LLM infrastructure and enables straightforward large-scale training. However, SMILES is not structure-native, weakening topology-based signals. String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (xia2025nature). Alternatively, graph-native foundation models directly process molecular graphs to preserve connectivity and local topology (you2021graph; liupre; you2020graph; xu2021self). High-performing methods typically employ encoder-only Graph Transformers (shehzad2024graph) designed for property prediction. While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (li2023knowledge; rong2020self). More recent structural foundation models integrate graphs with autoregressive components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (kong2025unimomo; lu2025uni; zhang2025unigenx).
2.2 Molecular Tokenization Strategies
Within a foundation model paradigm, tokenizer design becomes a key interface for scaling LLMs. In NLP, BPE yields strong compression via simple frequency-based merges without training an extra tokenizer model. Molecular tokenization spans string-based and graph-based methods (wang2019smiles; liu2023rethinking). String-based schemes operate on SMILES (weininger1988smiles) or related linearizations (krenn2022selfies), while graph-based schemes have evolved from basic node/edge-level units to fragment-level motifs (liu2023rethinking). Some rule-based motif tokenizations rely on limited handcrafted chemistry rules (zhang2021motif); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (li2023knowledge); ring-or-path-based (wollschlager2024expressivity) and tree-based (jin2018junction) schemes also expose substructures at different granularities; some other works learn tokenization via trainable rules (sun2024representing) or by training a GNN as a tokenizer (liu2023rethinking; luonode). Particularly noteworthy are data-driven motif mining approaches that adapt BPE idea to graphs to obtain reusable, multi-scale fragments with strong compression and minimal reliance on manually designed rules or additional trained models (geng2023novo; shen2024graphbpe; kong2022molecule). In addition, cross-level tokenization schemes, which combine motifs at different structural levels, also provide complementary design ideas (ji2022relmole; chenhierarchical).
3 Method
Our molecular representation must satisfy two requirements: (1) chemical fidelity, i.e., sensitivity to pharmacophore-level structures and their topology; and (2) autoregressive foundation model compatibility, i.e., a compressed interface naturally aligned with an NTP paradigm (touvron2023llama). Motivated by this, we propose Connection-Aware Motif Sequencing (CamS), which segments molecular graphs into reusable frequent motifs and, crucially, serializes the motif graph into multi-scale connection-aware causal sequences compatible with NTP pre-training (Section 3.1). Built on CamS, CamS-LLaMA performs large-scale autoregressive pre-training and injects a only lightweight fingerprint prior during downstream fine-tuning (Section 3.2). Finally, we analyze the characteristics and advantages of multi-scale CamS autoregressive modeling by comparing it with Graphformer-style modeling (Section 3.3).
3.1 CamS-Tokenizer
The CamS-tokenizer framework operates in three logical phases to bridge graph topology and scalable NTP. (1) CamS Vocabulary Construction: we maintain a robust vocabulary via BPE-style mining and Single-Atom Vocabulary Closure (SAVC) (Section 3.1.1). (2) Per-Scale Encoding: molecules are partitioned into motif graphs (BPEGraph) and serialized into single-scale CamS subsequences via scaffold-anchored BFS (Section 3.1.2). (3) Cross-Scale Concatenation: subsequences are organized in a fine-to-coarse order to form the final CamS sequence, enabling multiscale autoregressive modeling (Section 3.1.3). Full details are in Appendix A.
3.1.1 Substructure Level: CamS Vocabulary
BPE-style Mining. Following (geng2023novo; shen2024graphbpe), we first learn an ordered list of merge operations $\mathcal{O}$ from co-occurrence statistics. We then apply $\mathcal{O}$ to the corpus to mine the resulting subgraphs. This yields the BPEMotif set (frequent subgraphs with connectivity patterns) and the base BPE-SAV (Single-Atom Vocabulary with connectivity patterns) naturally captured during mining.
Single-Atom Vocabulary Closure (SAVC). While generation operates within a closed vocabulary, predictive encoding encounters unseen atom states. Standard GraphBPE tokenizers greedily match largest motifs and recursively back off; if a specific atom-connectivity form is absent, it collapses to [UNK]. This erases chemical details (e.g., element type) and creates Causal Information Breakpoints in the autoregressive stream. To prevent this, SAVC constructs the BPE-SAV Back-off via two strategies: (1) enumerating connection-aware tokens for all typical valences, and (2) mapping rare atypical forms of element X to <X_AltForm> rather than <UNK>.
The final CamS Vocabulary $\Sigma$ is the union of BPEMotif, the complete SAV (BPE-SAV $\cup$ Back-off), and special tokens mapped to unique integers. This guarantees that every atom is covered by either a high-level motif or a precise single-atom descriptor (details in Appendix A.1).
3.1.2 Per-Scale Encoding
Given a vocabulary scale $s$ (determined by the iteration count of merge operations), encoding a molecule $M$ involves three stages to generate a single-scale CamS subsequence.
BPEGraph Construction. We apply the prefix of merge operations to $M$, inducing a partition of atoms into non-overlapping fragments. This forms the BPEGraph $G_{s}=(V_{s},E_{s})$, where nodes $V_{s}\subset\Sigma$ are connection-aware motifs/atoms and $E_{s}$ represents chemical bonds between them.
Serialization (Intra-scale Order). To adapt the graph for decoder-only NTP, we serialize $V_{s}$ into an ordered list via Breadth-First Search (BFS). We select the motif with the largest atom count (typically the core scaffold) as the root. This strategy prioritizes global backbone structure before local substitutions (zhang2021motif), establishing a stable Center-to-Periphery causal order.
ID Extraction. Mapping the ordered nodes to their indices in $\Sigma$ yields the single-scale CamS subsequence $X^{(s)}=(x^{(s)}_{1},\dots,x^{(s)}_{L_{s}})$.
3.1.3 Cross-Scale Concatenation
Single-scale tokenization faces a trade-off: coarse scales capture global scaffolds but offer low resolution, while fine scales preserve atomic details but fragment high-level structures (e.g., pharmacophores). CamS resolves this by constructing a multi-scale causal context. We tokenize the molecule at multiple scales (e.g., $s\in\{1\mathrm{K},7\mathrm{K},27\mathrm{K},67\mathrm{K}\}$) by slicing $\mathcal{O}$. The resulting subsequences are concatenated in a fixed Fine-to-Coarse order (Inter-scale Order) to form the final CamS sequence $\mathbf{X}$:
$\mathbf{X}=[$ $\,\texttt{[BOS]},X^{(1\mathrm{K})},\texttt{[CONCAT]},X^{(7\mathrm{K})},\texttt{[CONCAT]},$ (1)
$\,X^{(27\mathrm{K})},\texttt{[CONCAT]},X^{(67\mathrm{K})},\texttt{[EOS]}].$
During training, this CamS token stream allows the model to leverage high-resolution local topology (fine scales) as a prefix to condition the prediction of global scaffolds (coarse scales). This effectively embeds bottom-up structural composition into the autoregressive objective without architectural modifications (analysis in Section 3.3).
3.2 CamS-LLaMA
Figure 1: The overall framework of CamS-LLaMA. (a) CamS-Tokenizer: Molecules are transformed into multi-scale causal sequences. Through Per-scale Encoding, motif graphs are serialized via Scaffold-rooted BFS (Intra-Scale Order) to establish CamS Subsequence, followed by Cross-Scale Concatenation to arrange subsequences into a final CamS Sequence from fine to coarse granularities (Inter-Scale Order). (b) Pre-training Pipeline: The model is pre-trained on CamS Token Lists using a standard LLaMA via NTP, learning to aggregate structural information. (c) Fine-tuning Pipeline: During downstream tasks, the pre-trained weights are transferred, and a lightweight fingerprint prior is injected (via projection, prepending in embedding layer and fusion in classifier or regressor) to take part in the prediction of specific properties.
CamS-LLaMA follows a representation-model co-design principle. Section 3.1 maps a molecular graph $G=(V,E)$ into a multi-scale, connection-aware causal sequence $X=(x_{1},\ldots,x_{T})$ (Eq. (1)) under vocabularies $\Sigma_{1\mathrm{K}}\subset\Sigma_{7\mathrm{K}}\subset\Sigma_{27\mathrm{K}}\subset\Sigma_{67\mathrm{K}}\triangleq\Sigma$, via BFS-ordered motif traversals $\mathcal{T}_{s}(G)$ and token mapping $x^{(s)}_{\ell}=\iota_{s}(\tau^{(s)}_{\ell})\in\Sigma_{s\mathrm{K}}\subseteq\Sigma$. The key point is that this graph-to-causal-sequence interface is intentionally designed to be transparent to a vanilla decoder-only Transformer: it exposes topology through a causal order and a fine-to-coarse hierarchy, so we can train with standard NTP without modifying the LLaMA backbone. Based on this interface, CamS-LLaMA follows the foundation model protocol: (1) self-supervised autoregressive pre-training on CamS sequences (figure 1 bottom right), and (2) supervised fine-tuning with a minimal structural prior (fingerprints) injected only at fine-tuning time (figure 1 bottom left).
3.2.1 Autoregressive Pre-training via NTP
CamS-LLaMA adopts a standard LLaMA-style decoder (touvron2023llama); in pre-training, every CamS token list is treated as an ordinary token sequence in the shared vocabulary $\Sigma$ and is consumed by the backbone under the usual causal mask, with no graph-specific architectural modification. Concretely, each molecule yields five CamS sequence views $\{X_{1},X_{7},X_{27},X_{67},X\}$: four single-scale BFS sequences tokenized with $\Sigma_{1\mathrm{K}}/\Sigma_{7\mathrm{K}}/\Sigma_{27\mathrm{K}}/\Sigma_{67\mathrm{K}}$, plus the concatenated multi-scale sequence $X$ in Eq. (1); we treat each view as one NTP training instance (details in Appendix A.3). Given any view $X=(x_{1},\ldots,x_{T})$, token embeddings form $\mathbf{H}^{(0)}$, which is transformed by $L$ causal self-attention blocks into $\mathbf{H}^{(L)}$; a tied output projection then yields $p_{\theta}(x_{t+1}\mid x_{\leq t})$.
Importantly, CamS changes what the model can condition on under the same NTP objective by reshaping the prefix context through fine-to-coarse concatenation (Eq. (1)). Let $\mathcal{I}_{\mathrm{tok}}\subseteq\{1,\ldots,T-1\}$ be valid prediction positions in $X$ (excluding [BOS], [CONCAT], [EOS]), and $x^{\star}_{t+1}$ be the ground-truth next token. The NTP loss is
$\mathcal{L}_{\mathrm{NTP}}(\theta;X)=-\sum_{t\in\mathcal{I}_{\mathrm{tok}}}\log p_{\theta}(x^{\star}_{t+1}\mid x_{\leq t}).$ (2)
Because $X$ concatenates BFS sequences from fine to coarse scales (Eq. (1)), the conditioning prefix $x_{\leq t}$ has an explicit hierarchy: for $t$ in $X_{67}$ (highest-merge segment), the model conditions on all finer-scale tokens $(X_{1},X_{7},X_{27})$ plus the coarse prefix. Thus, coarse-scale prediction errors propagate through attention to both coarse predecessors and the entire fine-scale prefix, providing a fine-to-coarse causal credit assignment under a fully standard decoder.
On the pre-training corpus $\mathcal{D}_{\mathrm{pre}}$, we minimize
$\min_{\theta}\;\mathbb{E}_{X\sim\mathcal{D}_{\mathrm{pre}}}\big[\mathcal{L}_{\mathrm{NTP}}(\theta;X)\big],$ (3)
learning general-purpose molecular representations purely from CamS sequences, without any downstream labels or extra descriptors.
3.2.2 Fine-tuning with Fingerprint Injection
Many molecular pre-training methods inject rich handcrafted descriptors early (e.g., during pre-training) (li2023knowledge). In contrast, we keep CamS as the primary representation and inject only one widely-used structural prior–a fingerprint–only during fine-tuning, to make the contribution of CamS representations more attributable and controlled.
We treat the fingerprint as a virtual global token. For each molecule, a fingerprint vector
$\mathbf{f}\in\mathbb{R}^{D_{\mathrm{fp}}},$ (4)
is projected to the model hidden size:
$\mathbf{f}^{\prime}=W_{\mathrm{fp}}\mathbf{f},\quad W_{\mathrm{fp}}\in\mathbb{R}^{H\times D_{\mathrm{fp}}},$ (5)
and prepended to token embeddings:
$\mathbf{H}^{(0)}_{\mathrm{ft}}=[\,\mathbf{f}^{\prime};\mathbf{e}_{1};\dots;\mathbf{e}_{T}\,]\in\mathbb{R}^{(T+1)\times H}.$ (6)
With a modified causal mask, all positions can attend to this global token. After $L$ layers:
$\mathbf{H}^{(L)}_{\mathrm{ft}}=[\,\mathbf{h}^{(L)}_{0};\mathbf{h}^{(L)}_{1};\dots;\mathbf{h}^{(L)}_{T}\,$]. (7)
We use the global token representation $\mathbf{u}^{(i)}=\mathbf{h}^{(L)}_{0}$ for prediction (a decoder-only analogue of [CLS]).
For classification:
$\begin{gathered}\hat{\mathbf{y}}^{(i)}=\mathrm{softmax}\!\big(W_{2}\,\sigma(W_{1}\mathbf{u}^{(i)})\big),\\ \mathcal{L}_{\mathrm{cls}}(\theta,\phi)=-\frac{1}{B}\sum_{i=1}^{B}\log\hat{\mathbf{y}}^{(i)}_{\,y^{(i)}}.\end{gathered}$ (8)
For regression:
$\begin{gathered}\hat{y}^{(i)}=W_{2}\,\sigma(W_{1}\mathbf{u}^{(i)}),\\ \mathcal{L}_{\mathrm{reg}}(\theta,\phi)=\frac{1}{B}\sum_{i=1}^{B}\big(\hat{y}^{(i)}-y^{(i)}\big)^{2}.\end{gathered}$ (9)
Fine-tuning minimizes
$\min_{\theta,\phi}\;\mathbb{E}_{(X,\mathbf{f},y)\sim\mathcal{D}_{\mathrm{task}}}\big[\mathcal{L}_{\mathrm{cls}/\mathrm{reg}}(\theta,\phi;X,\mathbf{f},y)\big].$ (10)
This keeps pre-training purely sequence/structure-driven, while using fingerprints as a lightweight and controlled prior only when optimizing for downstream targets.
3.3 CamS Representation vs. Graph Transformer
We analyze CamS-LLaMA and graph transformer-style models (e.g., KPGT) within a unified framework of Token-Level Graph-Structured Deep Feature Construction. We show that CamS yields (1) a denser and multi-view direct supervision signal and (2) a hierarchical inductive bias that are mathematically distinct from standard masked-node prediction (MNP). Detailed derivations and supporting discussions are provided in Appendix B.
3.3.1 Unified Framework.
Let $G=(V,E)$ be a molecular graph. Both paradigms construct a token graph $\mathcal{G}_{\mathrm{tok}}=(\mathcal{V}_{\mathrm{tok}},\mathcal{E}_{\mathrm{tok}})$ and perform $L$ layers of propagation:
$\begin{split}\mathbf{H}^{(l+1)}&=\mathrm{Agg}\big(\mathbf{A}^{(l)}\,\mathrm{Val}(\mathbf{H}^{(l)})\big),\\ \mathbf{A}^{(l)}_{ij}&\propto\exp\left(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d}}+\psi_{ij}\right).\end{split}$ (11)
where $\mathbf{H}^{(l)}$ are token embeddings, $\mathbf{q}_{i},\mathbf{k}_{j}$ are projections of $\mathbf{H}^{(l)}$, and $\psi_{ij}$ encodes the structural bias or causal mask.
Graphformer (Flat/Static Bias). $\mathcal{V}_{\text{tok}}$ corresponds to single-scale atoms or triplets. $\psi_{ij}$ encodes static structural biases (e.g., shortest-path-distance bias injected into attention) independent of token content, yielding an isotropic receptive field at a given graph distance (ying2021transformers).
CamS-LLaMA (Hierarchical/Causal Flow). $\mathcal{V}_{\text{tok}}$ comprises multi-scale motifs $\{v^{(s)}\}$, where the same molecule is re-represented across scales. $\psi_{ij}$ is determined by the causal mask $\mathbf{M}_{\text{causal}}$, enforcing a directed flow where coarse-scale tokens attend to fine-scale predecessors. This turns cross-scale aggregation into a learned, anisotropic connectivity along the sequence.
3.3.2 Information-Theoretic Context Analysis.
We contrast NTP with MNP. Let $x_{t}$ be a target token and let $Z_{t}$ denote its uncorrupted evidence set (e.g., the causal history for CamS coarse-scale tokens). MNP observes a stochastically corrupted context $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ via random masking.
Proposition 3.1 (Context Information Inequality).
For any non-trivial masking channel $\mathcal{M}$, we have
$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}).$ (12)
Sketch. The inequality follows from the Data Processing Inequality (DPI): masking is a stochastic post-processing of $Z_{t}$ that cannot increase mutual information (cover1999elements). In graph MNP, this loss of information is exacerbated by evidence-pattern uncertainty: the most relevant local-neighborhood evidence for a masked token may be co-masked, making the conditional distribution unstable across masking patterns. Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets / local subgraphs) or designing structure-guided masking, and by injecting global knowledge tokens (li2023knowledge; rong2020self; liu2024mask). In contrast, CamS NTP for coarse-scale tokens conditions on an uncorrupted fine-to-coarse history, avoiding such stochastic evidence loss. This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (yang2019xlnet).
3.3.3 Gradient Flow Density.
We define the Direct Supervision Density (SD) as the expected fraction of tokens serving as regression targets per update. For a multi-scale sequence with total length $T_{\text{all}}=\sum_{s=1}^{S}T_{s}$ and mask ratio $\rho$:
$\text{SD}_{\text{NTP}}\approx 1\quad\gg\quad\text{SD}_{\text{MNP}}=\rho.$ (13)
This trade-off (more predicted tokens vs. more corruption) is central to masked modeling (wettig2023should). Empirically, KPGT finds the best masking ratio can be as high as $\rho=0.5$ when using triplet-style tokens and global knowledge nodes to stabilize context (li2023knowledge). Meanwhile, dense token-level learning signals are known to substantially improve sample efficiency compared to sparse masked-only supervision (clark2020electra). CamS further strengthens this effect by providing multi-view supervision: the same molecule contributes token-level targets repeatedly across scales.
3.3.4 Hierarchical Inductive Bias.
Graphformers rely on implicit depth and static structural biases to propagate information. In contrast, CamS injects an explicit structural hierarchy via the multi-scale sequence $X=[X_{1},\dots,X_{T_{\text{all}}}]$, and the causal mask ensures coarse-scale predictions have deterministic access to fine-scale details. We hypothesize this is particularly beneficial for activity-cliff regimes, where small local motif edits trigger global property shifts, as it biases the model toward bottom-up composition rules.
4 Experiment
4.1 Experimental Setup
We decouple the pipeline into three stages, using the ChEMBL-34 Dataset for tokenizer training, the Enamine675M Dataset for NTP pre-training, and the MoleculeNet/MoleculeACE Benchmark for downstream fine-tuning and evaluation.
CamS-Tokenizer Training. CamS-tokenizer training is expensive because replaying $\mathcal{O}$ to materialize $\mathcal{V}$ repeatedly invokes RDKit’s subgraph-isomorphism matching (the dominant bottleneck) (landrum2021rdkit; ehrlich2011maximum) to extract motifs and attachment patterns, which is difficult to accelerate. We therefore train the tokenizer once on ChEMBL-34 (gaulton2017chembl), a moderately sized yet chemically diverse corpus, at the largest scale (67K). Concretely, we mine an ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K})$ with K=685K{=}685 merges, apply it to obtain $\mathcal{V}$, and then perform SAVC to ensure single-atom coverage for common elements. Smaller scales are obtained without retraining by prefix slicing: each scale uses a prefix $\mathcal{O}_{\leq k}$ together with its induced vocabulary subset (i.e., motifs mined by applying $\mathcal{O}_{\leq k}$), where 1K uses k=0k{=}0, 7K/27K use k=62/210k{=}62/210, and 67K uses k=685k{=}685. Implementation details and full procedures are provided in Appendix A.1.
Pre-training Dataset and Configuration. We instantiate a 16-layer LLaMA-style decoder (hidden size 720, 8 attention heads) using HuggingFace Transformers. For self-supervised pre-training, we use Enamine675M (675M molecules) (shivanyuk2007enamine) and optimize NTP (Section 3.2.1). For each molecule, we construct five CamS training sequences: four single-scale BFS sequences tokenized with 1K/7K/27K/67K vocabularies, plus the concatenated multi-scale sequence in Eq. (1). We treat each sequence as one training instance and uniformly shuffle all instances, resulting in roughly $5\times$ as many sequences as molecules. All optimization details (schedule, batch size, steps) are summarized in Appendix C.1.
Downstream Benchmarks and Baselines. We evaluate on MoleculeNet (11 tasks; broad endpoints that probe general-purpose molecular property prediction) (wu2018moleculenet) and MoleculeACE (30 activity-cliff tasks; stress-testing fine-grained structural sensitivity under subtle edits that induce large activity shifts) (van2022exposing). To enable an apples-to-apples comparison against a recent strong pre-training baseline, we adopt the data splits and evaluation protocol used in KPGT (li2023knowledge). On MoleculeNet, following KPGT’s setup, we compare primarily against self-supervised pre-training baselines to isolate the effect of (1) representation/tokenization and (2) the pre-training objective. On MoleculeACE, we report the benchmark’s comprehensive ML/DL baselines, since activity cliffs emphasize discriminating small structural changes where strong feature-based methods remain competitive. Unless stated otherwise, CamS-LLaMA is kept at $\sim$100M parameters (comparable to KPGT); fingerprints are injected only during fine-tuning as a minimal prior, while pre-training is purely sequence-based.
4.2 Main Results
We summarize the main comparisons in Table 1 (MoleculeNet) and Table 2 (MoleculeACE).
General property prediction (MoleculeNet). Table 1 shows that CamS-LLaMA achieves overall state-of-the-art on MoleculeNet under a comparable model scale to KPGT ($\sim$100M): it slightly improves the average AUROC in classification (0.845 vs. 0.843) and the average RMSE in regression (1.172 vs. 1.175). We also win 6/11 individual tasks (5/8 classification and 1/3 regression), and improve over the previous best self-supervised learning baselines by a clear margin (AVG-AUROC: 0.845 vs. 0.825; AVG-RMSE: 1.172 vs. 1.272). Notably, this is achieved with weaker prior injection: unlike KPGT, which explicitly feeds rich handcrafted descriptors, we keep pre-training purely sequence-based and inject only a lightweight fingerprint prior during downstream fine-tuning. Full per-baseline results and extended tables are provided in Appendix D.
[…]
Unified Takeaway. Taken together, Tables 12 indicate that CamS’s three-stage design (motif mining with single-atom completion $\rightarrow$ causal serialization $\rightarrow$ fine-to-coarse multi-scale concatenation) yields two complementary benefits: (1) general-purpose representation strength, by organizing functional-group semantics, attachment patterns, and topology into an NTP-friendly causal context; and (2) activity-cliff-oriented discriminability, by exposing cliff-driving edits through motif-level abstraction, position-sensitive causal contexts, and redundant multi-scale views. This motivates a targeted interpretability analysis below, where we test whether the model indeed allocates more attention to cliff-differential tokens than to shared scaffolds, and how this preference varies across scale regions.
[…]
Task Previous Best Baselines KPGT CamS-LLaMA
Method Score (Ours)
Classification Tasks (AUROC $\uparrow$)
bace GEM 0.857(0.016) 0.855(0.014) 0.870(0.013)
bbbp GEM 0.895(0.024) 0.908(0.012) 0.942(0.015)
clintox GEM 0.905(0.027) 0.946(0.026) 0.935(0.017)
estrogen GEM 0.894(0.048) 0.906(0.034) 0.917(0.050)
metstab GROV 0.876(0.046) 0.889(0.057) 0.891(0.059)
sider JOAO 0.640(0.012) 0.649(0.011) 0.655(0.016)
toxcast GEM 0.733(0.020) 0.745(0.003) 0.724(0.008)
[…]
AVG GEM 0.825 0.843 0.845
Regression Tasks (RMSE $\downarrow$)
esol GEM 0.803(0.051) 0.804(0.102) 0.761(0.046)
freesolv MolF 2.322(0.613) 2.121(1.025) 2.110(0.959)
lipo GROV 0.625(0.007) 0.600(0.012) 0.645(0.023)
AVG MolF 1.272 1.175 1.172
Table 2: Performance on MoleculeACE (RMSE). AVG: Average RMSE across all 30 tasks. Ranking: Bold (1st), Underline (2nd), Italic (3rd). Metric: $\downarrow$ denotes lower is better. Column: Previous Best Baselines denotes the best-performing method excluding KPGT; Type: ML (Machine Learning), DL (Deep Learning). Reference & Abbreviations: GMAE: GraphMAE (hou2022graphmae); MBRT: MoleBERT (xia2023mole); GROV: GROVER (rong2020self); SVM/RF/GBM: Support Vector Machine / Random Forest / Gradient Boosting Machine (van2022exposing). Descriptors: E: ECFP (rogers2010extended); M: MACCS (durant2002reoptimization).
Task Previous Best Baselines KPGT CamS-LLaMA
Method Type Score (Ours)
Regression Tasks (RMSE $\downarrow$)
[…]
4.3 Interpretability: Attention on Activity Cliffs
Interpretability setup and metric. To explain the activity-cliff advantage in Table 2, we test whether CamS-LLaMA preferentially attends to cliff-driving structural differences. Following the MoleculeACE protocol, we construct activity-cliff pairs on the test sets and identify differential vs. shared atom-level activity-cliff point; we then map them to tokens within each scale region of the concatenated CamS sequence (Appendix E). Given the final-layer attention distribution, we compute mean attention on differential tokens (MDTA) and shared tokens (MSTA) within a region and report the relative preference $\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100$, evaluated for each scale region and for the full concatenated sequence. We report the full attention metrics (MDTA/MSTA/Rel-DTAP) in Appendix D.
Findings and implications. Table 3 shows three patterns. (1) Overall preference is positive: on the concatenated sequence, Rel-DTAP is consistently >0>0 (14.79% w/o fingerprint; 14.23% w/ fingerprint), indicating that the model, on average, prioritizes cliff-driving differences over shared scaffolds. (2) Preference is scale-dependent: coarse regions (27K/67K) exhibit much stronger preference (24.33%/25.87% w/o fingerprint) than the fewest-merge region (1K, 8.08%), suggesting that coarse motifs provide a more salient interface for highlighting differential substructures, while fine scales contribute weaker but complementary local evidence. (3) Fingerprint acts as a stabilizer rather than a uniform booster: although fingerprint slightly reduces the overall preference, it corrects the anomalous 7K behavior (from -0.18% to 2.10%), consistent with fingerprint providing global guidance that resolves ambiguity at mid-scale tokenizations. Together, these results support the cliff-specific advantage of CamS: multi-scale causal serialization does not merely improve accuracy, but also steers attention toward structurally meaningful differences that drive activity cliffs.
Table 3: Relative Differential-Token Attention Preference (Rel-DTAP). Within each scale region of the concatenated multi-scale sequence, we compute per-molecule $\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100\%$, then average over all molecules from activity-cliff pairs (anchor and partner counted separately). Positive values mean higher per-token attention on differential tokens than on shared tokens.
Scale region Without FP With FP
1K 8.08% 4.72%
7K -0.18% 2.10%
27K 24.33% 23.60%
67K 25.87% 24.00%
CONCAT (all scales) 14.79% 14.23%
4.4 Ablation Study
Ablation setup. To isolate the contribution of each core component, we evaluate three variants against the full model (Table 4): (1) w/o FP removes fingerprint injection during fine-tuning; (2) Vocab-1K Only and (3) Vocab-67K Only remove multi-scale concatenation by restricting both pre-training and fine-tuning to a single scale. For fair comparison, single-scale variants keep the pre-training budget matched to the full model (details in Appendix C.1).
Multi-scale concatenation is indispensable. The full model consistently outperforms both single-scale variants across all benchmarks (e.g., MolNet-Cls 0.845 vs. 0.833/0.818; MolNet-Reg 1.172 vs. 1.215/1.329; MolACE 0.624 vs. 0.641/0.649), confirming that multi-scale concatenation is not a minor engineering choice but the key mechanism that lets coarse-scale reasoning build on fine-scale evidence within one causal context. This directly supports our central claim: multi-scale causal serialization is the representation-level ingredient that makes graph structure learnable by a vanilla decoder-only NTP backbone.
Why 1K-only can beat 67K-only under a fixed NTP budget. Under single-scale training, Vocab-67K Only is consistently weaker than Vocab-1K Only, especially on regression (1.329 vs. 1.215 RMSE on MolNet-Reg). We interpret this as an over-compression effect: at 67K, sequences for small molecules become very short and each token packs high compositional complexity, which reduces the number of NTP supervision steps per molecule and makes causal dependency learning harder under limited model capacity and training steps. In contrast, 1K produces longer sequences that provide denser autoregressive supervision and more stable causal credit assignment, albeit at the cost of fragmenting scaffold-level semantics. Concat resolves this trade-off by retaining the dense supervision of fine scales while reintroducing scaffold-level abstraction at coarse scales, without committing to a single “best” vocabulary size.
[…]
Table 4: Ablation study and ranking comparison. Scores represent the average performance across all tasks in the respective benchmarks from Tables 1 and 2. Rank: Global ranking of the specific variant when pooled with the complete set of baseline models. Previous Best: The highest-performing baseline excluding KPGT. Abbreviations: MolF: MolFormer
Method MolNet-Cls ($\uparrow$) MolNet-Reg ($\downarrow$) MolACE ($\downarrow$)
Score Rank RMSE Rank RMSE Rank
CamS-LLaMA (Full) 0.845 1 1.172 1 0.624 1
– w/o FP 0.838 3 1.195 3 0.650 5
– Vocab-1K Only 0.833 4 1.215 4 0.641 3
– Vocab-67K Only 0.818 6 1.329 7 0.649 4
KPGT 0.843 2 1.175 2 0.633 2
[…]
5 Conclusion
We introduced CamS, a graph-to-sequence interface that enables standard decoder-only Transformers to learn molecular topology via Next Token Prediction. By unifying motif mining, causal serialization, and fine-to-coarse concatenation, CamS constructs a hierarchical context where global scaffolds are conditioned on dense local evidence. The resulting CamS-LLaMA achieves state-of-the-art performance on MoleculeNet and MoleculeACE, surpassing strong graph baselines while relying on minimal priors. Mechanism analysis confirms that this multi-scale causal design is crucial, as it explicitly drives attention toward cliff-determining structural differences. Our work validates that with the right representation interface, generic autoregressive foundation models can serve as powerful, structure-native engines for molecular science.
Appendix A CamS-tokenizer and Graph-to-Sequence Construction
Overview. This section corresponds to the CamS-tokenizer description in Sec. 3.1 (including Secs. 3.1.1, 3.1.2, and 3.1.3) and the multi-view training-instance construction in Sec. 3.2.1. It supplements the main text with implementation-level details of (1) GraphBPE merge learning and Single-Atom Vocabulary Closure (SAVC)(Algs. 1 and Algs. 2), (2) connection-aware motif vocabulary construction and encoding-time unknown recovery (Algs.3 and Algs.4), (3) deterministic scaffold-rooted BFS serialization (Algs. 5), and (4) fine-to-coarse multi-scale concatenation together with NTP loss masking and length handling (Algs. 6)
A.1 Tokenizer Vocabulary Mining and Single-Atom Coverage (GraphBPE + SAVC)
Notation and objects.
We start from an RDKit molecule and construct the atom graph $G_{0}=(V_{0},E_{0})$, where each node is an atom and each edge is a chemical bond. A GraphBPE merge list is an ordered sequence of operations $\mathcal{O}=(o_{1},\dots,o_{K})$, where each operation $o_{t}$ is a canonical fragment code extracted from the union of two adjacent nodes (details below). Given a prefix $\mathcal{O}_{\leq k}$, applying merges on $G_{0}$ yields a motif graph $G=(V,E)$, where each motif node $v\in V$ stores the covered atom set $\mathrm{atom\_indices}(v)\subseteq V_{0}$.
Connection-aware motif representation.
Each motif token is stored in a pair form $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$. The first records the fragment SMILES of the core atoms (no connectors), while the second inserts wildcard connector atoms * at all attachment sites together with explicit bond types, making the token identity connection-aware. We materialize these motif pairs by replaying $\mathcal{O}_{\leq k}$ on a tokenizer corpus and inserting connector * atoms for every cut bond (Alg. 3).
Single-atom vocabulary closure (SAVC).
A key issue for property prediction is encoding coverage: if a rare but meaningful “single-atom + attachment-pattern” form is absent from the mined vocabulary, it would collapse to [UNK] and erase fine-grained distinctions. SAVC addresses this by (1) enumerating common connection-aware single-atom motifs for typical valence states, and (2) adding an element-wise fallback token X_AltForm to represent rare/atypical forms. At encoding time, if a queried motif contains exactly one core atom of element $X$ but the exact form is missing, we back off to X_AltForm instead of [UNK] (Alg. 2).
Encoding-time unknown recovery.
Encoding greedily applies $\mathcal{O}_{\leq k}$ and then maps each resulting motif to its token id. If a motif is still unknown after canonicalization, we recursively split it along its stored merge tree until known motifs are obtained; the recursion stops at single-atom leaves where SAVC backoff applies (Alg. 4).
Algorithm-to-component map.
Alg. 1 learns the ordered merge list $\mathcal{O}$ (Sec. 3.1.1). Alg. 3 materializes the connection-aware motif vocabulary by replaying $\mathcal{O}_{\leq k}$ and inserting connector atoms *. Alg. 2 specifies SAVC enumeration and X_AltForm backoff. Alg. 4 describes encoding-time motif extraction and recursive split for unknown recovery.
A.2 Graph-to-Causal-Sequence Serialization (Traversal and Root Selection)
Motif graph.
After applying $\mathcal{O}_{\leq k}$ to $G_{0}$, we obtain a motif graph $G=(V,E)$ where each node $v\in V$ corresponds to a merged fragment with $\mathrm{atom\_indices}(v)\subseteq V_{0}$. An edge $(u,v)\in E$ indicates that at least one original bond in $G_{0}$ connects atoms across the two fragments.
Scaffold-rooted BFS order.
CamS serializes $G$ into a token sequence by a deterministic scaffold-rooted BFS: we select as root the motif with the largest number of covered atoms (ties broken deterministically, e.g., by smaller node id), and traverse the motif graph breadth-first with a fixed neighbor ordering (e.g., ascending node id). The visitation order defines the final token order, implemented by relabeling motif nodes so that ids $0,1,\dots$ encode the BFS order (Alg. 5). This yields a stable core-to-periphery causal order aligned with prefix-based NTP (Sec. 3.1.2).
Why no explicit edges in the sequence.
We do not append edge/direction tokens: each motif token already includes explicit * attachment sites and bond types, so the attachment pattern is part of the token identity (Alg. 3). The BFS order then provides the autoregressive causal structure at the sequence level.
A.3 Multi-scale Concat Construction and Training-Instance Generation
Prefix slicing and scale definition.
GraphBPE yields a single merge list $\mathcal{O}=(o_{1},\dots,o_{K})$. Each vocabulary scale corresponds to a prefix $\mathcal{O}_{\leq k_{s}}$, which induces a sub-vocabulary $\Sigma_{\leq k_{s}}$ and a tokenizer $\iota_{s}$. This “train once, slice many” construction reuses the same learned merge statistics across all scales (Sec. 3.1.3).
Per-molecule views and fine-to-coarse concatenation.
For a molecule $s$, we generate $S$ single-scale views $\{X_{s_{1}},\dots,X_{s_{S}}\}$ by encoding under tokenizers ordered from fine to coarse. We then build one concatenated view by inserting [CONCAT] as separators and keeping only one global [BOS] and [EOS]. Alg. 6 provides the pseudo-code used in implementation (Sec. 3.2.1).
Loss masking and length handling.
For NTP training, we exclude [BOS], [EOS], and [CONCAT] positions from prediction targets (typically by setting labels to -100 at these indices). Sequences are padded with [PAD] up to the maximum length; if truncation is needed, we keep boundary tokens [BOS] and [EOS] and truncate the interior span consistently across views.
Algorithm 1 GraphBPE merging-operation learning (ordered merge list $\mathcal{O}$)
1:Input: tokenizer corpus $\mathcal{D}$ (RDKit molecules), max iterations $K$, min frequency $f_{\min}$
2:Output: ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K^{\prime}})$
3: Initialize each molecule $m\in\mathcal{D}$ with a working graph $G_{m}$ where every atom is a node and $\mathrm{atom\_indices}(i)=\{i\}$
4: Compute initial statistics: for each $m$ and each edge $(u,v)\in E(G_{m})$, set $c\leftarrow\mathrm{FragSmiles}(m,\mathrm{atom\_indices}(u)\cup\mathrm{atom\_indices}(v))$ and increment $\mathrm{stats}[c]$
5:for $t=1$ to $K$ do
6:  Select $o_{t}\leftarrow\arg\max_{c}\;(\mathrm{stats}[c],\,c)$ (tie-break by lexicographic $c$ as in code)
7:  if $\mathrm{stats}[o_{t}]<f_{\min}$ then
8:   break
9:  end if
10:  Append $o_{t}$ to $\mathcal{O}$
11:  Apply $o_{t}$ to all molecules that contain it: in each affected $G_{m}$, merge every edge $(u,v)$ whose union-fragment code equals $o_{t}$
12:  Update local pair statistics around merged nodes (remove old incident pairs; add new pairs)
13:  Set $\mathrm{stats}[o_{t}]\leftarrow 0$ (avoid re-selecting the same operation)
14:end for
Algorithm 2 Single-Atom Vocabulary Closure (SAVC) with AltForm backoff (implementation-level)
[…]
2:Output: basic single-atom token list $\Sigma_{\mathrm{atom}}$
3:for each element $X\in\mathcal{E}$ do
4:  Add X (standalone single atom) and X_AltForm (rare/atypical fallback)
5:  for each $v\in\mathrm{Val}(X)$ do
6:   for each attachment degree $d\in\{1,\dots,v\}$ do
7:    for each bond-type tuple $(b_{1},\dots,b_{d})\in\mathcal{B}^{d}$ do
8:     Construct a single-atom fragment SMILES with $d$ wildcard substituents (use branching if $d>1$), e.g. *X(*)(*)* with explicit bonds
9:     Add this motif_with_conn into $\Sigma_{\mathrm{atom}}$
10:    end for
11:   end for
12:  end for
13:end for
14:Encoding-time backoff (MotifVocab.__getitem__): if a queried motif SMILES is not found after RDKit canonicalization, and it contains exactly one core atom (non-H, non-*), return the id of X_AltForm; otherwise return [UNK] (code uses id =1=-1 as unknown sentinel).
Algorithm 3 Materializing the connection-aware motif vocabulary (motif mining + SAVC integration)
1:Input: tokenizer corpus $\mathcal{D}$ (SMILES), operation prefix $\mathcal{O}_{\leq k}$, special tokens $\{\texttt{[PAD]},\texttt{[BOS]},\texttt{[EOS]},\texttt{[UNK]}\}$
2:Output: motif-pair list $\mathcal{V}=\{(v_{\mathrm{noConn}},v_{\mathrm{withConn}})\}$ and token set $\Sigma$
3:for each SMILES $s\in\mathcal{D}$ do
4:  Parse $s$ to molecule $m$; build $G_{0}=(V_{0},E_{0})$ with atom/bond features
5:  Apply merges in order $\mathcal{O}_{\leq k}$ on a copy of $G_{0}$ to obtain motif graph $G=(V,E)$ where each node $v$ stores $\mathrm{atom\_indices}(v)\subseteq V_{0}$
6:  Assign each atom $i\in V_{0}$ to its motif node $\mathrm{bpe\_node}(i)\in V$
7:  for each original bond $(i,j)\in E_{0}$ do
8:   if $\mathrm{bpe\_node}(i)\neq\mathrm{bpe\_node}(j)$ then
9:    Add two connector nodes $i^{\star},j^{\star}$ with smarts=*; attach $i^{\star}$ to $i$ and $j^{\star}$ to $j$
10:    Set the bond types of $(i,i^{\star})$ and $(j,j^{\star})$ equal to the original bond type of $(i,j)$
11:    Include $i^{\star}$ in $\mathrm{atom\_indices}(\mathrm{bpe\_node}(i))$ and $j^{\star}$ in $\mathrm{atom\_indices}(\mathrm{bpe\_node}(j))$
12:   end if
13:  end for
14:  for each motif node $v\in V$ do
15:   $v_{\mathrm{noConn}}\leftarrow\mathrm{FragSmiles}(m,\mathrm{atom\_indices}(v)\cap V_{0})$
16:   $v_{\mathrm{withConn}}\leftarrow\mathrm{GraphSmiles}(G_{0}[\mathrm{atom\_indices}(v)])$ (contains * connectors)
17:   Add pair $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$ into a frequency counter
18:  end for
19:end for
20: Prepend special tokens; then integrate the SAVC basic vocabulary (Alg. 2) into the pair list
Algorithm 4 GraphBPE encoding with unknown-motif recovery (recursive split)
1:Input: SMILES $s$, operations $\mathcal{O}_{\leq k}$, motif vocabulary $\Sigma$ (lookup by canonical SMILES)
2:Output: ordered motif list $\big[(v_{\mathrm{noConn}},v_{\mathrm{withConn}})\big]$
3:if $s$ contains “.then
4:  Split $s$ into fragments and encode each fragment independently; concatenate motif lists
5:end if
6: Build initial atom graph $G_{0}$ and initialize a binary merge tree at each atom node
7: Apply $\mathcal{O}_{\leq k}$ in order: for each code $o$, merge every edge whose union-fragment code equals $o$ (merge also updates the node’s merge tree)
8: Relabel nodes to integers and reorder them by scaffold-rooted BFS (Alg. 5)
9: Add connector * nodes for every cut bond to build connection-aware subgraphs (Alg. 3)
10: Collect motif tokens in BFS order; if a motif is not in $\Sigma$, mark it as unknown
11:while unknown motifs exist do
12:  Pop one unknown motif node and recursively split it into its two children in the stored merge tree
13:  For each child: if it is in $\Sigma$, emit it; else split further; stop at leaves (single-atom motifs), where SAVC backoff applies
14:end while
Algorithm 5 Scaffold-rooted BFS serialization (root selection + neighbor order)
1:Input: motif graph $G=(V,E)$ with node attribute $\mathrm{atom\_indices}(v)$
2:Output: a deterministic BFS ordering $\pi$ (used to relabel nodes)
3: Select root $r\leftarrow\arg\max_{v\in V}\big(|\mathrm{atom\_indices}(v)|,\;-\mathrm{id}(v)\big)$ (tie-break: smaller id)
4: Initialize queue with $r$, mark $r$ visited
5:while queue not empty do
6:  Pop front node $u$, append $u$ to ordering $\pi$
7:  Let $\mathcal{N}(u)$ be neighbors of $u$ not yet visited, sorted by ascending node id
8:  Push neighbors in $\mathcal{N}(u)$ into the queue and mark them visited
9:end while
10: Relabel nodes by $\pi$ so that the first visited node becomes id 0, the second becomes id 11, etc.
Algorithm 6 Multi-scale CamS concatenation and training-view generation
[…]
2:Output: single-scale views $\{X_{s_{j}}\}_{j=1}^{S}$ and concatenated view $X_{\mathrm{concat}}$
3:for $j=1$ to $S$ do
4:  $X_{s_{j}}\leftarrow\iota_{s_{j}}.\mathrm{encode}(s).\mathrm{ids}$ (already includes [BOS] and [EOS])
5:end for
6: Build concatenated ids:
7:  $X_{\mathrm{concat}}\leftarrow[\texttt{bos\_id}]$
8:for $j=1$ to $S$ do
9:   Append the inner ids of $X_{s_{j}}$ (drop its BOS/EOS): $X_{s_{j}}[2{:}{-}1]$
10:  if $j<S$ then
11:    Append concat_id
12:  end if
13:end for
14:  Append eos_id
15:NTP loss masking: exclude [BOS], [EOS], [CONCAT] positions from prediction targets (implementation typically sets labels to -100 on these ids).
16:Padding/truncation: pad with [PAD] to max length; if truncating, keep the boundary tokens [BOS] and [EOS].
Appendix B Theoretical Derivations
Overview. This section corresponds to the theoretical analysis in Sec. 3.3 and the supporting arguments referenced in Secs. 3.1 and 3.2.1. It supplements the main text with derivation-level details of (1) the DPI-based proof for the information loss induced by stochastic corruption (Proposition 3.1; Sec. B.1), (2) the direct supervision density / target-coverage counting under NTP vs. masked modeling in the multi-scale setting (Sec. B.2), and (3) the explicit structural-bias formulation that contrasts graph-transformer static bias with CamS causal constraint plus learned aggregation (Eq. (11); Sec. B.3).
B.1 Proof of Proposition 3.1 (Context Information)
We formalize the conditioning-information gap between (1) uncorrupted evidence and (2) randomly masked evidence.
Setup.
For predicting token $x_{t}$, let $Z_{t}$ denote an unmasked evidence set (a collection of tokens) and let $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ be its stochastically masked version under a masking channel $\mathcal{M}$ (e.g., random node/substructure masking used in MLM/MNP). This yields a Markov chain
$x_{t}\;\longleftrightarrow\;Z_{t}\;\longrightarrow\;\tilde{Z}_{t}.$ (14)
Proof.
[…]
$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}),$ (15)
which proves Proposition 3.1. ∎
Remark (MLM limitations in NLP).
XLNet points out two intrinsic issues of corruption-based MLM: (1) the special token corruption (e.g., [MASK]) creates a pretrain–finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context (yang2019xlnet). Relatedly, ELECTRA shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions, while the [MASK] mismatch contributes a smaller but non-negligible part (clark2020electra).
Why this matters more on graphs (evidence-pattern uncertainty).
On graphs, the most predictive evidence for a token often lies in its local neighborhood (e.g., bond context, ring connectivity, functional-group surroundings). Random masking can therefore remove not only the target token but also its critical neighborhood evidence, making the missingness pattern itself a stochastic nuisance. This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability: (1) masking structured units and adding global knowledge nodes (KPGT) (li2023knowledge), (2) masking local subgraphs and predicting contextual properties (GROVER) (rong2020self), and (iii) structure-guided / curriculum masking rather than purely random masking (StructMAE) (liu2024mask). These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (wettig2023should).
CamS implication.
For CamS, coarse-scale token prediction under NTP conditions on an uncorrupted fine-to-coarse history by construction. Thus, relative to MNP which observes $\tilde{Z}_{t}$, CamS provides a higher-SNR conditioning set $Z_{t}$ that avoids stochastic evidence deletion. This does not claim that an arbitrary prefix always dominates the full bidirectional context; rather, it claims that masking as corruption strictly reduces the information of whatever evidence set is available, and CamS is designed so that the evidence set for coarse tokens is both uncorrupted and structurally aligned with the fine-grained neighborhood evidence.
B.2 Direct Supervision Density Analysis (Multi-Scale)
As a heuristic proxy for the richness of the learning signal, we compare the density of direct token-level supervision signals and how they accumulate over optimization.
Multi-scale notation.
Let the multi-scale CamS sequence consist of $S$ scales with lengths $\{T_{s}\}_{s=1}^{S}$ and total tokens $T_{\text{all}}=\sum_{s=1}^{S}T_{s}$.
Where gradients originate.
Let $\mathcal{L}$ be the token-level loss.
  • MNP: Direct loss terms (hence gradient sources) occur only on masked indices $i\in\mathcal{M}$; unmasked tokens receive gradients only indirectly via attention coupling.
  • NTP: Direct loss terms occur at (almost) all positions $t\in\{1,\dots,T_{\text{all}}-1\}$.
Direct supervision density and accumulation.
We define the Direct Supervision Density (SD) as the expected fraction of tokens that serve as prediction targets per update:
$\text{SD}_{\text{NTP}}\approx 1,\qquad\text{SD}_{\text{MNP}}=\rho.$ (16)
Equivalently, over $K$ optimization steps, the number of direct token-level targets scales as
$\mathcal{R}_{\text{NTP}}\propto K\cdot T_{\text{all}},\qquad\mathcal{R}_{\text{MNP}}\propto K\cdot\rho\,T_{\text{all}}.$ (17)
Thus, NTP provides about $1/\rho$ times more direct targets than MNP (e.g., $6.7\times$ when $\rho=0.15$).
Trade-off and practical masking rates.
Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off in masked modeling (wettig2023should). In molecular graph Transformers, KPGT reports that setting the masking rate to $\rho=0.5$ can be optimal when token design and global knowledge tokens help stabilize the conditioning context (li2023knowledge). More broadly, structured masking / task design in graph pretraining (e.g., GROVER) can be viewed as attempts to obtain more useful target coverage without excessively destabilizing the evidence set (rong2020self).
Why dense (and repeated) supervision can help.
Dense token-level objectives are known to improve sample efficiency compared to losses defined only on a small masked subset; ELECTRA explicitly attributes a major part of MLM’s inefficiency to learning from only masked tokens rather than all positions (clark2020electra). CamS strengthens this further via multi-view repetition: the same molecule contributes token-level targets repeatedly across scales, inducing cross-scale consistency pressure and providing redundant evidence pathways (a motif may appear as fine tokens at small scales and as a merged token at larger scales). This reduces sensitivity to any single tokenization granularity and makes the effective supervision more robust.
B.3 Structural Bias Formulation
We distinguish the structural injection mechanism in Eq. (11).
Graphformer (Hard Static Bias).
The bias $\psi_{ij}$ is defined via static structural encodings (e.g., shortest-path distance), injected into attention as an additive bias term (ying2021transformers):
$\psi_{ij}^{\text{Graph}}=b_{\phi(d_{G}(i,j))}.$ (18)
CamS (Hard Causal Constraint + Soft Learned Aggregation).
CamS uses a hard causal mask to enforce fine-to-coarse information flow:
$\psi_{ij}^{\text{CamS}}=\begin{cases}0&\text{if }(i\!\rightarrow\!j)\text{ is allowed by }\mathbf{M}_{\text{causal}}\\ -\infty&\text{otherwise.}\end{cases}$ (19)
Within the permitted region, connectivity is content-adaptive via $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$; hence coarse-scale tokens learn which fine-scale neighborhood details to aggregate, rather than uniformly mixing neighbors at a given distance.
Appendix C Training and Implementation Details
Overview. This section corresponds to the training and evaluation protocols in Secs. 3.2.1 and 4. It supplements the main text with implementation-level details of (1) the shared training framework and parallelization setup for pre-training vs. fine-tuning (Sec. C.1), (2) the complete pre-training configuration on Enamine675M (MultiScaler3375M) (Table 5), (3) the fixed fine-tuning setup used across MoleculeNet and MoleculeACE (Table 6), and (4) the unified hyperparameter search space (including strategy-specific grids) (Table 7).
C.1 Pre-training and Fine-tuning Hyperparameters
Implementation and framework.
All experiments are implemented in PyTorch using HuggingFace Transformers. Pre-training uses the built-in Trainer with DeepSpeed ZeRO-2 on multi-GPU data parallelism; fine-tuning is run as single-GPU jobs (one job per GPU) using the same Trainer API (or strategy-specific subclasses).
Pre-training on Enamine675M (MultiScaler3375M).
We pre-train a 16-layer LLaMA-style decoder initialized from a local config (i.e., random initialization; no pretrained weights are loaded). The training data are pre-tokenized and stored as HuggingFace datasets on disk with train/validation splits. We use FP16 training with a cosine learning-rate schedule and warmup, and evaluate/save checkpoints periodically during training.
Fine-tuning (MoleculeNet and MoleculeACE).
We fine-tune the pretrained backbone with fingerprint injection (fp_dim=2048) and a lightweight supervised head. All downstream runs use FP16, evaluate every epoch, and select the best checkpoint by the primary validation metric (higher-is-better for AUROC; lower-is-better for RMSE). We also tested two dropout application schemes in the MoleculeACE fine-tuning pipeline: (1) head-only dropout (applied only in the supervised head), and (2) backbone-synced dropout (also overriding backbone attention_dropout/hidden_dropout and all nn.Dropout modules). Both schemes share the same dropout-rate grid in Table 7.
Ablation compute budget.
For ablations (e.g., single-scale tokenization or removing fingerprint injection), we keep the optimizer, LR schedule, batch size, and training length the same as the corresponding main setting, and only change the ablated component.
Table 5: Pre-training hyperparameters on Enamine675M (MultiScaler3375M).
Setting Value
Backbone config 16 layers; hidden size 720; 8 heads; MLP 2880
Context length 4096 (max_position_embeddings)
Initialization new model from local config (no pretrained weights)
Tokenizer CamS-tokenizer; embedding resized to tokenizer.vocab_size
Training instances Multi-scale tokenized corpus (MultiScaler3375M; 5 views per molecule)
Parallelism torchrun 8 processes + DeepSpeed ZeRO stage 2
Precision FP16 (fp16=true in Trainer and DeepSpeed)
Micro-batch size 256 per GPU
Gradient accumulation 2
Global batch size 4096 (256 $\times$ 8 $\times$ 2)
Optimizer AdamW (Transformers Trainer default)
Learning rate $1\times 10^{-4}$
Weight decay $1\times 10^{-2}$
LR schedule cosine; warmup_steps =5000=5000
Gradient clipping 1.0 (DeepSpeed gradient_clipping)
Training length 10 epochs in planning (early stop after 1/10 epoch in actrual excution)
Eval / save / log eval_steps =33000=33000; save_steps =33000=33000; logging_steps =3300=3300
Checkpoint cap save_total_limit =100=100
Table 6: Fixed fine-tuning setup for MoleculeNet and MoleculeACE.
Setting MoleculeNet MoleculeACE
Task type classification / regression (per dataset) regression
Data split provided scaffold splits (same as KPGT), fold $\in\{0,1,2\}$ provided scaffold splits (same as KPGT)
Epochs 80 (search phase) 50
Batch size 32 per GPU (default) 32 per GPU (default; a small subset rerun with 64)
Grad accumulation 1 (default) 1(default)
Optimizer AdamW (Trainer default) AdamW (Trainer default)
LR schedule constant_with_warmup, warmup_ratio =0.1=0.1 constant_with_warmup, warmup_ratio =0.1=0.1
Evaluation per epoch per epoch
Primary metric AUROC / macro-AUROC; RMSE for regression tasks RMSE (overall_RMSE)
Table 7: Fine-tuning hyperparameter search space (shared).
Hyperparameter Candidates
Base grid (48 combos) learning_rate $\in\{10^{-6},3\!\times\!10^{-6},10^{-5},3\!\times\!10^{-5}\}$
dropout_rate $\in\{0,0.05,0.1,0.2\}$; weight_decay $\in\{0,10^{-6},10^{-4}\}$
Fine-tune strategy normal, l2sp, llrd, flag, reinit
LLRD layer_decay $\in\{0.95,0.9,0.85,0.8\}$
L2SP weight $\in\{10^{-1},10^{-2},10^{-3},10^{-4}\}$
FLAG $\epsilon\in\{10^{-3},3\!\times\!10^{-3},5\!\times\!10^{-3},10^{-2}\}$
Reinit reinit_num_layers $\in\{1,2,3,4\}$
Appendix D Detailed Experimental Results
Overview. This section corresponds to the benchmark results summarized in Sec. 4. It supplements the main text with the full detailed tables of (1) MoleculeNet overall performance of all baseline and their rank comparison (extended version of Table 1; Table 8), and (2) MoleculeACE results overall performance of all baseline and their rank comparison (extended version of Table 2; Table 9).
Table 8: Full Performance comparison on MoleculeNet benchmark. (Extended version of Table 1). Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting). Baseline Reference: Infomax (velivckovic2018deep); Edgepred (hamilton2017inductive); Attribute Masking (Masking) (hu2020strategies); Context Prediction (Contextpred) (hu2020strategies); GraphLoG (xu2021self); GraphCL (you2020graph); JOAO (you2021graph); GROVER (rong2020self); 3DInfomax (stark20223d); GraphMVP (liu2016pyridazinone); ImageMol (zeng2022accurate); MolFormer (wu2023molformer); GEM (fang2022geometry); GraphMAE (hou2022graphmae); MoleBert (xia2023mole); KPGT (li2023knowledge)
Classification Tasks Regression Tasks
Method AVG-AUROC $\uparrow$ Rank Method AVG-RMSE $\downarrow$ Rank
CamS-LLaMA 0.845 1 CamS-LLaMA 1.172 1
KPGT 0.843 2 KPGT 1.175 2
CamS-LLaMA (w/o FP) 0.838 3 CamS-LLaMA (w/o FP) 1.195 3
CamS-LLaMA (Vocab-1K Only) 0.833 4 CamS-LLaMA (Vocab-1K Only) 1.215 4
GEM 0.825 5 MolFormer 1.272 5
CamS-LLaMA (Vocab-67K Only) 0.818 6 GEM 1.285 6
GROVER 0.818 7 CamS-LLaMA (Vocab-67K Only) 1.329 7
ContextpredSup 0.807 8 GROVER 1.332 8
3DInfomax 0.804 9 3DInfomax 1.400 9
ImageMol 0.802 10 ContextpredSup 1.414 10
MoleBERT 0.802 11 ImageMol 1.501 11
MaskingSup 0.799 12 MoleBERT 1.559 12
EdgepredSup 0.795 13 Contextpred 1.563 13
InfomaxSup 0.794 14 EdgepredSup 1.576 14
JOAO 0.790 15 MaskingSup 1.620 15
Contextpred 0.790 16 GraphMVP 1.647 16
GraphMAE 0.788 17 InfomaxSup 1.661 17
Edgepred 0.785 18 GraphLoG 1.663 18
GraphCL 0.774 19 Edgepred 1.674 19
Infomax 0.773 20 Masking 1.697 20
Masking 0.773 21 GraphMAE 1.716 21
GraphLoG 0.769 22 Infomax 1.770 22
GraphMVP 0.769 23 GraphCL 1.822 23
MolFormer 0.744 24 JOAO 1.960 24
Table 9: Full Detailed results on 30 MoleculeACE activity-cliff tasks (Regression, RMSE $\downarrow$). (Extended version of Table 2). Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting). Descriptors: ECFP (rogers2010extended); MACCS (durant2002reoptimization); PHYSCHEM (walters2002prediction); WHIM (kubinyi19933d). Baseline Reference: support vector machines (SVM) (cristianini2002support); random forest (RF) (breiman1996bagging); gradient boosting machine (GBM) (friedman2001greedy); k-nearest neighbor (KNN) (fix1985discriminatory); message passing neural network (MPNN) (gilmer2017neural); graph attention network (GAT) (velivckovic2017graph); graph convolutional network (GCN) (kipf2016semi); AFP (xiong2019pushing); Convolutional Neural NetworkLong Short-Term Memory (CNN, specifically Maxsmi)  (kimber2021maxsmi); Long Short-Term Memory (LSTM, specifically CLM)  (moret2022perplexity); Transformer (specifically Chemberta) (chithrananda2020chemberta); KPGT (li2023knowledge)
MoleculeACE AVG-RMSE (Ranks 1–24) MoleculeACE AVG-RMSE (Ranks 25–48)
Method AVG-RMSE $\downarrow$ Rank Method AVG-RMSE $\downarrow$ Rank
CamS-LLaMA 0.624 1 EdgePredSup 0.764 25
KPGT 0.633 2 ContextpredSup 0.764 26
CamS-LLaMA (Vocab-1K Only) 0.641 3 GraphMVP 0.768 27
CamS-LLaMA (Vocab-67K Only) 0.649 4 Infomax 0.774 28
CamS-LLaMA (w/o FP) 0.650 5 Edgepred 0.774 29
SVMECFP 0.675 6 ImageMol 0.797 30
GROVER 0.680 7 3DInfomax 0.804 31
GBMECFP 0.701 8 GraphLoG 0.807 32
RFECFP 0.705 9 KNNMACCS 0.818 33
MolFormer 0.740 10 GEM 0.821 34
KNNECFP 0.741 11 Transformer 0.868 35
MLPECFP 0.742 12 RFPHYSCHEM 0.890 36
GBMMACCS 0.742 13 GBMPHYSCHEM 0.901 37
LSTM 0.744 14 KNNPHYSCHEM 0.923 38
Contextpred 0.747 15 SVMPHYSCHEM 0.935 39
RFMACCS 0.753 16 CNN 0.937 40
GraphMAE 0.753 17 MPNN 0.959 41
SVMMACCS 0.754 18 AFP 0.970 42
MaskingSup 0.755 19 RFWHIM 0.977 43
MoleBERT 0.755 20 GBMWHIM 0.992 44
InfomaxSup 0.756 21 SVMWHIM 1.003 45
JOAO 0.757 22 GCN 1.010 46
Masking 0.758 23 KNNWHIM 1.020 47
GraphCL 0.760 24 GAT 1.049 48
Appendix E Interpretability Details for Activity-Cliff Attention Analysis
Overview. This section corresponds to the activity-cliff attention analysis in Sec. 4.3. It supplements the main text with implementation-level details of (1) activity-cliff pair construction and differential/shared atom identification (Appendix E.1), Alg.7, (2) mapping atom-level diff/shared labels to CamS tokens within each scale region (Appendix E.2), and (3) attention extraction and metric computation/aggregation for MDTA/MSTA and Rel-DTAP (Appendix E.3).
E.1 Activity-cliff Pair Construction and Differential/Shared Atom Identification
Pair construction (MoleculeACE-style).
For each MoleculeACE sub-task dataset, we start from its per-molecule CSV containing smiles, exp_mean [nM], cliff_mol, and the split label. We treat molecules with cliff_mol=1 as anchors (by default restricted to the test split), and search partners among molecules with cliff_mol=0 (by default from the same split). A candidate pair is kept if it satisfies both (1) high structural similarity and (2) large potency change. Structural similarity is a “soft-consensus” rule: we compute (a) full-molecule ECFP Tanimoto, (b) generic Murcko-scaffold ECFP Tanimoto, and (c) SMILES Levenshtein similarity, and accept the pair if any of them is $\geq\tau_{\mathrm{sim}}$ (default $0.9$). Potency change is measured by the fold change on linear exp_mean [nM] values: $\mathrm{FC}=\frac{\max(y_{a},y_{p})}{\max(\min(y_{a},y_{p}),\epsilon)}$ (with $\epsilon{=}10^{-12}$), and we keep the pair if $\mathrm{FC}\geq\tau_{\mathrm{fold}}$ (default $10$). Each selected pair yields one record $(\text{anchor},\text{partner})$; for downstream attention statistics we count the two molecules (anchor and partner) separately, as stated in Sec. 4.3.
Differential vs. shared atoms.
Given an (anchor, partner) pair, we identify shared atoms via an RDKit maximum common substructure (MCS) query with ring constraints and chemistry-aware matching (elements compared by type; bonds compared by order; completeRingsOnly and ringMatchesRingOnly enabled). Let $\mathcal{A}$ and $\mathcal{B}$ be the atom-index sets of the two molecules and let $\mathcal{M}\subseteq\mathcal{A}$ and $\mathcal{M}^{\prime}\subseteq\mathcal{B}$ be the matched atom indices returned by the first substructure match of the MCS query. We define differential atoms as the complement sets $\mathcal{A}_{\Delta}=\mathcal{A}\setminus\mathcal{M}$ and $\mathcal{B}_{\Delta}=\mathcal{B}\setminus\mathcal{M}^{\prime}$, and shared atoms as $\mathcal{M}$ and $\mathcal{M}^{\prime}$. If MCS fails (or a SMILES cannot be parsed), we conservatively treat all atoms as differential.
Atom correspondence (optional for visualization).
For visualization/debugging we also record the ordered atom correspondence induced by the MCS query, i.e., a list of paired indices $\{(i,j)\}$ obtained by aligning the two match tuples in the MCS SMARTS query order. This mapping is not required for computing Rel-DTAP, but enables atom-level highlighting across the two molecules.
E.2 Mapping Atom-level Diff/Shared Labels to Tokens in Each Scale Region
Token-to-atom alignment.
We re-encode each molecule with MultiGraphBPETokenizerExplain, which returns, for every token position $t$, an atom-index set $S_{t}$ indicating which RDKit atom indices are covered by that token (special tokens such as [BOS], [EOS], and [CONCAT] have empty sets). Because GraphBPE encoding inserts auxiliary connector atoms (*) with indices outside the original RDKit atom range, we remove such indices by clipping $S_{t}$ to $\{0,\dots,N{-}1\}$ where $N$ is the number of atoms in the original molecule.
From atom-level labels to token-level labels.
Let $\mathcal{A}_{\Delta}$ be the differential-atom set for this molecule from Appendix E.1. We assign a token-level differential indicator by an any-diff rule: a token at position $t$ is marked as differential iff $S_{t}\cap\mathcal{A}_{\Delta}\neq\varnothing$; otherwise it is marked as shared/non-differential. Thus a token covering multiple atoms is counted as differential if it contains at least one differential atom.
Scale-region boundaries in the concatenated sequence.
For multi-scale CamS sequences, scale regions are delimited by the [CONCAT] separators. Concretely, given the full token-id list (including [BOS], [EOS], and [CONCAT]), we split it into consecutive spans between [CONCAT] tokens, and exclude [BOS]/[EOS]/[CONCAT] themselves from the per-region statistics. This yields the four scale regions (e.g., 1K/7K/27K/67K in the paper) used to report scale-wise attention preferences.
E.3 Attention Extraction and Metric Computation
Attention extraction protocol.
For each molecule, we run a forward pass with output_attentions=True and extract the last-layer attention tensor. We average over heads to obtain a single attention matrix $\bar{\mathbf{A}}\in\mathbb{R}^{S\times S}$, where $S$ is the sequence length. We use the attention distribution of the final token (the last position in the sequence) as a saliency-like weighting over the prefix tokens, i.e., the row vector $\bar{\mathbf{A}}_{S,:}$. We compute this distribution in two modes: (1) without fingerprint input by running the model without the prepended fingerprint token; (2) with fingerprint input by prepending the fingerprint embedding (sequence length becomes S+1S{+}1), extracting $\bar{\mathbf{A}}_{S{+}1,:}$, and then taking only the sub-vector over molecular tokens (excluding the fingerprint position) for diff/shared statistics.
MDTA/MSTA within a scale region.
Within a given scale region $s$ (defined by [CONCAT] boundaries), let $p_{i}$ denote the (renormalized) attention weight on token $i$ in that region, and let $d_{i}\in\{0,1\}$ be its differential indicator from Appendix E.2. We define the mean differential-token attention and mean shared-token attention as
$\mathrm{MDTA}_{s}=\frac{\sum_{i}p_{i}d_{i}}{\sum_{i}d_{i}},\qquad\mathrm{MSTA}_{s}=\frac{\sum_{i}p_{i}(1-d_{i})}{\sum_{i}(1-d_{i})},$
with the convention that the corresponding mean is set to 0 if the denominator is 0 (no tokens of that type in the region).
Rel-DTAP computation and aggregation.
For each molecule (anchor and partner counted separately), we compute
$\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100,$
using $\epsilon=10^{-12}$ for numerical stability. We report the final Rel-DTAP by averaging this quantity over all molecules from activity-cliff pairs, both for each scale region $s$ and for the full concatenated sequence (all regions combined).
Algorithm 7 Shared/differential fragment labeling for an activity-cliff pair
1:Input: SMILES pair $(s_{a},s_{b})$; RDKit MCS; tokenizer-explain encoder $\iota$ returning per-token atom sets
2:Output: shared/diff atom sets $(M_{a},\Delta_{a})$ and $(M_{b},\Delta_{b})$; token-level diff masks $\mathbf{d}^{(a)},\mathbf{d}^{(b)}$; optional atom map $\mathcal{P}$
[…]
5:     completeRingsOnly, ringMatchesRingOnly,
5:     CompareElements, CompareOrder).smartsString
6:if smartssmarts is empty then
7:  $M_{a}\leftarrow\varnothing$; $M_{b}\leftarrow\varnothing$
8:else
9:  $q\leftarrow\mathrm{MolFromSmarts}(smarts)$
10:  $\mathbf{t}_{a}\leftarrow m_{a}.\mathrm{GetSubstructMatch}(q)$; $\mathbf{t}_{b}\leftarrow m_{b}.\mathrm{GetSubstructMatch}(q)$ (take the first match)
11:  $M_{a}\leftarrow\mathrm{set}(\mathbf{t}_{a})$; $M_{b}\leftarrow\mathrm{set}(\mathbf{t}_{b})$
12:  $\mathcal{P}\leftarrow\{(\mathbf{t}_{a}[i],\mathbf{t}_{b}[i])\}_{i=1}^{|\mathbf{t}_{a}|}$ (ordered atom correspondence)
13:end if
[…]
15:for side $m\in\{a,b\}$ do
16:  Encode $s_{m}$ with $\iota$ to get tokens $(x^{(m)}_{1},\dots,x^{(m)}_{L_{m}})$ and atom sets $\{S^{(m)}_{t}\}_{t=1}^{L_{m}}$
17:  Clip connector atoms: $S^{(m)}_{t}\leftarrow S^{(m)}_{t}\cap\{0,\dots,N_{m}{-}1\}$ for all $t$
18:  for $t=1$ to $L_{m}$ do
19:   if $x^{(m)}_{t}$ is [BOS] or [EOS] or [CONCAT] then
20:    $d^{(m)}_{t}\leftarrow 0$
21:   else
22:    if $S^{(m)}_{t}\cap\Delta_{m}\neq\varnothing$ then
23:     $d^{(m)}_{t}\leftarrow 1$ (diff fragment token; any-diff rule)
24:    else
25:     $d^{(m)}_{t}\leftarrow 0$ (shared fragment token)
26:    end if
27:   end if
28:  end for
29:end for

### v2

@@
 1 Introduction
-Molecular property prediction is a core challenge in drug discovery (wu2018moleculenet) and has increasingly become a primary benchmark for “Foundation Models (FM) for Science.” (khan2025comprehensive) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (wang2024emu3; xia2025nature; zhang2025unigenx).
-Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (touvron2023llama).
-However, a critical gap remains: current NTP-based molecular models, which largely rely on 1D SMILES strings, fail to explicitly capture graph topology and consequently trail behind specialist graph models in predictive accuracy (xia2025nature).
-Conversely, graph-native methods explicitly model topology but typically depart from pure decoder-only NTP, instead adopting hybrid architectures or objectives that involve, to varying degrees, input corruption (e.g., masked node prediction) (li2023knowledge; shehzad2024graph; lu2025uni; kong2025unimomo).
-While sophisticated strategies—such as incorporating global knowledge nodes or avoiding random masking—have been developed to mitigate the information loss caused by masking (li2023knowledge; rong2020self; liu2024mask; you2020graph; xu2021self), these methods remain fundamentally bound by the corruption-coverage trade-off (wettig2023should).
+Molecular property prediction is a core challenge in drug discovery (Wu et al., 2018) and has increasingly become a primary benchmark for “Foundation Models (FM) for Science.” (Khan et al., 2025) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (Wang et al., 2024; Xia et al., 2025; Zhang et al., 2025).
+Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (Touvron et al., 2023).
+However, a critical gap remains: current NTP-based molecular models, which largely rely on 1D SMILES strings, fail to explicitly capture graph topology and consequently trail behind specialist graph models in predictive accuracy (Xia et al., 2025).
+Conversely, graph-native methods explicitly model topology but typically depart from pure decoder-only NTP, instead adopting hybrid architectures or objectives that involve, to varying degrees, input corruption (e.g., masked node prediction) (Li et al., 2023; Shehzad et al., 2024; Lu et al., 2025; Kong et al., 2025).
+While sophisticated strategies—such as incorporating global knowledge nodes or avoiding random masking—have been developed to mitigate the information loss caused by masking (Li et al., 2023; Rong et al., 2020; Liu et al., 2024; You et al., 2020; Xu et al., 2021), these methods remain fundamentally bound by the corruption-coverage trade-off (Wettig et al., 2023).
 This limitation is particularly acute in activity-cliff scenarios, where a single atomic change triggers a drastic property shift; masking the local neighborhood of a critical functional group effectively removes the precise evidence needed to discern such subtle differences.
-Standard NTP, by contrast, offers a dense supervision signal without corrupting the visible prefix context (clark2020electra), yet it lacks a representation interface that exposes graph connectivity as effectively as text.
+Standard NTP, by contrast, offers a dense supervision signal without corrupting the visible prefix context (Clark et al., 2020), yet it lacks a representation interface that exposes graph connectivity as effectively as text.
 To bridge this gap, we propose Connection-Aware Motif Sequencing (CamS), a tokenizer-level interface that makes molecular graphs directly learnable by standard decoder-only NTP through a three-stage design.
-First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (shibata1999byte; geng2023novo; shen2024graphbpe).
+First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (Shibata et al., 1999; Geng et al., 2023; Shen and Poczos, ).
 This produces motif vocabulary with frequent merge operation and introduces a tunable molecular scale that allows direct granularity control.
 To ensure faithful encoding of fine-grained chemical detail, we further integrate a Single-Atom Vocabulary Closure (SAVC) mechanism.
-Second, to make the graph autoregressive-ready, we partition the molecule into a BPEGraph at a specific scale and serialize it via scaffold-rooted breadth-first search (BFS)  (bundy1984breadth).
+Second, to make the graph autoregressive-ready, we partition the molecule into a BPEGraph at a specific scale and serialize it via scaffold-rooted breadth-first search (BFS)  (Bundy and Wallen, 1984).
 This produces a CamS subsequence with a stable Intra-scale Order, moving from the global core to peripheral functional groups.
@@
 Collectively, these steps unify substructure compression and serialization through a dual-causal ordering—sequencing motifs within each scale and stacking scales hierarchically—to enable the model to comprehend global scaffolds based on explicit, uncorrupted local structural evidence.
-Built on CamS, we instantiate CamS-LLaMA by pre-training a native decoder-only backbone with standard NTP on CamS sequences, keeping the architecture and objective unchanged (touvron2023llama).
-To keep attribution controlled, we inject only a lightweight molecular fingerprint (cereto2015molecular) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven.
-Across MoleculeNet (wu2018moleculenet) and the activity-cliff stress test MoleculeACE (van2022exposing), CamS-LLaMA achieves state-of-the-art performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (li2023knowledge).
+Built on CamS, we instantiate CamS-LLaMA by pre-training a native decoder-only backbone with standard NTP on CamS sequences, keeping the architecture and objective unchanged (Touvron et al., 2023).
+To keep attribution controlled, we inject only a lightweight molecular fingerprint (Cereto-Massagué et al., 2015) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven.
+Across MoleculeNet (Wu et al., 2018) and the activity-cliff stress test MoleculeACE (Van Tilborg et al., 2022), CamS-LLaMA achieves state-of-the-art performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (Li et al., 2023).
 Beyond accuracy, we provide mechanism-level evidence.
@@
 2.1 Molecular Property Prediction within Foundation Models Paradigm
-Molecular property prediction is increasingly adopting a foundation model approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (awais2025foundation; xia2025nature).
-One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (xia2025nature; cai2025chemfm).
+Molecular property prediction is increasingly adopting a foundation model approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (Awais et al., 2025; Xia et al., 2025).
+One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (Xia et al., 2025; Cai et al., 2025).
 This leverages mature LLM infrastructure and enables straightforward large-scale training.
 However, SMILES is not structure-native, weakening topology-based signals.
-String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (xia2025nature).
-Alternatively, graph-native foundation models directly process molecular graphs to preserve connectivity and local topology (you2021graph; liupre; you2020graph; xu2021self).
-High-performing methods typically employ encoder-only Graph Transformers (shehzad2024graph) designed for property prediction.
-While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (li2023knowledge; rong2020self).
-More recent structural foundation models integrate graphs with autoregressive components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (kong2025unimomo; lu2025uni; zhang2025unigenx).
+String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (Xia et al., 2025).
+Alternatively, graph-native foundation models directly process molecular graphs to preserve connectivity and local topology (You et al., 2021; Liu et al., ; You et al., 2020; Xu et al., 2021).
+High-performing methods typically employ encoder-only Graph Transformers (Shehzad et al., 2024) designed for property prediction.
+While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (Li et al., 2023; Rong et al., 2020).
+More recent structural foundation models integrate graphs with autoregressive components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (Kong et al., 2025; Lu et al., 2025; Zhang et al., 2025).
 2.2 Molecular Tokenization Strategies
@@
 In NLP, BPE yields strong compression via simple frequency-based merges without training an extra tokenizer model.
-Molecular tokenization spans string-based and graph-based methods (wang2019smiles; liu2023rethinking).
-String-based schemes operate on SMILES (weininger1988smiles) or related linearizations (krenn2022selfies), while graph-based schemes have evolved from basic node/edge-level units to fragment-level motifs (liu2023rethinking).
-Some rule-based motif tokenizations rely on limited handcrafted chemistry rules (zhang2021motif); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (li2023knowledge); ring-or-path-based (wollschlager2024expressivity) and tree-based (jin2018junction) schemes also expose substructures at different granularities; some other works learn tokenization via trainable rules (sun2024representing) or by training a GNN as a tokenizer (liu2023rethinking; luonode).
-Particularly noteworthy are data-driven motif mining approaches that adapt BPE idea to graphs to obtain reusable, multi-scale fragments with strong compression and minimal reliance on manually designed rules or additional trained models (geng2023novo; shen2024graphbpe; kong2022molecule).
-In addition, cross-level tokenization schemes, which combine motifs at different structural levels, also provide complementary design ideas (ji2022relmole; chenhierarchical).
+Molecular tokenization spans string-based and graph-based methods (Wang et al., 2019; Liu et al., 2023).
+String-based schemes operate on SMILES (Weininger, 1988) or related linearizations (Krenn et al., 2022), while graph-based schemes have evolved from basic node/edge-level units to fragment-level motifs (Liu et al., 2023).
+Some rule-based motif tokenizations rely on limited handcrafted chemistry rules (Zhang et al., 2021); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (Li et al., 2023); ring-or-path-based (Wollschläger et al., 2024) and tree-based (Jin et al., 2018) schemes also expose substructures at different granularities; some other works learn tokenization via trainable rules (Sun et al., 2024) or by training a GNN as a tokenizer (Liu et al., 2023; Luo et al., ).
+Particularly noteworthy are data-driven motif mining approaches that adapt BPE idea to graphs to obtain reusable, multi-scale fragments with strong compression and minimal reliance on manually designed rules or additional trained models (Geng et al., 2023; Shen and Poczos, ; Kong et al., 2022).
+In addition, cross-level tokenization schemes, which combine motifs at different structural levels, also provide complementary design ideas (Ji et al., 2022; Chen et al., ).
 3 Method
-Our molecular representation must satisfy two requirements: (1) chemical fidelity, i.e., sensitivity to pharmacophore-level structures and their topology; and (2) autoregressive foundation model compatibility, i.e., a compressed interface naturally aligned with an NTP paradigm (touvron2023llama).
+Our molecular representation must satisfy two requirements: (1) chemical fidelity, i.e., sensitivity to pharmacophore-level structures and their topology; and (2) autoregressive foundation model compatibility, i.e., a compressed interface naturally aligned with an NTP paradigm (Touvron et al., 2023).
 Motivated by this, we propose Connection-Aware Motif Sequencing (CamS), which segments molecular graphs into reusable frequent motifs and, crucially, serializes the motif graph into multi-scale connection-aware causal sequences compatible with NTP pre-training (Section 3.1).
@@
 BPE-style Mining.
-Following (geng2023novo; shen2024graphbpe), we first learn an ordered list of merge operations $\mathcal{O}$ from co-occurrence statistics.
+Following (Geng et al., 2023; Shen and Poczos, ), we first learn an ordered list of merge operations $\mathcal{O}$ from co-occurrence statistics.
 We then apply $\mathcal{O}$ to the corpus to mine the resulting subgraphs.
@@
 We select the motif with the largest atom count (typically the core scaffold) as the root.
-This strategy prioritizes global backbone structure before local substitutions (zhang2021motif), establishing a stable Center-to-Periphery causal order.
+This strategy prioritizes global backbone structure before local substitutions (Zhang et al., 2021), establishing a stable Center-to-Periphery causal order.
 ID Extraction. Mapping the ordered nodes to their indices in $\Sigma$ yields the single-scale CamS subsequence $X^{(s)}=(x^{(s)}_{1},\dots,x^{(s)}_{L_{s}})$.
@@
 3.2.1 Autoregressive Pre-training via NTP
-CamS-LLaMA adopts a standard LLaMA-style decoder (touvron2023llama); in pre-training, every CamS token list is treated as an ordinary token sequence in the shared vocabulary $\Sigma$ and is consumed by the backbone under the usual causal mask, with no graph-specific architectural modification.
+CamS-LLaMA adopts a standard LLaMA-style decoder (Touvron et al., 2023); in pre-training, every CamS token list is treated as an ordinary token sequence in the shared vocabulary $\Sigma$ and is consumed by the backbone under the usual causal mask, with no graph-specific architectural modification.
 Concretely, each molecule yields five CamS sequence views $\{X_{1},X_{7},X_{27},X_{67},X\}$: four single-scale BFS sequences tokenized with $\Sigma_{1\mathrm{K}}/\Sigma_{7\mathrm{K}}/\Sigma_{27\mathrm{K}}/\Sigma_{67\mathrm{K}}$, plus the concatenated multi-scale sequence $X$ in Eq.
@@
 3.2.2 Fine-tuning with Fingerprint Injection
-Many molecular pre-training methods inject rich handcrafted descriptors early (e.g., during pre-training) (li2023knowledge).
+Many molecular pre-training methods inject rich handcrafted descriptors early (e.g., during pre-training) (Li et al., 2023).
 In contrast, we keep CamS as the primary representation and inject only one widely-used structural prior–a fingerprint–only during fine-tuning, to make the contribution of CamS representations more attributable and controlled.
@@
 $\mathcal{V}_{\text{tok}}$ corresponds to single-scale atoms or triplets.
-$\psi_{ij}$ encodes static structural biases (e.g., shortest-path-distance bias injected into attention) independent of token content, yielding an isotropic receptive field at a given graph distance (ying2021transformers).
+$\psi_{ij}$ encodes static structural biases (e.g., shortest-path-distance bias injected into attention) independent of token content, yielding an isotropic receptive field at a given graph distance (Ying et al., 2021).
 CamS-LLaMA (Hierarchical/Causal Flow).
@@
 Sketch.
-The inequality follows from the Data Processing Inequality (DPI): masking is a stochastic post-processing of $Z_{t}$ that cannot increase mutual information (cover1999elements).
+The inequality follows from the Data Processing Inequality (DPI): masking is a stochastic post-processing of $Z_{t}$ that cannot increase mutual information (Cover, 1999).
 In graph MNP, this loss of information is exacerbated by evidence-pattern uncertainty: the most relevant local-neighborhood evidence for a masked token may be co-masked, making the conditional distribution unstable across masking patterns.
-Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets / local subgraphs) or designing structure-guided masking, and by injecting global knowledge tokens (li2023knowledge; rong2020self; liu2024mask).
+Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets / local subgraphs) or designing structure-guided masking, and by injecting global knowledge tokens (Li et al., 2023; Rong et al., 2020; Liu et al., 2024).
 In contrast, CamS NTP for coarse-scale tokens conditions on an uncorrupted fine-to-coarse history, avoiding such stochastic evidence loss.
-This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (yang2019xlnet).
+This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (Yang et al., 2019).
 3.3.3 Gradient Flow Density.
@@
 This trade-off (more predicted tokens vs.
-more corruption) is central to masked modeling (wettig2023should).
-Empirically, KPGT finds the best masking ratio can be as high as $\rho=0.5$ when using triplet-style tokens and global knowledge nodes to stabilize context (li2023knowledge).
-Meanwhile, dense token-level learning signals are known to substantially improve sample efficiency compared to sparse masked-only supervision (clark2020electra).
+more corruption) is central to masked modeling (Wettig et al., 2023).
+Empirically, KPGT finds the best masking ratio can be as high as $\rho=0.5$ when using triplet-style tokens and global knowledge nodes to stabilize context (Li et al., 2023).
+Meanwhile, dense token-level learning signals are known to substantially improve sample efficiency compared to sparse masked-only supervision (Clark et al., 2020).
 CamS further strengthens this effect by providing multi-view supervision: the same molecule contributes token-level targets repeatedly across scales.
@@
 CamS-Tokenizer Training.
-CamS-tokenizer training is expensive because replaying $\mathcal{O}$ to materialize $\mathcal{V}$ repeatedly invokes RDKit’s subgraph-isomorphism matching (the dominant bottleneck) (landrum2021rdkit; ehrlich2011maximum) to extract motifs and attachment patterns, which is difficult to accelerate.
-We therefore train the tokenizer once on ChEMBL-34 (gaulton2017chembl), a moderately sized yet chemically diverse corpus, at the largest scale (67K).
+CamS-tokenizer training is expensive because replaying $\mathcal{O}$ to materialize $\mathcal{V}$ repeatedly invokes RDKit’s subgraph-isomorphism matching (the dominant bottleneck) (Landrum et al., 2021; Ehrlich and Rarey, 2011) to extract motifs and attachment patterns, which is difficult to accelerate.
+We therefore train the tokenizer once on ChEMBL-34 (Gaulton et al., 2017), a moderately sized yet chemically diverse corpus, at the largest scale (67K).
 Concretely, we mine an ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K})$ with K=685K{=}685 merges, apply it to obtain $\mathcal{V}$, and then perform SAVC to ensure single-atom coverage for common elements.
@@
 We instantiate a 16-layer LLaMA-style decoder (hidden size 720, 8 attention heads) using HuggingFace Transformers.
-For self-supervised pre-training, we use Enamine675M (675M molecules) (shivanyuk2007enamine) and optimize NTP (Section 3.2.1).
+For self-supervised pre-training, we use Enamine675M (675M molecules) (Shivanyuk et al., 2007) and optimize NTP (Section 3.2.1).
 For each molecule, we construct five CamS training sequences: four single-scale BFS sequences tokenized with 1K/7K/27K/67K vocabularies, plus the concatenated multi-scale sequence in Eq.
@@
 Downstream Benchmarks and Baselines.
-We evaluate on MoleculeNet (11 tasks; broad endpoints that probe general-purpose molecular property prediction) (wu2018moleculenet) and MoleculeACE (30 activity-cliff tasks; stress-testing fine-grained structural sensitivity under subtle edits that induce large activity shifts) (van2022exposing).
-To enable an apples-to-apples comparison against a recent strong pre-training baseline, we adopt the data splits and evaluation protocol used in KPGT (li2023knowledge).
+We evaluate on MoleculeNet (11 tasks; broad endpoints that probe general-purpose molecular property prediction) (Wu et al., 2018) and MoleculeACE (30 activity-cliff tasks; stress-testing fine-grained structural sensitivity under subtle edits that induce large activity shifts) (Van Tilborg et al., 2022).
+To enable an apples-to-apples comparison against a recent strong pre-training baseline, we adopt the data splits and evaluation protocol used in KPGT (Li et al., 2023).
 On MoleculeNet, following KPGT’s setup, we compare primarily against self-supervised pre-training baselines to isolate the effect of (1) representation/tokenization and (2) the pre-training objective.
@@
 Column: Previous Best Baselines denotes the best-performing method excluding KPGT; Type: ML (Machine Learning), DL (Deep Learning).
-Reference & Abbreviations: GMAE: GraphMAE (hou2022graphmae); MBRT: MoleBERT (xia2023mole); GROV: GROVER (rong2020self); SVM/RF/GBM: Support Vector Machine / Random Forest / Gradient Boosting Machine (van2022exposing).
-Descriptors: E: ECFP (rogers2010extended); M: MACCS (durant2002reoptimization).
+Reference & Abbreviations: GMAE: GraphMAE (Hou et al., 2022); MBRT: MoleBERT (Xia et al., 2023); GROV: GROVER (Rong et al., 2020); SVM/RF/GBM: Support Vector Machine / Random Forest / Gradient Boosting Machine (Van Tilborg et al., 2022).
+Descriptors: E: ECFP (Rogers and Hahn, 2010); M: MACCS (Durant et al., 2002).
 Task Previous Best Baselines KPGT CamS-LLaMA
@@
 Remark (MLM limitations in NLP).
-XLNet points out two intrinsic issues of corruption-based MLM: (1) the special token corruption (e.g., [MASK]) creates a pretrain–finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context (yang2019xlnet).
-Relatedly, ELECTRA shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions, while the [MASK] mismatch contributes a smaller but non-negligible part (clark2020electra).
+XLNet points out two intrinsic issues of corruption-based MLM: (1) the special token corruption (e.g., [MASK]) creates a pretrain–finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context (Yang et al., 2019).
+Relatedly, ELECTRA shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions, while the [MASK] mismatch contributes a smaller but non-negligible part (Clark et al., 2020).
 Why this matters more on graphs (evidence-pattern uncertainty).
@@
 Random masking can therefore remove not only the target token but also its critical neighborhood evidence, making the missingness pattern itself a stochastic nuisance.
-This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability: (1) masking structured units and adding global knowledge nodes (KPGT) (li2023knowledge), (2) masking local subgraphs and predicting contextual properties (GROVER) (rong2020self), and (iii) structure-guided / curriculum masking rather than purely random masking (StructMAE) (liu2024mask).
-These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (wettig2023should).
+This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability: (1) masking structured units and adding global knowledge nodes (KPGT) (Li et al., 2023), (2) masking local subgraphs and predicting contextual properties (GROVER) (Rong et al., 2020), and (iii) structure-guided / curriculum masking rather than purely random masking (StructMAE) (Liu et al., 2024).
+These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (Wettig et al., 2023).
 CamS implication.
@@
 Trade-off and practical masking rates.
-Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off in masked modeling (wettig2023should).
-In molecular graph Transformers, KPGT reports that setting the masking rate to $\rho=0.5$ can be optimal when token design and global knowledge tokens help stabilize the conditioning context (li2023knowledge).
-More broadly, structured masking / task design in graph pretraining (e.g., GROVER) can be viewed as attempts to obtain more useful target coverage without excessively destabilizing the evidence set (rong2020self).
+Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off in masked modeling (Wettig et al., 2023).
+In molecular graph Transformers, KPGT reports that setting the masking rate to $\rho=0.5$ can be optimal when token design and global knowledge tokens help stabilize the conditioning context (Li et al., 2023).
+More broadly, structured masking / task design in graph pretraining (e.g., GROVER) can be viewed as attempts to obtain more useful target coverage without excessively destabilizing the evidence set (Rong et al., 2020).
 Why dense (and repeated) supervision can help.
-Dense token-level objectives are known to improve sample efficiency compared to losses defined only on a small masked subset; ELECTRA explicitly attributes a major part of MLM’s inefficiency to learning from only masked tokens rather than all positions (clark2020electra).
+Dense token-level objectives are known to improve sample efficiency compared to losses defined only on a small masked subset; ELECTRA explicitly attributes a major part of MLM’s inefficiency to learning from only masked tokens rather than all positions (Clark et al., 2020).
 CamS strengthens this further via multi-view repetition: the same molecule contributes token-level targets repeatedly across scales, inducing cross-scale consistency pressure and providing redundant evidence pathways (a motif may appear as fine tokens at small scales and as a merged token at larger scales).
@@
 Graphformer (Hard Static Bias).
-The bias $\psi_{ij}$ is defined via static structural encodings (e.g., shortest-path distance), injected into attention as an additive bias term (ying2021transformers):
+The bias $\psi_{ij}$ is defined via static structural encodings (e.g., shortest-path distance), injected into attention as an additive bias term (Ying et al., 2021):
 $\psi_{ij}^{\text{Graph}}=b_{\phi(d_{G}(i,j))}.$ (18)
@@
 Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting).
-Baseline Reference: Infomax (velivckovic2018deep); Edgepred (hamilton2017inductive); Attribute Masking (Masking) (hu2020strategies); Context Prediction (Contextpred) (hu2020strategies); GraphLoG (xu2021self); GraphCL (you2020graph); JOAO (you2021graph); GROVER (rong2020self); 3DInfomax (stark20223d); GraphMVP (liu2016pyridazinone); ImageMol (zeng2022accurate); MolFormer (wu2023molformer); GEM (fang2022geometry); GraphMAE (hou2022graphmae); MoleBert (xia2023mole); KPGT (li2023knowledge)
+Baseline Reference: Infomax (Veličković et al., 2018); Edgepred (Hamilton et al., 2017); Attribute Masking (Masking) (Hu et al., 2020); Context Prediction (Contextpred) (Hu et al., 2020); GraphLoG (Xu et al., 2021); GraphCL (You et al., 2020); JOAO (You et al., 2021); GROVER (Rong et al., 2020); 3DInfomax (Stärk et al., 2022); GraphMVP (Liu et al., 2016); ImageMol (Zeng et al., 2022); MolFormer (Wu et al., 2023); GEM (Fang et al., 2022); GraphMAE (Hou et al., 2022); MoleBert (Xia et al., 2023); KPGT (Li et al., 2023)
 Classification Tasks Regression Tasks
@@
 Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting).
-Descriptors: ECFP (rogers2010extended); MACCS (durant2002reoptimization); PHYSCHEM (walters2002prediction); WHIM (kubinyi19933d).
-Baseline Reference: support vector machines (SVM) (cristianini2002support); random forest (RF) (breiman1996bagging); gradient boosting machine (GBM) (friedman2001greedy); k-nearest neighbor (KNN) (fix1985discriminatory); message passing neural network (MPNN) (gilmer2017neural); graph attention network (GAT) (velivckovic2017graph); graph convolutional network (GCN) (kipf2016semi); AFP (xiong2019pushing); Convolutional Neural NetworkLong Short-Term Memory (CNN, specifically Maxsmi)  (kimber2021maxsmi); Long Short-Term Memory (LSTM, specifically CLM)  (moret2022perplexity); Transformer (specifically Chemberta) (chithrananda2020chemberta); KPGT (li2023knowledge)
+Descriptors: ECFP (Rogers and Hahn, 2010); MACCS (Durant et al., 2002); PHYSCHEM (Walters and Murcko, 2002); WHIM (Kubinyi, 1993).
+Baseline Reference: support vector machines (SVM) (Cristianini and Scholkopf, 2002); random forest (RF) (Breiman, 1996); gradient boosting machine (GBM) (Friedman, 2001); k-nearest neighbor (KNN) (Fix, 1985); message passing neural network (MPNN) (Gilmer et al., 2017); graph attention network (GAT) (Veličković et al., 2017); graph convolutional network (GCN) (Kipf, 2016); AFP (Xiong et al., 2019); Convolutional Neural NetworkLong Short-Term Memory (CNN, specifically Maxsmi)  (Kimber et al., 2021); Long Short-Term Memory (LSTM, specifically CLM)  (Moret et al., 2022); Transformer (specifically Chemberta) (Chithrananda et al., 2020); KPGT (Li et al., 2023)
 MoleculeACE AVG-RMSE (Ranks 1–24) MoleculeACE AVG-RMSE (Ranks 25–48)

### v3

@@
 1 Introduction
-Molecular property prediction is a core challenge in drug discovery (Wu et al., 2018) and has increasingly become a primary benchmark for “Foundation Models (FM) for Science.” (Khan et al., 2025) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (Wang et al., 2024; Xia et al., 2025; Zhang et al., 2025).
+Molecular property prediction is a core challenge in drug discovery (Wu et al., 2018) and has increasingly become a primary benchmark for “Foundation Models (FMs) of science.” (Khan et al., 2025) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (Wang et al., 2024; Xia et al., 2025; Zhang et al., 2025).
 Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (Touvron et al., 2023).
@@
 To bridge this gap, we propose Connection-Aware Motif Sequencing (CamS), a tokenizer-level interface that makes molecular graphs directly learnable by standard decoder-only NTP through a three-stage design.
-First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (Shibata et al., 1999; Geng et al., 2023; Shen and Poczos, ).
+First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (Shibata et al., 1999; Geng et al., 2023; Shen and Poczos, 2024).
 This produces motif vocabulary with frequent merge operation and introduces a tunable molecular scale that allows direct granularity control.
@@
 To keep attribution controlled, we inject only a lightweight molecular fingerprint (Cereto-Massagué et al., 2015) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven.
-Across MoleculeNet (Wu et al., 2018) and the activity-cliff stress test MoleculeACE (Van Tilborg et al., 2022), CamS-LLaMA achieves state-of-the-art performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (Li et al., 2023).
+Across MoleculeNet (Wu et al., 2018) and the activity-cliff stress test MoleculeACE (Van Tilborg et al., 2022), CamS-LLaMA achieves state-of-the-art (SOTA) performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (Li et al., 2023).
 Beyond accuracy, we provide mechanism-level evidence.
@@
 Contributions.
-(1) Methodological Framework: We introduce CamS, a unified graph-to-causal-sequence interface that resolves the conflict between graph topology preservation and scalable autoregressive training, enabling NTP to function as a structure-native objective.
-(2) Empirical Validation: We demonstrate that CamS-LLaMA achieves state-of-the-art performance on MoleculeNet and MoleculeACE benchmarks.
+(1) Methodological Framework: We introduce CamS, a unified graph-to-causal-sequence interface that resolves the conflict between graph topology preservation and scalable autoregressive (AR) training, enabling NTP to function as a structure-native objective.
+(2) Empirical Validation: We demonstrate that CamS-LLaMA achieves SOTA performance on MoleculeNet and MoleculeACE benchmarks.
 Mechanism analysis reveals that the multi-scale causal context explicitly improves attention focus on subtle, activity-determining structural edits.
-(3) Implementation Recipe: We provide a reproducible pipeline—from the CamS-tokenizer to the CamS-LLaMA—establishing a robust baseline for applying standard foundation model architectures to molecular science.
+(3) Implementation Recipe: We provide a reproducible pipeline—from the CamS-tokenizer to the CamS-LLaMA—establishing a robust baseline for applying standard AR FM architectures to molecular science.
 2 Related Work
-2.1 Molecular Property Prediction within Foundation Models Paradigm
-Molecular property prediction is increasingly adopting a foundation model approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (Awais et al., 2025; Xia et al., 2025).
+2.1 Molecular Property Prediction via FMs
+Molecular property prediction is increasingly adopting a FM approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (Awais et al., 2025; Xia et al., 2025).
 One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (Xia et al., 2025; Cai et al., 2025).
@@
 String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (Xia et al., 2025).
-Alternatively, graph-native foundation models directly process molecular graphs to preserve connectivity and local topology (You et al., 2021; Liu et al., ; You et al., 2020; Xu et al., 2021).
+Alternatively, graph-native FMs directly process molecular graphs to preserve connectivity and local topology (You et al., 2021; Liu et al., 2022; You et al., 2020; Xu et al., 2021).
 High-performing methods typically employ encoder-only Graph Transformers (Shehzad et al., 2024) designed for property prediction.
 While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (Li et al., 2023; Rong et al., 2020).
-More recent structural foundation models integrate graphs with autoregressive components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (Kong et al., 2025; Lu et al., 2025; Zhang et al., 2025).
+More recent structural FMs integrate graphs with AR components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (Kong et al., 2025; Lu et al., 2025; Zhang et al., 2025).
 2.2 Molecular Tokenization Strategies
-Within a foundation model paradigm, tokenizer design becomes a key interface for scaling LLMs.
-In NLP, BPE yields strong compression via simple frequency-based merges without training an extra tokenizer model.
-Molecular tokenization spans string-based and graph-based methods (Wang et al., 2019; Liu et al., 2023).
-String-based schemes operate on SMILES (Weininger, 1988) or related linearizations (Krenn et al., 2022), while graph-based schemes have evolved from basic node/edge-level units to fragment-level motifs (Liu et al., 2023).
-Some rule-based motif tokenizations rely on limited handcrafted chemistry rules (Zhang et al., 2021); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (Li et al., 2023); ring-or-path-based (Wollschläger et al., 2024) and tree-based (Jin et al., 2018) schemes also expose substructures at different granularities; some other works learn tokenization via trainable rules (Sun et al., 2024) or by training a GNN as a tokenizer (Liu et al., 2023; Luo et al., ).
-Particularly noteworthy are data-driven motif mining approaches that adapt BPE idea to graphs to obtain reusable, multi-scale fragments with strong compression and minimal reliance on manually designed rules or additional trained models (Geng et al., 2023; Shen and Poczos, ; Kong et al., 2022).
-In addition, cross-level tokenization schemes, which combine motifs at different structural levels, also provide complementary design ideas (Ji et al., 2022; Chen et al., ).
+Within a FM paradigm, tokenizer design becomes a key interface for scaling backbones.
+String-based schemes operate on SMILES (Weininger, 1988; Wang et al., 2019) or SELFIES (Krenn et al., 2022).
+Graph-based approaches employ diverse substructure definitions: rule-based tokenizations rely on limited handcrafted chemistry rules (Zhang et al., 2021); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (Li et al., 2023); ring-or-path-based (Wollschläger et al., 2024) and tree-based (Jin et al., 2018) schemes also expose substructures at different granularities; and other works learn tokenization via trainable rules (Sun et al., 2024) or graph neural networks (Liu et al., 2023; Luo et al., 2025).
+Particularly noteworthy are data-driven approaches that adapt BPE idea to graphs to obtain reusable fragments with strong compression and minimal reliance on rules or additional trained models (Geng et al., 2023; Shen and Poczos, 2024; Kong et al., 2022).
+In addition, cross-scale schemes, which combine motifs at different structural levels, also provide complementary design ideas (Ji et al., 2022; Chen et al., 2025).
 3 Method
-Our molecular representation must satisfy two requirements: (1) chemical fidelity, i.e., sensitivity to pharmacophore-level structures and their topology; and (2) autoregressive foundation model compatibility, i.e., a compressed interface naturally aligned with an NTP paradigm (Touvron et al., 2023).
-Motivated by this, we propose Connection-Aware Motif Sequencing (CamS), which segments molecular graphs into reusable frequent motifs and, crucially, serializes the motif graph into multi-scale connection-aware causal sequences compatible with NTP pre-training (Section 3.1).
-Built on CamS, CamS-LLaMA performs large-scale autoregressive pre-training and injects a only lightweight fingerprint prior during downstream fine-tuning (Section 3.2).
-Finally, we analyze the characteristics and advantages of multi-scale CamS autoregressive modeling by comparing it with Graphformer-style modeling (Section 3.3).
+To align chemical fidelity with AR compatibility, we propose Connection-Aware Motif Sequencing (CamS).
+This interface serializes molecular graphs into multi-scale, causal sequences naturally consumable by decoder-only models (Section 3.1).
+Instantiated on this representation, CamS-LLaMA performs standard NTP pre-training, injecting a lightweight fingerprint prior only during downstream fine-tuning (Section 3.2).
+We conclude by theoretically contrasting this CamS-LLaMA framework with Graph Transformers paradigm (Section 3.3).
 3.1 CamS-Tokenizer
-The CamS-tokenizer framework operates in three logical phases to bridge graph topology and scalable NTP.
-(1) CamS Vocabulary Construction: we maintain a robust vocabulary via BPE-style mining and Single-Atom Vocabulary Closure (SAVC) (Section 3.1.1).
-(2) Per-Scale Encoding: molecules are partitioned into motif graphs (BPEGraph) and serialized into single-scale CamS subsequences via scaffold-anchored BFS (Section 3.1.2).
-(3) Cross-Scale Concatenation: subsequences are organized in a fine-to-coarse order to form the final CamS sequence, enabling multiscale autoregressive modeling (Section 3.1.3).
-Full details are in Appendix A.
-3.1.1 Substructure Level: CamS Vocabulary
-BPE-style Mining.
-Following (Geng et al., 2023; Shen and Poczos, ), we first learn an ordered list of merge operations $\mathcal{O}$ from co-occurrence statistics.
-We then apply $\mathcal{O}$ to the corpus to mine the resulting subgraphs.
-This yields the BPEMotif set (frequent subgraphs with connectivity patterns) and the base BPE-SAV (Single-Atom Vocabulary with connectivity patterns) naturally captured during mining.
-Single-Atom Vocabulary Closure (SAVC).
-While generation operates within a closed vocabulary, predictive encoding encounters unseen atom states.
-Standard GraphBPE tokenizers greedily match largest motifs and recursively back off; if a specific atom-connectivity form is absent, it collapses to [UNK].
-This erases chemical details (e.g., element type) and creates Causal Information Breakpoints in the autoregressive stream.
-To prevent this, SAVC constructs the BPE-SAV Back-off via two strategies: (1) enumerating connection-aware tokens for all typical valences, and (2) mapping rare atypical forms of element X to <X_AltForm> rather than <UNK>.
-The final CamS Vocabulary $\Sigma$ is the union of BPEMotif, the complete SAV (BPE-SAV $\cup$ Back-off), and special tokens mapped to unique integers.
-This guarantees that every atom is covered by either a high-level motif or a precise single-atom descriptor (details in Appendix A.1).
-3.1.2 Per-Scale Encoding
-Given a vocabulary scale $s$ (determined by the iteration count of merge operations), encoding a molecule $M$ involves three stages to generate a single-scale CamS subsequence.
-BPEGraph Construction.
-We apply the prefix of merge operations to $M$, inducing a partition of atoms into non-overlapping fragments.
-This forms the BPEGraph $G_{s}=(V_{s},E_{s})$, where nodes $V_{s}\subset\Sigma$ are connection-aware motifs/atoms and $E_{s}$ represents chemical bonds between them.
-Serialization (Intra-scale Order).
-To adapt the graph for decoder-only NTP, we serialize $V_{s}$ into an ordered list via Breadth-First Search (BFS).
-We select the motif with the largest atom count (typically the core scaffold) as the root.
+The CamS-tokenizer framework bridges graph topology and scalable NTP through data-driven fragment compression at the substructure level and hierarchical serialization (both intra- and inter-scale) at the molecule level.
+Full implementation details are provided in Appendix A.
+CamS Vocabulary Construction.
+On the substructure level, we adopt a data-driven strategy grounded in the philosophy that compression implies understanding.
+Unlike rigid manual rules, we apply BPE-style Mining to capture statistically significant chemical semantics.
+Following (Geng et al., 2023; Shen and Poczos, 2024), we first learn an ordered list of merge operations $\mathcal{O}$, sorted by descending frequency based on co-occurrence statistics.
+Since BPE merges are iterative, later operations in $\mathcal{O}$ (lower frequency) build upon earlier ones to combine smaller fragments into larger complex structures.
+Thus, the list order represents a trajectory from fine-grained local substructures to coarse-grained global scaffolds.
+Applying the full list $\mathcal{O}$ yields the BPEMotif set (frequent subgraphs vocabulary) and the base BPE-SAV (single atoms vocabulary) naturally captured during mining.
+Crucially, distinct from prior works, we introduce Single-Atom Vocabulary Closure (SAVC) to address a critical flaw: while generation operates within a closed vocabulary, predictive encoding encounters unseen atom states.
+Standard tokenizers greedily back off to [UNK] when specific atom-connectivity forms are absent, erasing element details and creating “Causal Information Breakpoints” in the AR stream.
+To prevent this, SAVC constructs the BPE-SAV Back-off by: (1) enumerating connection-aware tokens for all typical valences, and (2) mapping rare atypical forms of element X to [X_AltForm] rather than [UNK].
+The final CamS Vocabulary $\Sigma$ is the rigorous union of BPEMotif, the complete SAV (BPE-SAV $\cup$ Back-off), and special tokens.
+This design not only guarantees that every atom is covered by either a high-level motif or a precise single-atom descriptor but also provides a pivotal manipulable dimension for downstream representation: the Motif Scale $s$.
+Formally, we define the Motif Scale $s$ as the effective vocabulary size induced by a specific granularity level.
+Each scale $s$ corresponds to a specific prefix length $k$ of the merge list $\mathcal{O}$.
+A small $s$ (requiring fewer merges) activates only high-frequency motifs, keeping the graph fragmented (Fine Scale); conversely, a large $s$ (using more merges) includes rare, late-stage operations that aggregate these fragments into larger scaffolds (Coarse Scale).
+Per-Scale Encoding.
+On the molecule level, utilizing the constructed CamS Vocabulary $\Sigma$, we can transform raw atom nodes into discrete substructure nodes containing rich connectivity, contextual compression and controllable scale.
+Specifically, given a molecular graph $G_{\text{mol}}$ and a target Motif Scale $s$ (associated with the operation prefix $\mathcal{O}_{\leq k}$), we apply the corresponding $k$ merge operations to $G_{\text{mol}}$.
+This process induces a partition of atoms into non-overlapping fragments, forming the BPEGraph $G_{s}=(V_{s},E_{s})$.
+Here, nodes $V_{s}\subset\Sigma$ denote connection-aware motifs/atoms (ranging from fine fragments to coarse scaffolds depending on $s$) and $E_{s}$ represents the chemical bonds preserving connectivity between them.
+Subsequently, a key challenge remains in how the NTP perceives this high-quality BPEGraph: the definition of causal order.
+While graphs permit various traversals (e.g., random walks (Lovász, 1993)), we establish a deterministic Intra-scale Order.
+We select the motif with the largest atom count (typically the core scaffold) as the root and serialize $V_{s}$ into an ordered list via Scaffold-Rooted BFS.
 This strategy prioritizes global backbone structure before local substitutions (Zhang et al., 2021), establishing a stable Center-to-Periphery causal order.
-ID Extraction. Mapping the ordered nodes to their indices in $\Sigma$ yields the single-scale CamS subsequence $X^{(s)}=(x^{(s)}_{1},\dots,x^{(s)}_{L_{s}})$.
-3.1.3 Cross-Scale Concatenation
-Single-scale tokenization faces a trade-off: coarse scales capture global scaffolds but offer low resolution, while fine scales preserve atomic details but fragment high-level structures (e.g., pharmacophores).
+Finally, we apply ID Extraction to map ordered nodes to their indices in $\Sigma$, yielding the single-scale CamS subsequence $X^{(s)}=(x^{(s)}_{1},\dots,x^{(s)}_{L_{s}})$.
+Cross-Scale Concatenation.
+The manipulable Motif Scale $s$ introduces an inherent resolution trade-off: coarse scales capture global scaffolds but offer low resolution, while fine scales preserve atomic details but fragment high-level pharmacophores.
 CamS resolves this by constructing a multi-scale causal context.
-We tokenize the molecule at multiple scales (e.g., $s\in\{1\mathrm{K},7\mathrm{K},27\mathrm{K},67\mathrm{K}\}$) by slicing $\mathcal{O}$.
-The resulting subsequences are concatenated in a fixed Fine-to-Coarse order (Inter-scale Order) to form the final CamS sequence $\mathbf{X}$:
-$\mathbf{X}=[$ $\,\texttt{[BOS]},X^{(1\mathrm{K})},\texttt{[CONCAT]},X^{(7\mathrm{K})},\texttt{[CONCAT]},$ (1)
-$\,X^{(27\mathrm{K})},\texttt{[CONCAT]},X^{(67\mathrm{K})},\texttt{[EOS]}].$
+We tokenize the molecule at a sequence of $M$ increasing scales $\mathcal{S}=\{s_{1},s_{2},\dots,s_{M}\}$.
+The resulting CamS Subsequences $\{X^{(s)}\mid s\in\mathcal{S}\}$ are concatenated in a fine-to-coarse order (i.e., the Inter-Scale Order) to form the final CamS sequence $\mathbf{X}$:
+$\mathbf{X}=[$ $\,\texttt{[BOS]},X^{(s_{1})},\texttt{[CONCAT]},X^{(s_{2})},\dots,$ (1)
+$\,\texttt{[CONCAT]},X^{(s_{M})},\texttt{[EOS]}].$
 During training, this CamS token stream allows the model to leverage high-resolution local topology (fine scales) as a prefix to condition the prediction of global scaffolds (coarse scales).
-This effectively embeds bottom-up structural composition into the autoregressive objective without architectural modifications (analysis in Section 3.3).
+This effectively embeds bottom-up structural composition into the AR objective without architectural modifications (analysis in Section 3.3).
+Figure 1: The overall framework of CamS.
+(a) CamS-Tokenizer: Transforming molecules into causal sequences via the manipulable Motif Scale $s$.
+First, Per-scale Encoding applies merge operations to construct motif graphs, which are serialized via Scaffold-rooted BFS to establish Intra-Scale Order.
+Subsequently, Cross-Scale Concatenation arranges these views from fine ($s=1\mathrm{K}$) to coarse ($s=67\mathrm{K}$) to establish Inter-Scale Order.
+The resultant CamS sequence serves as a native token stream for a vanilla LLaMA backbone.
+(b) CamS-LLaMA: The model is pre-trained on the resultant CamS Token Stream via standard NTP.
+Pre-trained weights are transferred, and a lightweight fingerprint prior is injected during fine-tuning.
 3.2 CamS-LLaMA
-Figure 1: The overall framework of CamS-LLaMA.
-(a) CamS-Tokenizer: Molecules are transformed into multi-scale causal sequences.
-Through Per-scale Encoding, motif graphs are serialized via Scaffold-rooted BFS (Intra-Scale Order) to establish CamS Subsequence, followed by Cross-Scale Concatenation to arrange subsequences into a final CamS Sequence from fine to coarse granularities (Inter-Scale Order).
-(b) Pre-training Pipeline: The model is pre-trained on CamS Token Lists using a standard LLaMA via NTP, learning to aggregate structural information.
-(c) Fine-tuning Pipeline: During downstream tasks, the pre-trained weights are transferred, and a lightweight fingerprint prior is injected (via projection, prepending in embedding layer and fusion in classifier or regressor) to take part in the prediction of specific properties.
-CamS-LLaMA follows a representation-model co-design principle.
-Section 3.1 maps a molecular graph $G=(V,E)$ into a multi-scale, connection-aware causal sequence $X=(x_{1},\ldots,x_{T})$ (Eq.
-(1)) under vocabularies $\Sigma_{1\mathrm{K}}\subset\Sigma_{7\mathrm{K}}\subset\Sigma_{27\mathrm{K}}\subset\Sigma_{67\mathrm{K}}\triangleq\Sigma$, via BFS-ordered motif traversals $\mathcal{T}_{s}(G)$ and token mapping $x^{(s)}_{\ell}=\iota_{s}(\tau^{(s)}_{\ell})\in\Sigma_{s\mathrm{K}}\subseteq\Sigma$.
-The key point is that this graph-to-causal-sequence interface is intentionally designed to be transparent to a vanilla decoder-only Transformer: it exposes topology through a causal order and a fine-to-coarse hierarchy, so we can train with standard NTP without modifying the LLaMA backbone.
-Based on this interface, CamS-LLaMA follows the foundation model protocol: (1) self-supervised autoregressive pre-training on CamS sequences (figure 1 bottom right), and (2) supervised fine-tuning with a minimal structural prior (fingerprints) injected only at fine-tuning time (figure 1 bottom left).
-3.2.1 Autoregressive Pre-training via NTP
-CamS-LLaMA adopts a standard LLaMA-style decoder (Touvron et al., 2023); in pre-training, every CamS token list is treated as an ordinary token sequence in the shared vocabulary $\Sigma$ and is consumed by the backbone under the usual causal mask, with no graph-specific architectural modification.
-Concretely, each molecule yields five CamS sequence views $\{X_{1},X_{7},X_{27},X_{67},X\}$: four single-scale BFS sequences tokenized with $\Sigma_{1\mathrm{K}}/\Sigma_{7\mathrm{K}}/\Sigma_{27\mathrm{K}}/\Sigma_{67\mathrm{K}}$, plus the concatenated multi-scale sequence $X$ in Eq.
-(1); we treat each view as one NTP training instance (details in Appendix A.3).
-Given any view $X=(x_{1},\ldots,x_{T})$, token embeddings form $\mathbf{H}^{(0)}$, which is transformed by $L$ causal self-attention blocks into $\mathbf{H}^{(L)}$; a tied output projection then yields $p_{\theta}(x_{t+1}\mid x_{\leq t})$.
-Importantly, CamS changes what the model can condition on under the same NTP objective by reshaping the prefix context through fine-to-coarse concatenation (Eq.
-(1)).
-Let $\mathcal{I}_{\mathrm{tok}}\subseteq\{1,\ldots,T-1\}$ be valid prediction positions in $X$ (excluding [BOS], [CONCAT], [EOS]), and $x^{\star}_{t+1}$ be the ground-truth next token.
-The NTP loss is
-$\mathcal{L}_{\mathrm{NTP}}(\theta;X)=-\sum_{t\in\mathcal{I}_{\mathrm{tok}}}\log p_{\theta}(x^{\star}_{t+1}\mid x_{\leq t}).$ (2)
-Because $X$ concatenates BFS sequences from fine to coarse scales (Eq.
-(1)), the conditioning prefix $x_{\leq t}$ has an explicit hierarchy: for $t$ in $X_{67}$ (highest-merge segment), the model conditions on all finer-scale tokens $(X_{1},X_{7},X_{27})$ plus the coarse prefix.
-Thus, coarse-scale prediction errors propagate through attention to both coarse predecessors and the entire fine-scale prefix, providing a fine-to-coarse causal credit assignment under a fully standard decoder.
-On the pre-training corpus $\mathcal{D}_{\mathrm{pre}}$, we minimize
-$\min_{\theta}\;\mathbb{E}_{X\sim\mathcal{D}_{\mathrm{pre}}}\big[\mathcal{L}_{\mathrm{NTP}}(\theta;X)\big],$ (3)
-learning general-purpose molecular representations purely from CamS sequences, without any downstream labels or extra descriptors.
-3.2.2 Fine-tuning with Fingerprint Injection
-Many molecular pre-training methods inject rich handcrafted descriptors early (e.g., during pre-training) (Li et al., 2023).
-In contrast, we keep CamS as the primary representation and inject only one widely-used structural prior–a fingerprint–only during fine-tuning, to make the contribution of CamS representations more attributable and controlled.
-We treat the fingerprint as a virtual global token. For each molecule, a fingerprint vector
-$\mathbf{f}\in\mathbb{R}^{D_{\mathrm{fp}}},$ (4)
-is projected to the model hidden size:
-$\mathbf{f}^{\prime}=W_{\mathrm{fp}}\mathbf{f},\quad W_{\mathrm{fp}}\in\mathbb{R}^{H\times D_{\mathrm{fp}}},$ (5)
-and prepended to token embeddings:
-$\mathbf{H}^{(0)}_{\mathrm{ft}}=[\,\mathbf{f}^{\prime};\mathbf{e}_{1};\dots;\mathbf{e}_{T}\,]\in\mathbb{R}^{(T+1)\times H}.$ (6)
-With a modified causal mask, all positions can attend to this global token. After $L$ layers:
-$\mathbf{H}^{(L)}_{\mathrm{ft}}=[\,\mathbf{h}^{(L)}_{0};\mathbf{h}^{(L)}_{1};\dots;\mathbf{h}^{(L)}_{T}\,$]. (7)
-We use the global token representation $\mathbf{u}^{(i)}=\mathbf{h}^{(L)}_{0}$ for prediction (a decoder-only analogue of [CLS]).
-For classification:
-$\begin{gathered}\hat{\mathbf{y}}^{(i)}=\mathrm{softmax}\!\big(W_{2}\,\sigma(W_{1}\mathbf{u}^{(i)})\big),\\ \mathcal{L}_{\mathrm{cls}}(\theta,\phi)=-\frac{1}{B}\sum_{i=1}^{B}\log\hat{\mathbf{y}}^{(i)}_{\,y^{(i)}}.\end{gathered}$ (8)
-For regression:
-$\begin{gathered}\hat{y}^{(i)}=W_{2}\,\sigma(W_{1}\mathbf{u}^{(i)}),\\ \mathcal{L}_{\mathrm{reg}}(\theta,\phi)=\frac{1}{B}\sum_{i=1}^{B}\big(\hat{y}^{(i)}-y^{(i)}\big)^{2}.\end{gathered}$ (9)
-Fine-tuning minimizes
-$\min_{\theta,\phi}\;\mathbb{E}_{(X,\mathbf{f},y)\sim\mathcal{D}_{\mathrm{task}}}\big[\mathcal{L}_{\mathrm{cls}/\mathrm{reg}}(\theta,\phi;X,\mathbf{f},y)\big].$ (10)
-This keeps pre-training purely sequence/structure-driven, while using fingerprints as a lightweight and controlled prior only when optimizing for downstream targets.
-3.3 CamS Representation vs. Graph Transformer
+Following a representation-model co-design principle, CamS-LLaMA leverages the graph-to-sequence interface established in Section 3.1 to enable standard AR modeling on molecular graphs.
+As illustrated in Figure 1(a), the multi-scale CamS sequence acts as a transparent bridge, exposing topology to a vanilla LLaMA backbone without architectural modifications.
+This facilitates a scalable two-stage FM pipeline (Figure 1(b)): (1) self-supervised pre-training via NTP to capture structural generative laws, and (2) property-guided fine-tuning with a controlled fingerprint prior.
+Autoregressive Pre-training via NTP.
+CamS-LLaMA adopts a standard LLaMA-style decoder (Touvron et al., 2023).
+In pre-training, CamS token lists are treated as ordinary sequences in the shared vocabulary $\Sigma$ and consumed under the usual causal mask.
+Concretely, to maximize structural learning signals via data augmentation, each molecule yields five sequence views (four single-scale + one multi-scale); we treat each view as an independent NTP training instance (details in Appendix A.3).
+Conditioned on the prefix context reshaped by fine-to-coarse concatenation (Eq.
+(1)), the model learns to predict the next token $x^{\star}_{t+1}$.
+The NTP loss is minimized on the pre-training corpus $\mathcal{D}_{\mathrm{pre}}$:
+$\mathcal{L}_{\mathrm{NTP}}(\theta)=-\mathbb{E}_{X\sim\mathcal{D}_{\mathrm{pre}}}\sum_{t\in\mathcal{I}_{\mathrm{tok}}}\log p_{\theta}(x^{\star}_{t+1}\mid x_{\leq t}).$ (2)
+Crucially, because coarse-scale tokens appear after fine-scale ones in $X$, prediction errors propagate gradients through the entire fine-scale prefix, effectively providing a fine-to-coarse causal credit assignment that embeds bottom-up compositional logic into the standard decoder.
+Fine-tuning with Fingerprint Injection.
+Many methods inject rich handcrafted descriptors early during pre-training (Li et al., 2023).
+In contrast, we keep CamS as the primary representation and inject a lightweight fingerprint prior only during fine-tuning via a dual-path strategy (Early Injection + Late Fusion).
+Given a molecule’s fingerprint $\mathbf{f}\in\mathbb{R}^{D_{\mathrm{fp}}}$, we project it to $\mathbf{f}^{\prime}=W_{\mathrm{fp}}\mathbf{f}$ using a learnable projection $W_{\mathrm{fp}}\in\mathbb{R}^{H\times D_{\mathrm{fp}}}$.
+First, via Early Injection, $\mathbf{f}^{\prime}$ is prepended to the input embeddings as a global prompt:
+$\mathbf{H}^{(0)}_{\mathrm{ft}}=[\,\mathbf{f}^{\prime};\mathbf{e}_{\text{BOS}};\dots;\mathbf{e}_{\text{EOS}}\,]\in\mathbb{R}^{(T+2)\times H}.$ (3)
+After $L$ layers, we extract the representation of the [EOS] token, $\mathbf{h}_{\text{EOS}}^{(L)}$, which aggregates the AR structural context.
+Second, to allow the backbone to focus on deep implicit reasoning (by offloading explicit feature extraction to the direct path), we perform Late Fusion by concatenating $\mathbf{h}_{\text{EOS}}^{(L)}$ with the original projected fingerprint $\mathbf{f}^{\prime}$ to form the final representation $\mathbf{u}$:
+$\mathbf{u}=[\,\mathbf{h}_{\text{EOS}}^{(L)}\,;\,\mathbf{f}^{\prime}\,]\in\mathbb{R}^{2H}.$ (4)
+This shortcut not only preserves chemical fidelity but also serves as a stabilizer for fine-grained decision making (as analyzed in Section 4.4), ensuring predictions remain grounded even when deep features fluctuate.
+This fused vector is fed into task-specific heads.
+For classification ($\hat{\mathbf{y}}$) and regression ($\hat{y}$), the objectives are:
+$\mathcal{L}_{\mathrm{cls}}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\big(\mathrm{softmax}(W_{2}\,\sigma(W_{1}\mathbf{u}))\big)_{y^{(i)}},$ (5)
+$\mathcal{L}_{\mathrm{reg}}$ $=\frac{1}{B}\sum_{i=1}^{B}\big(W_{2}\,\sigma(W_{1}\mathbf{u})-y^{(i)}\big)^{2}.$
+This strategy balances the pre-trained structural knowledge with the explicit fingerprint prior, optimizing $\mathbb{E}_{\mathcal{D}_{\mathrm{task}}}[\mathcal{L}_{\mathrm{task}}]$.
+3.3 CamS LLaMA vs. Graph Transformer
 We analyze CamS-LLaMA and Graph Transformer-style models (e.g., KPGT) within a unified framework of Token-Level Graph-Structured Deep Feature Construction.
-We show that CamS yields (1) a denser and multi-view direct supervision signal and (2) a hierarchical inductive bias that are mathematically distinct from standard masked-node prediction (MNP).
-Detailed derivations and supporting discussions are provided in Appendix B.
-3.3.1 Unified Framework.
-Let $G=(V,E)$ be a molecular graph. Both paradigms construct a token graph $\mathcal{G}_{\mathrm{tok}}=(\mathcal{V}_{\mathrm{tok}},\mathcal{E}_{\mathrm{tok}})$ and perform $L$ layers of propagation:
+We show that CamS yields (1) a denser multi-view direct supervision signal and (2) a hierarchical inductive bias that are mathematically distinct from standard masked-node prediction (MNP).
+Detailed derivations are in Appendix B.
+Unified Framework.
+Let $G=(V,E)$ be a molecular graph.
+Both paradigms construct a token graph $\mathcal{G}_{\mathrm{tok}}=(\mathcal{V}_{\mathrm{tok}},\mathcal{E}_{\mathrm{tok}})$ and perform $L$ layers of propagation:
 $\begin{split}\mathbf{H}^{(l+1)}&=\mathrm{Agg}\big(\mathbf{A}^{(l)}\,\mathrm{Val}(\mathbf{H}^{(l)})\big),\\ \mathbf{A}^{(l)}_{ij}&\propto\exp\left(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d}}+\psi_{ij}\right).\end{split}$ (6)
-where $\mathbf{H}^{(l)}$ are token embeddings, $\mathbf{q}_{i},\mathbf{k}_{j}$ are projections of $\mathbf{H}^{(l)}$, and $\psi_{ij}$ encodes the structural bias or causal mask.
-Graphformer (Flat/Static Bias).
-$\mathcal{V}_{\text{tok}}$ corresponds to single-scale atoms or triplets.
-$\psi_{ij}$ encodes static structural biases (e.g., shortest-path-distance bias injected into attention) independent of token content, yielding an isotropic receptive field at a given graph distance (Ying et al., 2021).
-CamS-LLaMA (Hierarchical/Causal Flow).
-$\mathcal{V}_{\text{tok}}$ comprises multi-scale motifs $\{v^{(s)}\}$, where the same molecule is re-represented across scales.
-$\psi_{ij}$ is determined by the causal mask $\mathbf{M}_{\text{causal}}$, enforcing a directed flow where coarse-scale tokens attend to fine-scale predecessors.
-This turns cross-scale aggregation into a learned, anisotropic connectivity along the sequence.
-3.3.2 Information-Theoretic Context Analysis.
+where $\mathbf{H}^{(l)}$ are token embeddings and $\psi_{ij}$ encodes structural bias.
+For Graph Transformers (Flat/Static Bias), $\mathcal{V}_{\mathrm{tok}}$ corresponds to single-scale atoms or triplets.
+$\psi_{ij}$ encodes static biases (e.g., shortest-path distance) independent of token content, yielding an isotropic receptive field (Ying et al., 2021).
+For CamS-LLaMA (Hierarchical/Causal Flow), $\mathcal{V}_{\mathrm{tok}}$ comprises multi-scale connection-aware motifs $\{v^{(s)}\}$.
+$\psi_{ij}$ is determined by the causal mask $\mathbf{M}_{\mathrm{causal}}$, enforcing a directed flow where coarse-scale tokens attend to fine-scale predecessors (via Inter-Scale Order).
+This turns cross-scale aggregation into a learned, anisotropic connectivity.
+Information-Theoretic Context Analysis.
 We contrast NTP with MNP.
-Let $x_{t}$ be a target token and let $Z_{t}$ denote its uncorrupted evidence set (e.g., the causal history for CamS coarse-scale tokens).
-MNP observes a stochastically corrupted context $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ via random masking.
+Let $x_{t}$ be a target token and $Z_{t}$ its uncorrupted evidence set.
+MNP observes a stochastically corrupted context $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$.
 Proposition 3.1 (Context Information Inequality).
-For any non-trivial masking channel $\mathcal{M}$, we have
-$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}).$ (12)
+For any non-trivial masking channel $\mathcal{M}$, we have $I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t})$.
 Sketch.
@@
 In graph MNP, this loss of information is exacerbated by evidence-pattern uncertainty: the most relevant local-neighborhood evidence for a masked token may be co-masked, making the conditional distribution unstable across masking patterns.
-Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets / local subgraphs) or designing structure-guided masking, and by injecting global knowledge tokens (Li et al., 2023; Rong et al., 2020; Liu et al., 2024).
+Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets or local subgraphs), designing structure-guided masking, or injecting global knowledge tokens (Li et al., 2023; Rong et al., 2020; Liu et al., 2024).
 In contrast, CamS NTP for coarse-scale tokens conditions on an uncorrupted fine-to-coarse history, avoiding such stochastic evidence loss.
 This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (Yang et al., 2019).
-3.3.3 Gradient Flow Density.
-We define the Direct Supervision Density (SD) as the expected fraction of tokens serving as regression targets per update.
-For a multi-scale sequence with total length $T_{\text{all}}=\sum_{s=1}^{S}T_{s}$ and mask ratio $\rho$:
-$\text{SD}_{\text{NTP}}\approx 1\quad\gg\quad\text{SD}_{\text{MNP}}=\rho.$ (13)
-This trade-off (more predicted tokens vs.
-more corruption) is central to masked modeling (Wettig et al., 2023).
-Empirically, KPGT finds the best masking ratio can be as high as $\rho=0.5$ when using triplet-style tokens and global knowledge nodes to stabilize context (Li et al., 2023).
-Meanwhile, dense token-level learning signals are known to substantially improve sample efficiency compared to sparse masked-only supervision (Clark et al., 2020).
-CamS further strengthens this effect by providing multi-view supervision: the same molecule contributes token-level targets repeatedly across scales.
-3.3.4 Hierarchical Inductive Bias.
-Graphformers rely on implicit depth and static structural biases to propagate information.
-In contrast, CamS injects an explicit structural hierarchy via the multi-scale sequence $X=[X_{1},\dots,X_{T_{\text{all}}}]$, and the causal mask ensures coarse-scale predictions have deterministic access to fine-scale details.
-We hypothesize this is particularly beneficial for activity-cliff regimes, where small local motif edits trigger global property shifts, as it biases the model toward bottom-up composition rules.
+Gradient Flow Density and Hierarchical Inductive Bias.
+We define the Direct Supervision Density (SD) as the expected fraction of tokens serving as targets per update.
+With mask ratio $\rho$, we have $\text{SD}_{\text{NTP}}\approx 1\gg\text{SD}_{\text{MNP}}=\rho$.
+Masked models face a dilemma: increasing $\rho$ improves density but exacerbates context corruption (Wettig et al., 2023); for instance, KPGT relies on global nodes to stabilize optimization at $\rho=0.5$ (Li et al., 2023).
+In contrast, CamS achieves maximal density on uncorrupted contexts and the same structure contributes gradients repeatedly across scales, yielding substantially higher sample efficiency than sparse masked supervision (Clark et al., 2020).
+From another perspective, Graph Transformers rely on implicit depth to propagate information.
+In contrast, CamS injects an explicit structural hierarchy via the multi-scale sequence $\mathbf{X}$, where the causal mask ensures coarse-scale predictions have deterministic access to fine-scale details.
+We hypothesize this bias is particularly beneficial for stabilizing predictions in activity-cliff regimes (as supported by ablation study in Section 4.4).
 4 Experiment
 4.1 Experimental Setup
-We decouple the pipeline into three stages, using the ChEMBL-34 Dataset for tokenizer training, the Enamine675M Dataset for NTP pre-training, and the MoleculeNet/MoleculeACE Benchmark for downstream fine-tuning and evaluation.
-CamS-Tokenizer Training.
-CamS-tokenizer training is expensive because replaying $\mathcal{O}$ to materialize $\mathcal{V}$ repeatedly invokes RDKit’s subgraph-isomorphism matching (the dominant bottleneck) (Landrum et al., 2021; Ehrlich and Rarey, 2011) to extract motifs and attachment patterns, which is difficult to accelerate.
-We therefore train the tokenizer once on ChEMBL-34 (Gaulton et al., 2017), a moderately sized yet chemically diverse corpus, at the largest scale (67K).
-Concretely, we mine an ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K})$ with K=685K{=}685 merges, apply it to obtain $\mathcal{V}$, and then perform SAVC to ensure single-atom coverage for common elements.
-Smaller scales are obtained without retraining by prefix slicing: each scale uses a prefix $\mathcal{O}_{\leq k}$ together with its induced vocabulary subset (i.e., motifs mined by applying $\mathcal{O}_{\leq k}$), where 1K uses k=0k{=}0, 7K/27K use k=62/210k{=}62/210, and 67K uses k=685k{=}685.
-Implementation details and full procedures are provided in Appendix A.1.
-Pre-training Dataset and Configuration.
-We instantiate a 16-layer LLaMA-style decoder (hidden size 720, 8 attention heads) using HuggingFace Transformers.
-For self-supervised pre-training, we use Enamine675M (675M molecules) (Shivanyuk et al., 2007) and optimize NTP (Section 3.2.1).
-For each molecule, we construct five CamS training sequences: four single-scale BFS sequences tokenized with 1K/7K/27K/67K vocabularies, plus the concatenated multi-scale sequence in Eq.
-(1).
-We treat each sequence as one training instance and uniformly shuffle all instances, resulting in roughly $5\times$ as many sequences as molecules.
-All optimization details (schedule, batch size, steps) are summarized in Appendix C.1.
+CamS-Tokenizer Configuration.
+We train the CamS-Tokenizer on ChEMBL-34 (Gaulton et al., 2017) to balance diversity with feasibility, as vocabulary materialization involves expensive subgraph-isomorphism matching (Landrum et al., 2021; Ehrlich and Rarey, 2011) that prohibits direct training on billion-scale corpora.
+We learn the merge list $\mathcal{O}$ once with K=685K{=}685 operations to materialize the full CamS Vocabulary ($67\mathrm{K}$).
+To instantiate the multi-scale context (M=4M{=}4), we slice $\mathcal{O}$ at prefix indices $k\in\{0,62,210,685\}$, inducing Motif Scales $s\in\{1\mathrm{K},7\mathrm{K},27\mathrm{K},67\mathrm{K}\}$.
+These scales are selected to span the granularity spectrum: the $1\mathrm{K}$ scale serves as the atomic baseline (fine limit), the $67\mathrm{K}$ scale represents the experiment-maximal coarse limit, while $7\mathrm{K}$ and $27\mathrm{K}$ provide necessary intermediate resolutions to bridge the gap.
+Full details are in Appendix A.1.
+Pre-training Implementation.
+We instantiate a 16-layer LLaMA decoder (hidden size 720, 8 heads; $\sim$100M parameters, comparable to KPGT).
+Pre-training uses Enamine675M (675M molecules) (Shivanyuk et al., 2007).
+Following Section 3.2, we generate five sequence views per molecule, treating them as independent instances and uniformly shuffling the resulting dataset.
+We optimize the NTP objective (Eq.
+2) using hyperparameters summarized in Appendix C.1.
 Downstream Benchmarks and Baselines.
-We evaluate on MoleculeNet (11 tasks; broad endpoints that probe general-purpose molecular property prediction) (Wu et al., 2018) and MoleculeACE (30 activity-cliff tasks; stress-testing fine-grained structural sensitivity under subtle edits that induce large activity shifts) (Van Tilborg et al., 2022).
-To enable an apples-to-apples comparison against a recent strong pre-training baseline, we adopt the data splits and evaluation protocol used in KPGT (Li et al., 2023).
-On MoleculeNet, following KPGT’s setup, we compare primarily against self-supervised pre-training baselines to isolate the effect of (1) representation/tokenization and (2) the pre-training objective.
-On MoleculeACE, we report the benchmark’s comprehensive ML/DL baselines, since activity cliffs emphasize discriminating small structural changes where strong feature-based methods remain competitive.
-Unless stated otherwise, CamS-LLaMA is kept at $\sim$100M parameters (comparable to KPGT); fingerprints are injected only during fine-tuning as a minimal prior, while pre-training is purely sequence-based.
-4.2 Main Results
-We summarize the main comparisons in Table 1 (MoleculeNet) and Table 2 (MoleculeACE).
+We evaluate on MoleculeNet (11 general property tasks) (Wu et al., 2018) and MoleculeACE (30 activity-cliff tasks) (Van Tilborg et al., 2022), strictly adhering to the data splits and protocols established by the SOTA baseline KPGT (Li et al., 2023) (details in Appendix C.2).
+We benchmark against KPGT and the baselines adopted in KPGT’s evaluation.
+Crucially, distinct from KPGT’s heavy reliance on rich descriptor at all stages, we inject lightweight fingerprints only during fine-tuning, maintaining a purely structural pre-training.
+On MoleculeNet, we include comprehensive graph-based self-supervised baselines, such as MolFormer (MolF) (Wu et al., 2023), ContextPred (Ctxt) (Hu et al., 2020), GROVER (GROV) (Rong et al., 2020), JOAO (You et al., 2020), GEM (Fang et al., 2022), GraphMAE (GMAE) (Hou et al., 2022), and MoleBERT (MBRT) (Xia et al., 2023).
+Additionally, we benchmark against the SMILES-NTP FM NatureLM (Xia et al., 2025) (comparison in Appendix D).
+On MoleculeACE, we additionally include strong classical machine learning baselines (SVM, RF, GBM (Van Tilborg et al., 2022) with ECFP (E) (Rogers and Hahn, 2010) or MACCS (M) (Durant et al., 2002) descriptors), which are known to outperform deep learning methods on the activity-cliff stress test.
+4.2 Results on Downstream Tasks
 General Property Prediction (MoleculeNet).
-Table 1 shows that CamS-LLaMA achieves overall state-of-the-art on MoleculeNet under a comparable model scale to KPGT ($\sim$100M): it slightly improves the average AUROC in classification (0.845 vs.
-0.843) and the average RMSE in regression (1.172 vs.
+Table 1 demonstrates that CamS-LLaMA establishes a new SOTA on MoleculeNet.
+While maintaining a comparable model scale to KPGT ($\sim$100M) with fewer priors, CamS-LLaMA surpasses it in both classification (AVG-AUROC: 0.845 vs.
+0.843) and regression (AVG-RMSE: 1.172 vs.
 1.175).
-We also win 6/11 individual tasks (5/8 classification and 1/3 regression), and improve over the previous best self-supervised learning baselines by a clear margin (AVG-AUROC: 0.845 vs.
-0.825; AVG-RMSE: 1.172 vs.
-1.272).
-Notably, this is achieved with weaker prior injection: unlike KPGT, which explicitly feeds rich handcrafted descriptors, we keep pre-training purely sequence-based and inject only a lightweight fingerprint prior during downstream fine-tuning.
-Full per-baseline results and extended tables are provided in Appendix D.
+Crucially, these results underscore superior scalability and efficiency: unlike KPGT’s descriptor-imposed bottleneck on data scaling or NatureLM’s reliance on massive parameters ($\sim$56B, see Appendix D), CamS leverages a purely sequence-driven pre-training to efficiently scale to 675M pure molecular data (discussion in Appendix E) and achieve SOTA with a compact backbone.
+By winning 6 out of 11 individual tasks, CamS-LLaMA proves that properly serialized causal modeling captures molecular properties as effectively as descriptor-enhanced graph methods, without their reliance on external domain priors.
+Full per-baseline results are provided in Appendix F.
 […]
 Unified Takeaway.
-Taken together, Tables 12 indicate that CamS’s three-stage design (motif mining with single-atom completion $\rightarrow$ causal serialization $\rightarrow$ fine-to-coarse multi-scale concatenation) yields two complementary benefits: (1) general-purpose representation strength, by organizing functional-group semantics, attachment patterns, and topology into an NTP-friendly causal context; and (2) activity-cliff-oriented discriminability, by exposing cliff-driving edits through motif-level abstraction, position-sensitive causal contexts, and redundant multi-scale views.
-This motivates a targeted interpretability analysis below, where we test whether the model indeed allocates more attention to cliff-differential tokens than to shared scaffolds, and how this preference varies across scale regions.
+The results across Tables 1 and 2 reveal a distinct advantage of the CamS framework: it combines general-purpose robustness with fine-grained structural discriminability.
+The three-stage design—(1) motif mining, (2) causal serialization, and (3) multi-scale concatenation—allows the model to encode both global scaffolds and local variations effectively.
+By doing so, CamS-LLaMA achieves what KPGT attempts via explicit descriptors: it forces the model to attend to chemically significant substructures, but does so intrinsically through the vocabulary and causal objective rather than extrinsically through handcrafted features.
+This motivates the targeted interpretability analysis in Section 4.3, where we verify whether the model indeed allocates sharper attention to cliff-differential tokens compared to baselines.
 […]
-Task Previous Best Baselines KPGT CamS-LLaMA
-Method Score (Ours)
-Classification Tasks (AUROC $\uparrow$)
-bace GEM 0.857(0.016) 0.855(0.014) 0.870(0.013)
-bbbp GEM 0.895(0.024) 0.908(0.012) 0.942(0.015)
-clintox GEM 0.905(0.027) 0.946(0.026) 0.935(0.017)
-estrogen GEM 0.894(0.048) 0.906(0.034) 0.917(0.050)
-metstab GROV 0.876(0.046) 0.889(0.057) 0.891(0.059)
-sider JOAO 0.640(0.012) 0.649(0.011) 0.655(0.016)
-toxcast GEM 0.733(0.020) 0.745(0.003) 0.724(0.008)
-[…]
-AVG GEM 0.825 0.843 0.845
-Regression Tasks (RMSE $\downarrow$)
-esol GEM 0.803(0.051) 0.804(0.102) 0.761(0.046)
-freesolv MolF 2.322(0.613) 2.121(1.025) 2.110(0.959)
-lipo GROV 0.625(0.007) 0.600(0.012) 0.645(0.023)
-AVG MolF 1.272 1.175 1.172
-Table 2: Performance on MoleculeACE (RMSE).
+Task Prev. Best KPGT Ours Task Prev. Best KPGT Ours
+Method Score Method Score
+Classification Tasks (AUROC $\uparrow$) Classification Tasks Cont.
+BACE GEM 0.857(0.016) 0.855(0.014) 0.870(0.013) Tox21 Ctxt 0.840(0.028) 0.848(0.017) 0.827(0.028)
+BBBP GEM 0.895(0.024) 0.908(0.012) 0.942(0.015) AVG (Cls) GEM 0.825 0.843 0.845
+ClinTox GEM 0.905(0.027) 0.946(0.026) 0.935(0.017) Regression Tasks (RMSE $\downarrow$)
+Estrogen GEM 0.894(0.048) 0.906(0.034) 0.917(0.050) ESOL GEM 0.803(0.051) 0.804(0.102) 0.761(0.046)
+Metstab GROV 0.876(0.046) 0.889(0.057) 0.891(0.059) FreeSolv MolF 2.322(0.613) 2.121(1.025) 2.110(0.959)
+SIDER JOAO 0.640(0.012) 0.649(0.011) 0.655(0.016) Lipo GROV 0.625(0.007) 0.600(0.012) 0.645(0.023)
+ToxCast GEM 0.733(0.020) 0.745(0.003) 0.724(0.008) AVG (Reg) MolF 1.272 1.175 1.172
+Table 2: Performance on MoleculeACE (RMSE $\downarrow$).
 AVG: Average RMSE across all 30 tasks.
-Ranking: Bold (1st), Underline (2nd), Italic (3rd).
-Metric: $\downarrow$ denotes lower is better.
-Column: Previous Best Baselines denotes the best-performing method excluding KPGT; Type: ML (Machine Learning), DL (Deep Learning).
-Reference & Abbreviations: GMAE: GraphMAE (Hou et al., 2022); MBRT: MoleBERT (Xia et al., 2023); GROV: GROVER (Rong et al., 2020); SVM/RF/GBM: Support Vector Machine / Random Forest / Gradient Boosting Machine (Van Tilborg et al., 2022).
-Descriptors: E: ECFP (Rogers and Hahn, 2010); M: MACCS (Durant et al., 2002).
-Task Previous Best Baselines KPGT CamS-LLaMA
-Method Type Score (Ours)
-Regression Tasks (RMSE $\downarrow$)
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Bold: 1st; Underline: 2nd; Italic 3rd.
+Prev.
+Best: Best baseline excluding KPGT.
+Type: ML denotes traditional machine learning; DL denotes deep learning.
+Task Prev. Best KPGT Ours Task Prev. Best KPGT Ours
+Method Type Score Method Type Score
 […]
@@
 Interpretability Setup and Metric.
-To explain the activity-cliff advantage in Table 2, we test whether CamS-LLaMA preferentially attends to cliff-driving structural differences.
-Following the MoleculeACE protocol, we construct activity-cliff pairs on the test sets and identify differential vs.
-shared atom-level activity-cliff point; we then map them to tokens within each scale region of the concatenated CamS sequence (Appendix E).
-Given the final-layer attention distribution, we compute mean attention on differential tokens (MDTA) and shared tokens (MSTA) within a region and report the relative preference $\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100$, evaluated for each scale region and for the full concatenated sequence.
-We report the full attention metrics (MDTA/MSTA/Rel-DTAP) in Appendix D.
+To mechanistically explain the activity-cliff superiority observed in Table 2, we investigate whether the CamS-LLaMA attention mechanism inherently prioritizes cliff-driving structural variations.
+Aligned with the CamS-Tokenizer framework (Section 3.1), we map atom-level differences in activity-cliff pairs to tokens within specific Motif Scales of the concatenated CamS Sequence $\mathbf{X}$.
+Specifically, for each cliff pair, we identify differential versus shared atoms and project these labels onto the CamS Sequence $\mathbf{X}$.
+Using the final-layer attention distribution, we compute the Mean Attention on Differential Tokens (MDTA) and Shared Tokens (MSTA) within each scale region $s$.
+We define the Relative Differential-Token Attention Preference (Rel-DTAP) as:
+$\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100\%,$ (7)
+averaged over all test pairs.
+A positive value indicates that the model allocates disproportionately higher attention to the subtle structural edits driving the activity cliff.
+Full implementation details are in Appendix G.
 Findings and Implications.
-Table 3 shows three patterns.
-(1) Overall preference is positive: on the concatenated sequence, Rel-DTAP is consistently >0>0 (14.79% w/o fingerprint; 14.23% w/ fingerprint), indicating that the model, on average, prioritizes cliff-driving differences over shared scaffolds.
-(2) Preference is scale-dependent: coarse regions (27K/67K) exhibit much stronger preference (24.33%/25.87% w/o fingerprint) than the fewest-merge region (1K, 8.08%), suggesting that coarse motifs provide a more salient interface for highlighting differential substructures, while fine scales contribute weaker but complementary local evidence.
-(3) Fingerprint acts as a stabilizer rather than a uniform booster: although fingerprint slightly reduces the overall preference, it corrects the anomalous 7K behavior (from -0.18% to 2.10%), consistent with fingerprint providing global guidance that resolves ambiguity at mid-scale tokenizations.
-Together, these results support the cliff-specific advantage of CamS: multi-scale causal serialization does not merely improve accuracy, but also steers attention toward structurally meaningful differences that drive activity cliffs.
+Table 3 reveals three patterns that corroborate our design principles: (1) Intrinsic Structural Discriminability: On the full sequence, Rel-DTAP is consistently positive ($\sim$14%), confirming that even without explicit guidance, the NTP-driven causal objective naturally steers attention toward cliff-driving motifs.
+(2) Scale-Dependent Hierarchy: Coarse scales (27K/67K) exhibit significantly stronger preference ($\sim$25%) than the fine-grained baseline (1K, $\sim$8%).
+This validates our hypothesis in Section 3.1 that coarse motifs act as high-level ”semantic anchors,” making structural edits more salient compared to the fragmented 1K view.
+(3) Fingerprint as a Stabilizer: Fingerprint injection (Section 3.2) does not uniformly boost attention but acts as a corrective stabilizer.
+It specifically rectifies the anomalous attention pattern at the intermediate 7K scale (from -0.18% to 2.10%), resolving ambiguity where motif granularity may be suboptimal.
+Collectively, these findings provide a mechanistic basis for CamS’s dominance on MoleculeACE: the multi-scale architecture explicitly exposes subtle edits as distinct multi-granular tokens, preventing them from being obscured by the local averaging mechanisms common in standard graph encoders.
 Table 3: Relative Differential-Token Attention Preference (Rel-DTAP).
-Within each scale region of the concatenated multi-scale sequence, we compute per-molecule $\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100\%$, then average over all molecules from activity-cliff pairs (anchor and partner counted separately).
-Positive values mean higher per-token attention on differential tokens than on shared tokens.
-Scale region Without FP With FP
-1K 8.08% 4.72%
+We report the average Rel-DTAP across all activity-cliff test pairs.
+A positive value indicates that the model attends more intensively to differential tokens.
+Scale Region ($s$): Indicates the model’s attention to regions in CamS Sequence corresponding to specific Motif Scales.With FP / Without FP: Indicates whether the fingerprint was injected.
+Scale Region ($s$) Without FP With FP
+1K (Fine) 8.08% 4.72%
 7K -0.18% 2.10%
 27K 24.33% 23.60%
-67K 25.87% 24.00%
-CONCAT (all scales) 14.79% 14.23%
+67K (Coarse) 25.87% 24.00%
+CONCAT (All) 14.79% 14.23%
+[…]
+[…]
 4.4 Ablation Study
 Ablation setup.
-To isolate the contribution of each core component, we evaluate three variants against the full model (Table 4): (1) w/o FP removes fingerprint injection during fine-tuning; (2) Vocab-1K Only and (3) Vocab-67K Only remove multi-scale concatenation by restricting both pre-training and fine-tuning to a single scale.
-For fair comparison, single-scale variants keep the pre-training budget matched to the full model (details in Appendix C.1).
-Multi-scale concatenation is indispensable.
-The full model consistently outperforms both single-scale variants across all benchmarks (e.g., MolNet-Cls 0.845 vs.
-0.833/0.818; MolNet-Reg 1.172 vs.
-1.215/1.329; MolACE 0.624 vs.
-0.641/0.649), confirming that multi-scale concatenation is not a minor engineering choice but the key mechanism that lets coarse-scale reasoning build on fine-scale evidence within one causal context.
-This directly supports our central claim: multi-scale causal serialization is the representation-level ingredient that makes graph structure learnable by a vanilla decoder-only NTP backbone.
-Why 1K-only can beat 67K-only under a fixed NTP budget.
-Under single-scale training, Vocab-67K Only is consistently weaker than Vocab-1K Only, especially on regression (1.329 vs.
-1.215 RMSE on MolNet-Reg).
-We interpret this as an over-compression effect: at 67K, sequences for small molecules become very short and each token packs high compositional complexity, which reduces the number of NTP supervision steps per molecule and makes causal dependency learning harder under limited model capacity and training steps.
-In contrast, 1K produces longer sequences that provide denser autoregressive supervision and more stable causal credit assignment, albeit at the cost of fragmenting scaffold-level semantics.
-Concat resolves this trade-off by retaining the dense supervision of fine scales while reintroducing scaffold-level abstraction at coarse scales, without committing to a single “best” vocabulary size.
-[…]
-Table 4: Ablation study and ranking comparison.
-Scores represent the average performance across all tasks in the respective benchmarks from Tables 1 and 2.
-Rank: Global ranking of the specific variant when pooled with the complete set of baseline models.
-Previous Best: The highest-performing baseline excluding KPGT.
-Abbreviations: MolF: MolFormer
-Method MolNet-Cls ($\uparrow$) MolNet-Reg ($\downarrow$) MolACE ($\downarrow$)
-Score Rank RMSE Rank RMSE Rank
-CamS-LLaMA (Full) 0.845 1 1.172 1 0.624 1
+We evaluate three variants against the full model (summary in Table 4, full breakdown in Appendix H) to isolate core components: (1) w/o FP completely removes fingerprint injection; (2) $1\mathrm{K}$ Only and (3) $67\mathrm{K}$ Only use single fixed scales.
+Single-scale variants have pre-training and fine-tuning budgets matched to the full model.
+Indispensability of Multi-Scale Context.
+The full model consistently outperforms single-scale variants (e.g., MolACE 0.624 vs.
+0.649), confirming that Cross-Scale Concatenation (Section 3.1) is fundamental.
+By enabling coarse reasoning to ground on fine evidence, this design validates our central thesis: multi-scale causal serialization is the key representation-level innovation that unlocks vanilla NTP for molecular graphs modeling.
+Pitfall of Coarse-Scale Over-Compression.
+Single-scale results show that coarse modeling ($s=67\mathrm{K}$) consistently lags behind fine-grained modeling ($s=1\mathrm{K}$).
+We attribute this to supervision sparsity: over-compressed sequences provide insufficient AR steps for effective learning.
+The full model circumvents this by concatenation, successfully fusing the dense supervision of fine scales with the coarse-scale scaffold-level abstraction, without committing to a specific “best” resolution.
+Fingerprint as the Maximum-Scale Global Token.
+The pure-sequence model (w/o FP) proves remarkably robust, particularly in classification where it statistically rivals KPGT (0.838 vs.
+0.843).
+While KPGT resorts to heavy priors to patch MNP’s limitations (e.g., over-smoothing), our use of the fingerprint is fundamentally different: we strategically integrate it as the maximum-scale global token (equivalent to Motif Scale $s=\infty$) to complete the topological hierarchy.
+Acting as a stabilizer (Section 4.3), it plays a significantly larger role in regression tasks, effectively completing the final link of our multi-scale structural context.
+Thus, the full model’s superiority stems from a unified causal formulation: $Global+(Fine\to Coarse)$, where the fingerprint serves as the holistic structural anchor driving the fine-to-coarse reasoning.
+Table 4: Ablation study. Scores represent the average performance across all tasks in the respective benchmarks from Tables 1 and 2. R: Global ranking in complete baselines.
+Method MolNet-Cls MolNet-Reg MolACE
+AUROC ($\uparrow$) R RMSE ($\downarrow$) R RMSE ($\downarrow$) R
+CamS-LLaMA 0.845 1 1.172 1 0.624 1
 – w/o FP 0.838 3 1.195 3 0.650 5
-– Vocab-1K Only 0.833 4 1.215 4 0.641 3
-– Vocab-67K Only 0.818 6 1.329 7 0.649 4
+– 1K Only 0.833 4 1.215 4 0.641 3
+– 67K Only 0.818 6 1.329 7 0.649 4
 KPGT 0.843 2 1.175 2 0.633 2
@@
 5 Conclusion
-We introduced CamS, a graph-to-sequence interface that enables standard decoder-only Transformers to learn molecular topology via Next Token Prediction.
-By unifying motif mining, causal serialization, and fine-to-coarse concatenation, CamS constructs a hierarchical context where global scaffolds are conditioned on dense local evidence.
-The resulting CamS-LLaMA achieves state-of-the-art performance on MoleculeNet and MoleculeACE, surpassing strong graph baselines while relying on minimal priors.
-Mechanism analysis confirms that this multi-scale causal design is crucial, as it explicitly drives attention toward cliff-determining structural differences.
-Our work validates that with the right representation interface, generic autoregressive foundation models can serve as powerful, structure-native engines for molecular science.
+We presented CamS, a graph-to-sequence interface that enables standard decoder-only Transformers to learn molecular topology via Next Token Prediction.
+Its FM prototype CamS-LLaMA achieve SOTA on MoleculeNet and MoleculeACE.
+Our multi-scale causal serialization is fundamental to this success, as mechanistic analysis confirms it explicitly drives attention toward cliff-driving structural differences.
+Ultimately, CamS validates generic AR FMs as powerful engines for structure-native molecular science given the right representation interface.
 Appendix A CamS-Tokenizer and Graph-to-Sequence Construction
 Overview.
-This section corresponds to the CamS-tokenizer description in Sec.
-3.1 (including Secs.
-3.1.1, 3.1.2, and 3.1.3) and the multi-view training-instance construction in Sec.
-3.2.1.
-It supplements the main text with implementation-level details of (1) GraphBPE merge learning and Single-Atom Vocabulary Closure (SAVC)(Algs.
-1 and Algs.
-2), (2) connection-aware motif vocabulary construction and encoding-time unknown recovery (Algs.3 and Algs.4), (3) deterministic scaffold-rooted BFS serialization (Algs.
-5), and (4) fine-to-coarse multi-scale concatenation together with NTP loss masking and length handling (Algs.
-6)
-A.1 Tokenizer Vocabulary Mining and Single-Atom Coverage (GraphBPE + SAVC)
+This section supplements the CamS-Tokenizer framework description in Sec.
+3.1 and the CamS-LLaMA pipeline in Sec.
+3.2.
+It provides implementation-level details for: (1) Vocabulary Mining: Learning merge operations and constructing the Single-Atom Vocabulary Closure (SAVC) (Algs.
+1 and 2); (2) Materialization: Building the connection-aware motif vocabulary (Alg.
+3); (3) Per-Scale Encoding: The recursive encoding process with unknown recovery (Alg.
+4) and the deterministic Scaffold-Rooted BFS serialization (Alg.
+5); (4) Cross-Scale Concatenation: Constructing multi-view training instances for NTP (Alg.
+6).
+A.1 Tokenizer Vocabulary Mining and Single-Atom Coverage
 Notation and Objects.
-We start from an RDKit molecule and construct the atom graph $G_{0}=(V_{0},E_{0})$, where each node is an atom and each edge is a chemical bond.
-A GraphBPE merge list is an ordered sequence of operations $\mathcal{O}=(o_{1},\dots,o_{K})$, where each operation $o_{t}$ is a canonical fragment code extracted from the union of two adjacent nodes (details below).
-Given a prefix $\mathcal{O}_{\leq k}$, applying merges on $G_{0}$ yields a motif graph $G=(V,E)$, where each motif node $v\in V$ stores the covered atom set $\mathrm{atom\_indices}(v)\subseteq V_{0}$.
+We start from an RDKit molecule and construct the atom graph $G_{0}=(V_{0},E_{0})$, where nodes are atoms and edges are chemical bonds.
+A merge list is an ordered sequence of operations $\mathcal{O}=(o_{1},\dots,o_{K})$, where each operation $o_{t}$ is a canonical fragment code extracted from the union of two adjacent nodes.
+The Motif Scale $s$ corresponds to a prefix length $k_{s}$, such that applying $\mathcal{O}_{\leq k_{s}}$ to $G_{0}$ yields a motif graph $G_{s}=(V_{s},E_{s})$, where each node $v\in V_{s}$ covers a subset of atoms $\mathrm{atom\_indices}(v)\subseteq V_{0}$.
 Connection-Aware Motif Representation.
 Each motif token is stored in a pair form $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$.
-The first records the fragment SMILES of the core atoms (no connectors), while the second inserts wildcard connector atoms * at all attachment sites together with explicit bond types, making the token identity connection-aware.
-We materialize these motif pairs by replaying $\mathcal{O}_{\leq k}$ on a tokenizer corpus and inserting connector * atoms for every cut bond (Alg.
+To ensure a deterministic and unique identifier for every substructure, we employ RDKit’s canonicalization routine (Landrum et al., 2021).
+The first component, $v_{\mathrm{noConn}}$, is the standard canonical SMILES of the fragment.
+The second component, $v_{\mathrm{withConn}}$, explicitly encodes connectivity information to differentiate chemically identical fragments with different attachment contexts (e.g., a pyridine ring attached at the 2-position vs.
+the 3-position).
+Specifically, at each attachment site (severed bond), we insert a dummy atom (wildcard *) preserving the original bond type.
+To guarantee a unique string representation invariant to atom ordering, we generate the canonical SMILES of this wildcard-augmented fragment.
+Thus, two motifs are identical if and only if they share the same graph topology, atom types, and attachment configurations.
+We materialize these pairs by replaying $\mathcal{O}_{\leq k_{s}}$ on the tokenizer corpus (Alg.
 3).
 Single-Atom Vocabulary Closure (SAVC).
-A key issue for property prediction is encoding coverage: if a rare but meaningful “single-atom + attachment-pattern” form is absent from the mined vocabulary, it would collapse to [UNK] and erase fine-grained distinctions.
-SAVC addresses this by (1) enumerating common connection-aware single-atom motifs for typical valence states, and (2) adding an element-wise fallback token X_AltForm to represent rare/atypical forms.
-At encoding time, if a queried motif contains exactly one core atom of element $X$ but the exact form is missing, we back off to X_AltForm instead of [UNK] (Alg.
+A key challenge for predictive encoding is coverage: if a rare “single-atom + attachment-pattern” form is absent from the mined vocabulary, standard tokenizers collapse it to [UNK], creating causal information breakpoints.
+SAVC addresses this by: (1) enumerating common connection-aware single-atom motifs for all typical valences, and (2) introducing an element-wise fallback token [X_AltForm] to represent rare/atypical forms.
+During encoding, if a queried motif contains exactly one core atom of element $X$ but the exact form is missing, we back off to [X_AltForm] instead of [UNK] (Alg.
 2).
 Encoding-Time Unknown Recovery.
-Encoding greedily applies $\mathcal{O}_{\leq k}$ and then maps each resulting motif to its token id.
-If a motif is still unknown after canonicalization, we recursively split it along its stored merge tree until known motifs are obtained; the recursion stops at single-atom leaves where SAVC backoff applies (Alg.
+Encoding greedily applies $\mathcal{O}_{\leq k_{s}}$ and maps resulting motifs to token IDs.
+If a motif is unknown (not in $\Sigma$), we recursively split it along its stored merge tree until known sub-motifs are obtained.
+The recursion guarantees termination at single-atom leaves, where SAVC back-off logic applies (Alg.
 4).
-Algorithm-to-component map.
-Alg.
-1 learns the ordered merge list $\mathcal{O}$ (Sec.
-3.1.1).
-Alg.
-3 materializes the connection-aware motif vocabulary by replaying $\mathcal{O}_{\leq k}$ and inserting connector atoms *.
-Alg.
-2 specifies SAVC enumeration and X_AltForm backoff.
-Alg.
-4 describes encoding-time motif extraction and recursive split for unknown recovery.
-A.2 Graph-to-Causal-Sequence Serialization (Traversal and Root Selection)
-Motif graph.
-After applying $\mathcal{O}_{\leq k}$ to $G_{0}$, we obtain a motif graph $G=(V,E)$ where each node $v\in V$ corresponds to a merged fragment with $\mathrm{atom\_indices}(v)\subseteq V_{0}$.
-An edge $(u,v)\in E$ indicates that at least one original bond in $G_{0}$ connects atoms across the two fragments.
-Scaffold-rooted BFS order.
-CamS serializes $G$ into a token sequence by a deterministic scaffold-rooted BFS: we select as root the motif with the largest number of covered atoms (ties broken deterministically, e.g., by smaller node id), and traverse the motif graph breadth-first with a fixed neighbor ordering (e.g., ascending node id).
-The visitation order defines the final token order, implemented by relabeling motif nodes so that ids $0,1,\dots$ encode the BFS order (Alg.
+A.2 Graph-to-Causal-Sequence Serialization
+Motif Graph Construction.
+After applying $\mathcal{O}_{\leq k_{s}}$ to $G_{0}$, we obtain the motif graph $G_{s}=(V_{s},E_{s})$.
+An edge $(u,v)\in E_{s}$ exists if any original bond in $G_{0}$ connects atoms across fragments $u$ and $v$.
+Scaffold-Rooted BFS Order (Intra-Scale Order).
+CamS serializes $G_{s}$ into a sequence via a deterministic Scaffold-Rooted BFS.
+To ensure that the serialization is invariant to the input atom permutation (i.e., canonical), we perform a standard molecule-level canonicalization (via RDKit) before any processing.
+This establishes a canonical ordering of the original atom indices $0,\dots,N-1$.
+During BFS on the motif graph $G_{s}$, we define the node_id of a motif $v$ as the minimum canonical atom index among its constituent atoms ($\min_{a\in\mathrm{atoms}(v)}\mathrm{idx}(a)$).
+We then select the root as the motif with the largest atom count (breaking ties by the smallest node_id) and traverse $G_{s}$ breadth-first.
+When expanding neighbors, we sort them by their node_id in ascending order.
+This strategy prioritizes global backbone structure before local substituents, establishing a stable Center-to-Periphery causal order (Alg.
 5).
-This yields a stable core-to-periphery causal order aligned with prefix-based NTP (Sec.
-3.1.2).
-Why no explicit edges in the sequence.
-We do not append edge/direction tokens: each motif token already includes explicit * attachment sites and bond types, so the attachment pattern is part of the token identity (Alg.
-3).
-The BFS order then provides the autoregressive causal structure at the sequence level.
-A.3 Multi-scale Concat Construction and Training-Instance Generation
-Prefix slicing and scale definition.
-GraphBPE yields a single merge list $\mathcal{O}=(o_{1},\dots,o_{K})$.
-Each vocabulary scale corresponds to a prefix $\mathcal{O}_{\leq k_{s}}$, which induces a sub-vocabulary $\Sigma_{\leq k_{s}}$ and a tokenizer $\iota_{s}$.
-This “train once, slice many” construction reuses the same learned merge statistics across all scales (Sec.
-3.1.3).
-Per-molecule views and fine-to-coarse concatenation.
-For a molecule $s$, we generate $S$ single-scale views $\{X_{s_{1}},\dots,X_{s_{S}}\}$ by encoding under tokenizers ordered from fine to coarse.
-We then build one concatenated view by inserting [CONCAT] as separators and keeping only one global [BOS] and [EOS].
-Alg.
-6 provides the pseudo-code used in implementation (Sec.
-3.2.1).
-Loss masking and length handling.
-For NTP training, we exclude [BOS], [EOS], and [CONCAT] positions from prediction targets (typically by setting labels to -100 at these indices).
-Sequences are padded with [PAD] up to the maximum length; if truncation is needed, we keep boundary tokens [BOS] and [EOS] and truncate the interior span consistently across views.
-Algorithm 1 GraphBPE merging-operation learning (ordered merge list $\mathcal{O}$)
-1:Input: tokenizer corpus $\mathcal{D}$ (RDKit molecules), max iterations $K$, min frequency $f_{\min}$
+A.3 Multi-Scale Concatenation and Training Views
+Scale Definition.
+We define a set of Motif Scales $S=\{s_{1},s_{2},\dots,s_{M}\}$ (e.g., $1\mathrm{K},7\mathrm{K},\dots$), each corresponding to a prefix $\mathcal{O}_{\leq s_{i}}$.
+This “train once, slice many” strategy allows reusing the same learned merge statistics across all scales.
+Data Augmentation via Views.
+For each molecule, we generate $M$ single-scale views $\{X^{(s_{1})},\dots,X^{(s_{M})}\}$ and one concatenated multi-scale view $\mathbf{X}$.
+The multi-scale view $\mathbf{X}$ is constructed by concatenating single-scale sequences in Fine-to-Coarse Order (Alg.
+6), enabling the model to leverage fine-grained details as context for coarse-grained predictions.
+NTP Loss Masking.
+For AR training, we compute loss on all motif tokens but exclude special tokens [BOS], [EOS], and [CONCAT] from prediction targets (typically by setting labels to -100).
+Algorithm 1 BPE-style Merge Operation Learning (List $\mathcal{O}$)
+1:Input: Tokenizer corpus $\mathcal{D}$, max iterations $K$, min frequency $f_{\min}$
 2:Output: Ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K^{\prime}})$
-3: Initialize each molecule $m\in\mathcal{D}$ with a working graph $G_{m}$ where every atom is a node and $\mathrm{atom\_indices}(i)=\{i\}$
-4: Compute initial statistics: for each $m$ and each edge $(u,v)\in E(G_{m})$, set $c\leftarrow\mathrm{FragSmiles}(m,\mathrm{atom\_indices}(u)\cup\mathrm{atom\_indices}(v))$ and increment $\mathrm{stats}[c]$
+3: Initialize graph $G_{m}$ for each molecule $m\in\mathcal{D}$ (every atom is a node)
+4: Compute initial pair statistics for all adjacent edges
 5:for $t=1$ to $K$ do
-6:  Select $o_{t}\leftarrow\arg\max_{c}\;(\mathrm{stats}[c],\,c)$ (tie-break by lexicographic $c$ as in code)
+6:  Select operation $o_{t}\leftarrow\arg\max_{c}\;(\mathrm{stats}[c],\,c)$
 7:  if $\mathrm{stats}[o_{t}]<f_{\min}$ then
@@
 10:  Append $o_{t}$ to $\mathcal{O}$
-11:  Apply $o_{t}$ to all molecules that contain it: in each affected $G_{m}$, merge every edge $(u,v)$ whose union-fragment code equals $o_{t}$
-12:  Update local pair statistics around merged nodes (remove old incident pairs; add new pairs)
-13:  Set $\mathrm{stats}[o_{t}]\leftarrow 0$ (avoid re-selecting the same operation)
-14:end for
-Algorithm 2 Single-Atom Vocabulary Closure (SAVC) with AltForm backoff (implementation-level)
-[…]
-2:Output: basic single-atom token list $\Sigma_{\mathrm{atom}}$
+11:  Apply $o_{t}$ to all applicable edges in $\mathcal{D}$, merging nodes into larger motifs
+12:  (Implementation Note: Merges are performed greedily. In case of overlaps, edges with smaller canonical atom indices are prioritized to ensure determinism.)
+13:  Update local pair statistics around merged nodes
+14:  Reset $\mathrm{stats}[o_{t}]\leftarrow 0$
+15:end for
+Algorithm 2 Single-Atom Vocabulary Closure (SAVC) & Back-off
+1:Input: Element set $\mathcal{E}$, valences $\mathrm{Val}(X)$, bond types $\mathcal{B}$
+2:Output: Basic single-atom vocabulary $\Sigma_{\mathrm{atom}}$
 3:for each element $X\in\mathcal{E}$ do
-4:  Add X (standalone single atom) and X_AltForm (rare/atypical fallback)
-5:  for each $v\in\mathrm{Val}(X)$ do
-6:   for each attachment degree $d\in\{1,\dots,v\}$ do
-7:    for each bond-type tuple $(b_{1},\dots,b_{d})\in\mathcal{B}^{d}$ do
-8:     Construct a single-atom fragment SMILES with $d$ wildcard substituents (use branching if $d>1$), e.g. *X(*)(*)* with explicit bonds
-9:     Add this motif_with_conn into $\Sigma_{\mathrm{atom}}$
+4:  Add [X] (standalone) and [X_AltForm] (fallback) to $\Sigma_{\mathrm{atom}}$
+5:  for each valence state and attachment pattern do
+6:   Construct connection-aware SMILES with wildcards (e.g., *X(*)*)
+7:   Add to $\Sigma_{\mathrm{atom}}$
+8:  end for
+9:end for
+10:Encoding-Time Back-off Logic:
+11:if queried motif is unknown AND contains exactly 1 core atom (element $X$) then
+12:  return ID of [X_AltForm]
+13:else
+14:  return [UNK]
+15:end if
+Algorithm 3 Connection-Aware Vocabulary Materialization
+1:Input: Corpus $\mathcal{D}$, operation prefix $\mathcal{O}_{\leq k}$, SAVC vocabulary $\Sigma_{\mathrm{atom}}$
+2:Output: Full vocabulary $\Sigma$ (Map: Canonical SMILES $\to$ ID)
+3:for each molecule $m\in\mathcal{D}$ do
+4:  Apply $\mathcal{O}_{\leq k}$ to partition atoms into motif nodes $V$
+5:  for each motif node $v\in V$ do
+6:   Extract core fragment SMILES $v_{\mathrm{noConn}}$
+7:   Extract connection-aware SMILES $v_{\mathrm{withConn}}$ (inserting * at cut bonds)
+8:   Add $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$ to candidate set
+9:  end for
+10:end for
+11: Filter candidates by frequency; Union with $\Sigma_{\mathrm{atom}}$ and Special Tokens to form $\Sigma$
+Algorithm 4 Encoding with Recursive Unknown Recovery
+1:Input: Molecule $M$, operations $\mathcal{O}_{\leq k}$, Vocabulary $\Sigma$
+2:Output: Ordered token list
+3:// Phase 1: Construction (Over-merge)
+4: Initialize atom graph $G$. Assign each node a leaf BPETreeNode.
+5:for operation $o\in\mathcal{O}_{\leq k}$ do
+6:  Find all edges matching $o$.
+7:  for edge $(u,v)$ in matches (ordered by node index) do
+8:   Merge $u,v$ into $w$.
+9:   Record merge history: $w.\text{tree}\leftarrow\text{Node}(\text{children}=[u.\text{tree},v.\text{tree}])$.
 10:  end for
 11:end for
-12:  end for
-13:end for
-14:Encoding-time backoff (MotifVocab.__getitem__): if a queried motif SMILES is not found after RDKit canonicalization, and it contains exactly one core atom (non-H, non-*), return the id of X_AltForm; otherwise return [UNK] (code uses id =1=-1 as unknown sentinel).
-Algorithm 3 Materializing the connection-aware motif vocabulary (motif mining + SAVC integration)
-1:Input: tokenizer corpus $\mathcal{D}$ (SMILES), operation prefix $\mathcal{O}_{\leq k}$, special tokens $\{\texttt{[PAD]},\texttt{[BOS]},\texttt{[EOS]},\texttt{[UNK]}\}$
-2:Output: motif-pair list $\mathcal{V}=\{(v_{\mathrm{noConn}},v_{\mathrm{withConn}})\}$ and token set $\Sigma$
-3:for each SMILES $s\in\mathcal{D}$ do
-4:  Parse $s$ to molecule $m$; build $G_{0}=(V_{0},E_{0})$ with atom/bond features
-5:  Apply merges in order $\mathcal{O}_{\leq k}$ on a copy of $G_{0}$ to obtain motif graph $G=(V,E)$ where each node $v$ stores $\mathrm{atom\_indices}(v)\subseteq V_{0}$
-6:  Assign each atom $i\in V_{0}$ to its motif node $\mathrm{bpe\_node}(i)\in V$
-7:  for each original bond $(i,j)\in E_{0}$ do
-8:   if $\mathrm{bpe\_node}(i)\neq\mathrm{bpe\_node}(j)$ then
-9:    Add two connector nodes $i^{\star},j^{\star}$ with smarts=*; attach $i^{\star}$ to $i$ and $j^{\star}$ to $j$
-10:    Set the bond types of $(i,i^{\star})$ and $(j,j^{\star})$ equal to the original bond type of $(i,j)$
-11:    Include $i^{\star}$ in $\mathrm{atom\_indices}(\mathrm{bpe\_node}(i))$ and $j^{\star}$ in $\mathrm{atom\_indices}(\mathrm{bpe\_node}(j))$
+12:// Phase 2: Serialization & Recovery
+13: Serialize nodes $V_{s}$ via Scaffold-Rooted BFS (Alg. 5).
+14: Initialize empty list $L$.
+15:for node $v$ in BFS order do
+16:  $L.\text{append}(\text{RecursiveResolve}(v,\Sigma))$
+17:end for
+18:return $L$
+19:Function RecursiveResolve(node $v$, Vocabulary $\Sigma$):
+20:if $v\in\Sigma$ then
+21:  return [ID($v$)]
+22:else
+23:  // Unknown motif: backtrack using the BPE tree built in Phase 1
+24:  Let $c_{1},c_{2}\leftarrow v.\text{tree}.\text{children}$
+25:  return RecursiveResolve($c_{1},\Sigma$) + RecursiveResolve($c_{2},\Sigma$)
+26:  // Base case: if single atom is unknown, apply SAVC back-off (Alg. 2)
+27:end if
+Algorithm 5 Scaffold-Rooted BFS (Intra-Scale Order)
+1:Input: Motif Graph $G_{s}=(V_{s},E_{s})$
+2:Output: Ordered list of nodes
+3: Select root $r\leftarrow\arg\max_{v\in V_{s}}(\text{atom\_count}(v),-\text{node\_id}(v))$
+4: Initialize Queue $Q\leftarrow[r]$, Visited $\leftarrow\{r\}$, Order $\leftarrow[]$
+5:while $Q$ is not empty do
+6:  $u\leftarrow Q.\text{pop\_front}()$
+7:  Append $u$ to Order
+8:  Get neighbors $\mathcal{N}(u)$, sort by node_id
+9:  for $v\in\mathcal{N}(u)$ do
+10:   if $v\notin\text{Visited}$ then
+11:    $Q.\text{push\_back}(v)$, Visited.add($v$)
 12:   end if
 13:  end for
-14:  for each motif node $v\in V$ do
-15:   $v_{\mathrm{noConn}}\leftarrow\mathrm{FragSmiles}(m,\mathrm{atom\_indices}(v)\cap V_{0})$
-16:   $v_{\mathrm{withConn}}\leftarrow\mathrm{GraphSmiles}(G_{0}[\mathrm{atom\_indices}(v)])$ (contains * connectors)
-17:   Add pair $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$ into a frequency counter
-18:  end for
-19:end for
-20: Prepend special tokens; then integrate the SAVC basic vocabulary (Alg. 2) into the pair list
-Algorithm 4 GraphBPE encoding with unknown-motif recovery (recursive split)
-1:Input: SMILES $s$, operations $\mathcal{O}_{\leq k}$, motif vocabulary $\Sigma$ (lookup by canonical SMILES)
-2:Output: ordered motif list $\big[(v_{\mathrm{noConn}},v_{\mathrm{withConn}})\big]$
-3:if $s$ contains “.then
-4:  Split $s$ into fragments and encode each fragment independently; concatenate motif lists
-5:end if
-6: Build initial atom graph $G_{0}$ and initialize a binary merge tree at each atom node
-7: Apply $\mathcal{O}_{\leq k}$ in order: for each code $o$, merge every edge whose union-fragment code equals $o$ (merge also updates the node’s merge tree)
-8: Relabel nodes to integers and reorder them by scaffold-rooted BFS (Alg. 5)
-9: Add connector * nodes for every cut bond to build connection-aware subgraphs (Alg. 3)
-10: Collect motif tokens in BFS order; if a motif is not in $\Sigma$, mark it as unknown
-11:while unknown motifs exist do
-12:  Pop one unknown motif node and recursively split it into its two children in the stored merge tree
-13:  For each child: if it is in $\Sigma$, emit it; else split further; stop at leaves (single-atom motifs), where SAVC backoff applies
 14:end while
-Algorithm 5 Scaffold-rooted BFS serialization (root selection + neighbor order)
-1:Input: motif graph $G=(V,E)$ with node attribute $\mathrm{atom\_indices}(v)$
-2:Output: a deterministic BFS ordering $\pi$ (used to relabel nodes)
-3: Select root $r\leftarrow\arg\max_{v\in V}\big(|\mathrm{atom\_indices}(v)|,\;-\mathrm{id}(v)\big)$ (tie-break: smaller id)
-4: Initialize queue with $r$, mark $r$ visited
-5:while queue not empty do
-6:  Pop front node $u$, append $u$ to ordering $\pi$
-7:  Let $\mathcal{N}(u)$ be neighbors of $u$ not yet visited, sorted by ascending node id
-8:  Push neighbors in $\mathcal{N}(u)$ into the queue and mark them visited
-9:end while
-10: Relabel nodes by $\pi$ so that the first visited node becomes id 0, the second becomes id 11, etc.
-Algorithm 6 Multi-scale CamS concatenation and training-view generation
+15:return Order
+Algorithm 6 Multi-Scale Concatenation (Inter-Scale Order)
+1:Input: Molecule $M$, scales $\{s_{1},\dots,s_{M}\}$ (Fine $\to$ Coarse)
+2:Output: Single-scale views $\{X^{(s_{j})}\}$, Multi-scale view $\mathbf{X}$
+3:for $j=1$ to $M$ do
+4:  $X^{(s_{j})}\leftarrow\text{Encode}(M,\mathcal{O}_{\leq s_{j}})$ (using Alg. 4)
+5:end for
+6:Construct Concatenated View:
 […]
-2:Output: single-scale views $\{X_{s_{j}}\}_{j=1}^{S}$ and concatenated view $X_{\mathrm{concat}}$
-3:for $j=1$ to $S$ do
-4:  $X_{s_{j}}\leftarrow\iota_{s_{j}}.\mathrm{encode}(s).\mathrm{ids}$ (already includes [BOS] and [EOS])
-5:end for
-6: Build concatenated ids:
-7:  $X_{\mathrm{concat}}\leftarrow[\texttt{bos\_id}]$
-8:for $j=1$ to $S$ do
-9:   Append the inner ids of $X_{s_{j}}$ (drop its BOS/EOS): $X_{s_{j}}[2{:}{-}1]$
-10:  if $j<S$ then
-11:    Append concat_id
+8:for $j=1$ to $M$ do
+9:  Append tokens of $X^{(s_{j})}$ (excluding its own BOS/EOS)
+10:  if $j<M$ then
+11:   Append [CONCAT]
 12:  end if
 13:end for
-14:  Append eos_id
-15:NTP loss masking: exclude [BOS], [EOS], [CONCAT] positions from prediction targets (implementation typically sets labels to -100 on these ids).
-16:Padding/truncation: pad with [PAD] to max length; if truncating, keep the boundary tokens [BOS] and [EOS].
+14: Append [EOS]
+15:return $\{X^{(s_{1})},\dots,X^{(s_{M})},\mathbf{X}\}$
 Appendix B Theoretical Derivations
@@
 This section corresponds to the theoretical analysis in Sec.
-3.3 and the supporting arguments referenced in Secs.
-3.1 and 3.2.1.
-It supplements the main text with derivation-level details of (1) the DPI-based proof for the information loss induced by stochastic corruption (Proposition 3.1; Sec.
-B.1), (2) the direct supervision density / target-coverage counting under NTP vs.
-masked modeling in the multi-scale setting (Sec.
-B.2), and (3) the explicit structural-bias formulation that contrasts graph-transformer static bias with CamS causal constraint plus learned aggregation (Eq.
-(11); Sec.
+3.3.
+It supplements the main text with derivation-level details of: (1) The DPI-based proof for the information loss induced by stochastic corruption, including a discussion on Graph-specific evidence uncertainty (Sec.
+B.1); (2) The quantitative analysis of Direct Supervision Density, contrasting NTP vs.
+MNP and quantifying the gain from multi-view augmentation (Sec.
+B.2); (3) The explicit structural-bias formulation that contrasts graph-transformer static bias with CamS causal constraint plus learned aggregation (Sec.
 B.3).
@@
 We formalize the conditioning-information gap between (1) uncorrupted evidence and (2) randomly masked evidence.
-Setup.
-For predicting token $x_{t}$, let $Z_{t}$ denote an unmasked evidence set (a collection of tokens) and let $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ be its stochastically masked version under a masking channel $\mathcal{M}$ (e.g., random node/substructure masking used in MLM/MNP).
-This yields a Markov chain
+Setup and Markov Chain.
+For predicting token $x_{t}$, let $Z_{t}$ denote an unmasked evidence set (e.g., the full causal history in CamS) and let $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ be its stochastically masked version under a masking channel $\mathcal{M}$ (e.g., random node/substructure masking used in MLM/MNP).
+This yields a Markov chain:
 $x_{t}\;\longleftrightarrow\;Z_{t}\;\longrightarrow\;\tilde{Z}_{t}.$ (8)
 Proof.
-[…]
-$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}),$ (15)
-which proves Proposition 3.1. ∎
-Remark (MLM limitations in NLP).
-XLNet points out two intrinsic issues of corruption-based MLM: (1) the special token corruption (e.g., [MASK]) creates a pretrain–finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context (Yang et al., 2019).
-Relatedly, ELECTRA shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions, while the [MASK] mismatch contributes a smaller but non-negligible part (Clark et al., 2020).
-Why this matters more on graphs (evidence-pattern uncertainty).
-On graphs, the most predictive evidence for a token often lies in its local neighborhood (e.g., bond context, ring connectivity, functional-group surroundings).
-Random masking can therefore remove not only the target token but also its critical neighborhood evidence, making the missingness pattern itself a stochastic nuisance.
-This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability: (1) masking structured units and adding global knowledge nodes (KPGT) (Li et al., 2023), (2) masking local subgraphs and predicting contextual properties (GROVER) (Rong et al., 2020), and (iii) structure-guided / curriculum masking rather than purely random masking (StructMAE) (Liu et al., 2024).
-These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (Wettig et al., 2023).
-CamS implication.
-For CamS, coarse-scale token prediction under NTP conditions on an uncorrupted fine-to-coarse history by construction.
-Thus, relative to MNP which observes $\tilde{Z}_{t}$, CamS provides a higher-SNR conditioning set $Z_{t}$ that avoids stochastic evidence deletion.
-This does not claim that an arbitrary prefix always dominates the full bidirectional context; rather, it claims that masking as corruption strictly reduces the information of whatever evidence set is available, and CamS is designed so that the evidence set for coarse tokens is both uncorrupted and structurally aligned with the fine-grained neighborhood evidence.
-B.2 Direct Supervision Density Analysis (Multi-Scale)
-As a heuristic proxy for the richness of the learning signal, we compare the density of direct token-level supervision signals and how they accumulate over optimization.
-Multi-scale notation.
-Let the multi-scale CamS sequence consist of $S$ scales with lengths $\{T_{s}\}_{s=1}^{S}$ and total tokens $T_{\text{all}}=\sum_{s=1}^{S}T_{s}$.
-Where gradients originate.
+By the Data Processing Inequality (DPI), for any Markov chain $X\!\leftrightarrow\!Z\!\rightarrow\!\tilde{Z}$, the data processing step $\tilde{Z}$ cannot increase the mutual information with the target $X$.
+Formally:
+$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}).$ (9)
+This proves Proposition 3.1. ∎
+Remark1: MLM limitations in NLP.
+Our analysis aligns with classic critiques of Masked Language Modeling (MLM).
+XLNet (Yang et al., 2019) points out two intrinsic issues of corruption-based objectives: (1) the special token corruption (e.g., [MASK]) creates a pretrain-finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context.
+Relatedly, ELECTRA (Clark et al., 2020) shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions.
+Remark 2: Graph-Specific Evidence Instability.
+While the DPI inequality holds generally, the loss is particularly severe on graphs due to Evidence-Pattern Uncertainty.
+The most predictive evidence for a token often lies in its immediate local neighborhood (e.g., bond context, ring connectivity, functional-group surroundings).
+Random masking $\mathcal{M}$ frequently removes not only the target token but also its critical neighborhood evidence (”co-masked”), making the conditional distribution $P(x_{t}|\tilde{Z}_{t})$ highly unstable across different masking patterns.
+This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability:
+  • KPGT (Li et al., 2023): Masks structured units and adds global knowledge nodes to anchor the context.
+  • GROVER (Rong et al., 2020): Masks local subgraphs and predicts contextual properties.
+  • StructMAE (Liu et al., 2024): Uses structure-guided / curriculum masking rather than purely random masking.
+These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (Wettig et al., 2023), but they still fundamentally operate on corrupted inputs.
+CamS Implication: In contrast, CamS conditions coarse-scale predictions on a structurally complete and uncorrupted fine-grained history.
+This ensures evidence stability and maximizes information gain without requiring auxiliary anchors.
+B.2 Direct Supervision Density Analysis (Decomposition)
+As a heuristic proxy for the richness of the learning signal, we compare the density of direct token-level supervision signals.
+To address the distinction between objective efficiency and data augmentation, we decompose the density gain into two factors: Intrinsic Efficiency and Systemic Multiplier.
+Gradient Origins.
 Let $\mathcal{L}$ be the token-level loss.
@@
   • NTP: Direct loss terms occur at (almost) all positions $t\in\{1,\dots,T_{\text{all}}-1\}$.
-Direct supervision density and accumulation.
+Factor 1: Intrinsic Objective Efficiency ($\times 1/\rho$).
 We define the Direct Supervision Density (SD) as the expected fraction of tokens that serve as prediction targets per update:
 $\text{SD}_{\text{NTP}}\approx 1,\qquad\text{SD}_{\text{MNP}}=\rho.$ (10)
-Equivalently, over $K$ optimization steps, the number of direct token-level targets scales as
-$\mathcal{R}_{\text{NTP}}\propto K\cdot T_{\text{all}},\qquad\mathcal{R}_{\text{MNP}}\propto K\cdot\rho\,T_{\text{all}}.$ (17)
-Thus, NTP provides about $1/\rho$ times more direct targets than MNP (e.g., $6.7\times$ when $\rho=0.15$).
+Thus, per pass, NTP provides about $1/\rho$ times more direct targets than MNP.
+For a typical masking rate $\rho=0.15$, this is a factor of $6.7\times$.
+This advantage is inherent to the NTP objective.
+Factor 2: Systemic Augmentation Multiplier ($\times M$).
+Crucially, the sequence representation of CamS naturally supports lightweight multi-view augmentation.
+We utilize 5 views per molecule (4 single-scale + 1 multi-scale) during pre-training.
+Let $\mathcal{R}$ be the total number of supervision targets provided by one molecule in one epoch:
+$\mathcal{R}_{\text{MNP}}$ $\approx 1\times(\rho\cdot T_{\text{avg}})$ (11)
+$\mathcal{R}_{\text{CamS}}$ $\approx 5\times(1\cdot T_{\text{avg}})$ (12)
+Even against aggressive masking ($\rho=0.5$ as in KPGT), CamS provides $10\times$ more signals per molecule.
+While MNP can ostensibly adopt augmentation, re-encoding massive graphs for every view is computationally costlier than re-slicing sequences.
+CamS leverages the product of these factors, resulting in substantially higher sample efficiency.
 Trade-off and Practical Masking Rates.
-Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off in masked modeling (Wettig et al., 2023).
-In molecular graph Transformers, KPGT reports that setting the masking rate to $\rho=0.5$ can be optimal when token design and global knowledge tokens help stabilize the conditioning context (Li et al., 2023).
-More broadly, structured masking / task design in graph pretraining (e.g., GROVER) can be viewed as attempts to obtain more useful target coverage without excessively destabilizing the evidence set (Rong et al., 2020).
-Why dense (and repeated) supervision can help.
-Dense token-level objectives are known to improve sample efficiency compared to losses defined only on a small masked subset; ELECTRA explicitly attributes a major part of MLM’s inefficiency to learning from only masked tokens rather than all positions (Clark et al., 2020).
-CamS strengthens this further via multi-view repetition: the same molecule contributes token-level targets repeatedly across scales, inducing cross-scale consistency pressure and providing redundant evidence pathways (a motif may appear as fine tokens at small scales and as a merged token at larger scales).
-This reduces sensitivity to any single tokenization granularity and makes the effective supervision more robust.
+Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off (Wettig et al., 2023).
+KPGT reports that setting $\rho=0.5$ is only viable when global knowledge tokens help stabilize the conditioning context (Li et al., 2023).
+CamS bypasses this dilemma entirely, achieving maximum density ($\text{SD}\approx 1$) with zero corruption.
 B.3 Structural Bias Formulation
-We distinguish the structural injection mechanism in Eq. (11).
-Graphformer (Hard Static Bias).
-The bias $\psi_{ij}$ is defined via static structural encodings (e.g., shortest-path distance), injected into attention as an additive bias term (Ying et al., 2021):
-$\psi_{ij}^{\text{Graph}}=b_{\phi(d_{G}(i,j))}.$ (18)
+We distinguish the structural injection mechanism in Eq. (6) based on the bias term $\psi_{ij}$.
+Graph Transformer (Hard Static Bias).
+Standard Graph Transformers inject structure via static encodings, such as Shortest Path Distance (SPD), added to the attention scores (Ying et al., 2021):
+$\psi_{ij}^{\text{Graph}}=b_{\text{SPD}(i,j)}.$ (13)
+This imposes an isotropic prior: all atoms at distance $k$ are treated equally by the structural bias, regardless of their semantic content.
 CamS (Hard Causal Constraint + Soft Learned Aggregation).
-CamS uses a hard causal mask to enforce fine-to-coarse information flow:
-$\psi_{ij}^{\text{CamS}}=\begin{cases}0&\text{if }(i\!\rightarrow\!j)\text{ is allowed by }\mathbf{M}_{\text{causal}}\\ -\infty&\text{otherwise.}\end{cases}$ (19)
-Within the permitted region, connectivity is content-adaptive via $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$; hence coarse-scale tokens learn which fine-scale neighborhood details to aggregate, rather than uniformly mixing neighbors at a given distance.
-Appendix C Training and Implementation Details
+CamS uses a hard causal mask to enforce fine-to-coarse information flow, as defined by the Inter-Scale Order:
+$\psi_{ij}^{\text{CamS}}=\begin{cases}0&\text{if }i\in\text{Prefix}(j)\\ -\infty&\text{otherwise.}\end{cases}$ (14)
+  • Bottom-up Composition: The Inter-Scale Order ensures that coarse motifs attend to their fine-grained constituents.
+  • Anisotropic Learning: Within the permitted prefix ($\psi_{ij}=0$), connectivity is content-adaptive via $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$.
+This allows the model to dynamically select relevant fine-scale neighborhood details (e.g., focusing on a specific pharmacophore) rather than uniformly mixing neighbors, yielding a learned, anisotropic aggregation.
+Appendix C Training Details and Benchmark Descriptions
 Overview.
-This section corresponds to the training and evaluation protocols in Secs.
-3.2.1 and 4.
-It supplements the main text with implementation-level details of (1) the shared training framework and parallelization setup for pre-training vs.
-fine-tuning (Sec.
-C.1), (2) the complete pre-training configuration on Enamine675M (MultiScaler3375M) (Table 5), (3) the fixed fine-tuning setup used across MoleculeNet and MoleculeACE (Table 6), and (4) the unified hyperparameter search space (including strategy-specific grids) (Table 7).
-C.1 Pre-training and Fine-tuning Hyperparameters
-Implementation and framework.
+This section supplements the experimental setup in Sec.
+4.
+It provides: (1) Implementation details for the shared training framework (Sec.
+C.1); (2) Complete hyperparameter configurations for pre-training (Table 5) and fine-tuning (Tables 6 and 7); (3) Detailed descriptions of the downstream tasks in MoleculeNet and MoleculeACE (Sec.
+C.2).
+C.1 Pre-training and Fine-tuning Implementation
+Framework.
 All experiments are implemented in PyTorch using HuggingFace Transformers.
-Pre-training uses the built-in Trainer with DeepSpeed ZeRO-2 on multi-GPU data parallelism; fine-tuning is run as single-GPU jobs (one job per GPU) using the same Trainer API (or strategy-specific subclasses).
-Pre-training on Enamine675M (MultiScaler3375M).
-We pre-train a 16-layer LLaMA-style decoder initialized from a local config (i.e., random initialization; no pretrained weights are loaded).
-The training data are pre-tokenized and stored as HuggingFace datasets on disk with train/validation splits.
-We use FP16 training with a cosine learning-rate schedule and warmup, and evaluate/save checkpoints periodically during training.
-Fine-tuning (MoleculeNet and MoleculeACE).
-We fine-tune the pretrained backbone with fingerprint injection (fp_dim=2048) and a lightweight supervised head.
-All downstream runs use FP16, evaluate every epoch, and select the best checkpoint by the primary validation metric (higher-is-better for AUROC; lower-is-better for RMSE).
-We also tested two dropout application schemes in the MoleculeACE fine-tuning pipeline: (1) head-only dropout (applied only in the supervised head), and (2) backbone-synced dropout (also overriding backbone attention_dropout/hidden_dropout and all nn.Dropout modules).
-Both schemes share the same dropout-rate grid in Table 7.
-Ablation compute budget.
-For ablations (e.g., single-scale tokenization or removing fingerprint injection), we keep the optimizer, LR schedule, batch size, and training length the same as the corresponding main setting, and only change the ablated component.
-Table 5: Pre-training hyperparameters on Enamine675M (MultiScaler3375M).
-Setting Value
-Backbone config 16 layers; hidden size 720; 8 heads; MLP 2880
-Context length 4096 (max_position_embeddings)
-Initialization new model from local config (no pretrained weights)
-Tokenizer CamS-tokenizer; embedding resized to tokenizer.vocab_size
-Training instances Multi-scale tokenized corpus (MultiScaler3375M; 5 views per molecule)
-Parallelism torchrun 8 processes + DeepSpeed ZeRO stage 2
-Precision FP16 (fp16=true in Trainer and DeepSpeed)
+Pre-training utilizes the Trainer API with DeepSpeed ZeRO-Stage 2 for multi-GPU data parallelism.
+Fine-tuning runs as single-GPU jobs using strategy-specific trainer subclasses.
+Pre-training Configuration.
+We train a 16-layer LLaMA-style decoder initialized from scratch (random initialization).
+The training corpus (Enamine675M augmented to $\sim$3.4B views) is pre-tokenized and stored as memory-mapped datasets.
+We use FP16 mixed precision, a cosine learning rate schedule with warmup, and periodic evaluation.
+The specific hyperparameters are listed in Table 5.
+Fine-tuning Protocol.
+We fine-tune the pre-trained backbone using the Dual-Path Strategy (Section 3.2), injecting fingerprints (specifically RDKit topological fingerprints, default parameters, 2048 bits) via a linear projection layer (fp_dim=2048).
+We employ two dropout schemes during fine-tuning (searched via grid search): (1) Head-only Dropout: Applied only to the task-specific prediction head.
+(2) Backbone-synced Dropout: Applied to the head and overriding all backbone dropout rates (attention and hidden states).
+Metrics are task-dependent: AUROC for classification (higher is better) and RMSE for regression (lower is better).
+Table 5: Pre-training hyperparameters on Enamine675M.
+Hyperparameter Value
+Backbone Architecture LLaMA Decoder (16 layers, 720 hidden size, 8 heads, 2880 MLP dim)
+Context Length 4096 tokens
+Initialization Random (from local config)
+Tokenizer CamS-Tokenizer (Vocab size $\approx$ 67K)
+Data Augmentation 5 views per molecule (4 single-scale + 1 multi-scale)
+Parallelism DeepSpeed ZeRO Stage 2 (8 GPUs)
+Precision FP16 Mixed Precision
 Micro-batch Size 256 per GPU
-Gradient accumulation 2
-Global batch size 4096 (256 $\times$ 8 $\times$ 2)
-Optimizer AdamW (Transformers Trainer default)
-Learning rate $1\times 10^{-4}$
+Gradient Accumulation 2 steps
+Global Batch Size 4096 sequences ($256\times 8\times 2$)
+Optimizer AdamW ($\beta_{1}=0.9,\beta_{2}=0.999$)
+Peak Learning Rate $1\times 10^{-4}$
 Weight Decay $1\times 10^{-2}$
-LR schedule cosine; warmup_steps =5000=5000
-Gradient clipping 1.0 (DeepSpeed gradient_clipping)
-Training length 10 epochs in planning (early stop after 1/10 epoch in actrual excution)
-Eval / save / log eval_steps =33000=33000; save_steps =33000=33000; logging_steps =3300=3300
-Checkpoint cap save_total_limit =100=100
-Table 6: Fixed fine-tuning setup for MoleculeNet and MoleculeACE.
+LR Schedule Cosine with 5000 warmup steps
+Gradient Clipping 1.0
+Training Duration Two-stage training: Stage 1 (1 epoch) + Stage 2 (0.1 epoch with adjusted LR). Total samples seen $\approx$ 3.7B (where 1 epoch = 3.4B augmented views).
+Table 6: Fine-tuning setup for downstream tasks.
 Setting MoleculeNet MoleculeACE
-Task type classification / regression (per dataset) regression
-Data split provided scaffold splits (same as KPGT), fold $\in\{0,1,2\}$ provided scaffold splits (same as KPGT)
-Epochs 80 (search phase) 50
-Batch size 32 per GPU (default) 32 per GPU (default; a small subset rerun with 64)
-Grad accumulation 1 (default) 1(default)
-Optimizer AdamW (Trainer default) AdamW (Trainer default)
-LR schedule constant_with_warmup, warmup_ratio =0.1=0.1 constant_with_warmup, warmup_ratio =0.1=0.1
-Evaluation per epoch per epoch
-Primary metric AUROC / macro-AUROC; RMSE for regression tasks RMSE (overall_RMSE)
-Table 7: Fine-tuning hyperparameter search space (shared).
-Hyperparameter Candidates
-Base grid (48 combos) learning_rate $\in\{10^{-6},3\!\times\!10^{-6},10^{-5},3\!\times\!10^{-5}\}$
-dropout_rate $\in\{0,0.05,0.1,0.2\}$; weight_decay $\in\{0,10^{-6},10^{-4}\}$
-Fine-tune strategy normal, l2sp, llrd, flag, reinit
-LLRD layer_decay $\in\{0.95,0.9,0.85,0.8\}$
-L2SP weight $\in\{10^{-1},10^{-2},10^{-3},10^{-4}\}$
-FLAG $\epsilon\in\{10^{-3},3\!\times\!10^{-3},5\!\times\!10^{-3},10^{-2}\}$
-Reinit reinit_num_layers $\in\{1,2,3,4\}$
-Appendix D Detailed Experimental Results
+Task Type Classification / Regression Regression
+Split Strategy Scaffold Split (KPGT protocol) Scaffold Split (Benchmark default)
+Epochs 80 50
+Batch Size 32 (per GPU) 32 (per GPU)
+Optimizer AdamW AdamW
+LR Schedule Linear decay (warmup ratio 0.1) Linear decay (warmup ratio 0.1)
+Metric ROC-AUC (Cls) / RMSE (Reg) RMSE
+Table 7: Hyperparameter search space for fine-tuning.
+Hyperparameter Search Grid
+Learning Rate $\{1\!\times\!10^{-6},3\!\times\!10^{-6},1\!\times\!10^{-5},3\!\times\!10^{-5}\}$
+Dropout Rate $\{0.0,0.05,0.1,0.2\}$
+Weight Decay $\{0.0,1\!\times\!10^{-6},1\!\times\!10^{-4}\}$
+Fine-tune Strategy Standard, L2SP, LLRD, FLAG, Reinit
+C.2 Benchmark Task Descriptions
+We evaluate CamS-LLaMA on two complementary benchmarks to assess both general property prediction and structural sensitivity.
+MoleculeNet (General Properties) (Wu et al., 2018).
+This benchmark covers a diverse range of molecular properties. We select 11 datasets that cover physiology, biophysics, and physical chemistry. Consistent with our results table, the tasks are:
+  • Classification Tasks (8):
+    • BACE: Inhibition of $\beta$-secretase (a key Alzheimer’s therapeutic target).
+    • BBBP: Blood-brain barrier penetration (permeability).
+    • ClinTox: Clinical toxicity (distinguishing FDA-approved drugs from toxic compounds).
+    • Estrogen: Estrogen receptor ($\alpha$, $\beta$) binding activity (endocrine disruption potential).
+    • Metstab: Metabolic stability (half-life duration in liver microsomes).
+    • SIDER: Adverse drug reactions (side effects) of marketed medicines.
+    • ToxCast: High-throughput toxicology screening data.
+    • Tox21: Toxicity testing across 12 biological targets (nuclear receptors/stress pathways).
+  • Regression Tasks (3):
+    • ESOL: Water solubility (log solubility in mols per litre).
+    • FreeSolv: Hydration free energy (experimental vs calculated).
+    • Lipo: Lipophilicity (octanol/water distribution coefficient, logD).
+All datasets follow the scaffold splitting strategy (as used in KPGT) to strictly evaluate the model’s generalization capability to chemically distinct structures.
+MoleculeACE (Activity Cliffs) (Van Tilborg et al., 2022).
+MoleculeACE is designed to stress-test models on Activity Cliffs—pairs of molecules with high structural similarity but large differences in potency.
+It consists of 30 datasets derived from ChEMBL, each targeting a specific biological protein target (e.g., CHEMBL204, CHEMBL240).
+  • Challenge: Unlike MoleculeNet, which often rewards global scaffold recognition, MoleculeACE requires the model to identify fine-grained structural edits (e.g., methylation, halogenation) that trigger drastic activity shifts.
+  • Metric: Performance is measured by RMSE on the test set (scaffold split). Lower RMSE indicates better ability to capture the non-smooth structure-activity landscape.
+Evaluation Protocols and Statistical Reporting.
+[…]
+Appendix D Comparison with Large-scale SMILES FMs
+[…]
+Table 8: Comparison with a SOTA SMILES-based Natural Language Science Foundation Models (NatureLM (Xia et al., 2025)).
+Results for NatureLM are taken directly from the original paper (Table 19).
+CamS-LLaMA ($\sim$0.1B) significantly outperforms NatureLM (up to 56B parameters) on common benchmarks, demonstrating superior representational efficiency.
+Model Params BBBP BACE Tox21
+NatureLM 1B 0.711 0.794 0.683
+NatureLM 8B 0.702 0.820 0.698
+[…]
+CamS-LLaMA 0.1B 0.942 0.870 0.827
+Appendix E Extended Discussion on Data Scale, Fairness, and Efficiency
+A prominent distinction between CamS-LLaMA and graph-native baselines (e.g., KPGT) is the pre-training corpus size (675M vs.
+550K).
+We emphasize that this disparity does not constitute experimental unfairness, but rather demonstrates a critical methodological advantage of our sequence-based framework: Scalability.
+Graph-native foundation models typically rely on heavy auxiliary inputs or targets to compensate for the lack of semantic density in raw graphs. This creates distinct scaling barriers:
+  • Descriptor Calculation (e.g., KPGT): KPGT requires computing over 200 RDKit descriptors and fingerprints for every training instance to serve as ”knowledge” targets.
+Scaling this dense annotation to the 675M-scale Enamine dataset (let alone augmented views) imposes significant data engineering and storage overheads.
+  • 3D Conformer Generation (e.g., GEM, Uni-Mol): Other baselines like GEM explicitly rely on 3D geometric views.
+Generating conformers (via ETKDG or DFT) allows for richer physics but is computationally prohibitive at the billion-scale.
+These pre-processing costs limit such methods to smaller datasets (e.g., $\sim$500K) by design necessity.
+In contrast, CamS transforms molecular graphs into causal sequences via purely logical operations (BPE merge and BFS traversal).
+These operations involve only string manipulation and standard graph traversal, which are CPU-efficient and easily parallelizable.
+CamS operates directly on pure molecular topology without requiring external descriptor injection or 3D optimization during pre-training.
+This efficiency removes the data-scaling bottleneck, allowing our method to leverage billion-scale augmented datasets as a standard feature of its training pipeline.
+Diversity over Repetition (High-Coverage Training).
+Our model was trained for $\sim$1 epoch on the augmented Enamine dataset (5 views per molecule).
+While baselines typically train on a small dataset (550K) for many epochs (High Repetition), CamS processes a vast number of unique structures (675M) with limited repetition per structure (High Diversity).
+The performance superiority of CamS, therefore, stems from its ability to efficiently access and learn from the breadth of the chemical space.
+This ”Diversity over Repetition” regime is unlocked specifically by the efficient nature of the CamS representation, which would be computationally infeasible for descriptor-heavy graph baselines.
+Appendix F Detailed Experimental Results
 Overview.
@@
 GraphCL 0.760 24 GAT 1.049 48
-Appendix E Interpretability Details for Activity-Cliff Attention Analysis
+Appendix G Interpretability Details for Activity-Cliff Attention Analysis
 Overview.
-This section corresponds to the activity-cliff attention analysis in Sec.
-4.3.
-It supplements the main text with implementation-level details of (1) activity-cliff pair construction and differential/shared atom identification (Appendix E.1), Alg.7, (2) mapping atom-level diff/shared labels to CamS tokens within each scale region (Appendix E.2), and (3) attention extraction and metric computation/aggregation for MDTA/MSTA and Rel-DTAP (Appendix E.3).
-E.1 Activity-cliff Pair Construction and Differential/Shared Atom Identification
+This section provides implementation details for the attention analysis presented in Section 4.3.
+It complements the main text by specifying: (1) the exact algorithm for identifying differential atoms in activity-cliff pairs (Appendix G.1, Algorithm 7); (2) the mapping from atom-level labels to CamS tokens across different Motif Scales (Appendix G.2); and (3) the detailed computation of the Rel-DTAP metric (Appendix G.3).
+G.1 Activity-cliff Pair Construction and Differential/Shared Atom Identification
 Pair construction (MoleculeACE-style).
@@
 This mapping is not required for computing Rel-DTAP, but enables atom-level highlighting across the two molecules.
-E.2 Mapping Atom-level Diff/Shared Labels to Tokens in Each Scale Region
+G.2 Mapping Atom-level Diff/Shared Labels to Tokens in Each Scale Region
 Token-to-atom alignment.
@@
 From atom-level labels to token-level labels.
-Let $\mathcal{A}_{\Delta}$ be the differential-atom set for this molecule from Appendix E.1.
+Let $\mathcal{A}_{\Delta}$ be the differential-atom set for this molecule from Appendix G.1.
 We assign a token-level differential indicator by an any-diff rule: a token at position $t$ is marked as differential iff $S_{t}\cap\mathcal{A}_{\Delta}\neq\varnothing$; otherwise it is marked as shared/non-differential.
@@
 This yields the four scale regions (e.g., 1K/7K/27K/67K in the paper) used to report scale-wise attention preferences.
-E.3 Attention Extraction and Metric Computation
+G.3 Attention Extraction and Metric Computation
 Attention extraction protocol.
@@
 MDTA/MSTA within a scale region.
-Within a given scale region $s$ (defined by [CONCAT] boundaries), let $p_{i}$ denote the (renormalized) attention weight on token $i$ in that region, and let $d_{i}\in\{0,1\}$ be its differential indicator from Appendix E.2.
+Within a given scale region $s$ (defined by [CONCAT] boundaries), let $p_{i}$ denote the (renormalized) attention weight on token $i$ in that region, and let $d_{i}\in\{0,1\}$ be its differential indicator from Appendix G.2.
 We define the mean differential-token attention and mean shared-token attention as
@@
 29:end for
+G.4 Case-Study Pair Selection Protocol
+To ensure the representativeness and reproducibility of the qualitative analysis in Section 4.3, we employ a rigorous, log-driven selection pipeline rather than manual cherry-picking.
+We build case studies from the test-set activity-cliff pairs identified in Appendix G.1.
+We implement three complementary selection modes:
+  1. 1.
+    Mode A: “Similar-structure” cases (Targeting Minimal Edits). This mode isolates pairs that differ by only a few atoms but yield a large potency change.
+    • Atom-count prefilter: We require the atom count difference $\Delta N\leq 3$ and pair size $N_{\max}\leq 100$.
+    • MCS-based scoring: We compute the MCS and define the edit size $d_{\max}$ as the number of non-MCS atoms. We prioritize pairs with $d_{\max}\leq 5$.
+    • Ranking: Pairs are ranked by increasing edit size (more similar first) and then by decreasing fold change.
+  2. 2.
+    Mode B: “Largest fold-change” cases (Targeting Extreme Cliffs).
+This mode targets the most extreme potency shifts regardless of edit size.
+We sort eligible pairs by potency fold-change in descending order and select top candidates.
+  3. 3.
+    Mode C: “Relatively larger molecules” cases (Targeting Complexity).
+This mode ensures coverage of complex molecules.
+We filter for pairs with $N_{\max}\geq 40$ and small $\Delta N$, selecting those with the largest molecule sizes.
+Justification of Token-Level Labeling (Holistic Chemical Semantics).
+A potential concern is whether the ”any-diff” rule (marking a coarse motif as differential if it contains any modified atom) biases metrics toward larger tokens.
+We argue this design is grounded in holistic chemical semantics.
+In medicinal chemistry, modifying a single atom within a ring system or functional group (e.g., H $\to$ F on a benzene ring) fundamentally alters the electronic and steric properties of the entire substructure.
+In CamS, such a modification results in a completely distinct Token ID for the coarse motif.
+Therefore, high attention to this ”diff-containing” coarse token reflects a valid recognition of the macro-semantic shift of the functional group as a whole, rather than a metric artifact.
+G.5 Note on Baselines.
+The Rel-DTAP metric is specifically designed to evaluate the multi-granular attention allocation inherent to the CamS hierarchy (mapping attention to specific $1\mathrm{K}$ vs.
+$67\mathrm{K}$ token regions).
+Since standard graph baselines (e.g., KPGT, GROVER) operate on fixed input granularities (atoms or triplets) and lack this explicit hierarchical tokenization, this metric is not directly applicable to them.
+Thus, our analysis focuses on the intrinsic mechanism of CamS.
+G.6 Additional Case Studies
+We provide additional visualizations for pairs selected via Mode A (Minimal Edits) and Mode B (Extreme Cliffs) to further substantiate the findings.
+Figure 3: Additional Case Study 1 (Minimal Edit). Attention heatmap for pairs (from CHEMBL234Ki) with minimal-atom substitution causing activity cliffs.
+Figure 4: Additional Case Study 2 (Extreme Cliff). Attention heatmap for pairs with a scaffold-hopping modification yielding a $4000\times$ shift.
+Appendix H Detailed Ablation Analysis
+In this section, we provide the complete breakdown of the ablation study.
+Budget Matching Protocol.
+To ensure a fair comparison, all single-scale variants were trained with a computational budget strictly matched to the full model.
+Specifically, we controlled the total number of parameter update steps (and thus the total number of training samples seen) to be identical across all settings, eliminating training duration as a confounding factor.
+We compare the following variants across all tasks:
+  • CamS (Full): The proposed multi-scale model with fingerprint injection.
+  • w/o FP: The pure multi-scale sequence model without fingerprint injection.
+  • 1K Only: Single-scale model using fine-grained motifs.
+  • 67K Only: Single-scale model using coarse-grained motifs.
+1.
+Impact of Multi-Scale Concatenation.
+As shown in Table 11, the CamS (Full) model consistently outperforms the single-scale variants (1K Only and 67K Only).
+Notably, the 67K Only variant suffers significant performance degradation in regression tasks (AVG RMSE 1.329), confirming that excessive compression reduces supervision density.
+2.
+Pure Sequence vs.
+Fingerprint Injection.
+On tasks such as Estrogen and ESOL, the w/o FP variant achieves the best performance (bolded).
+This indicates that the intrinsic structural features learned by CamS are sometimes superior to explicit fingerprints.
+However, as shown in Table 12, FP injection is crucial for stability on activity-cliff tasks.
+[…]
+Task KPGT CamS (Full) w/o FP 1K Only 67K Only
+Classification (AUROC $\uparrow$)
+BACE 0.855(0.014) 0.870(0.013) 0.850(0.014) 0.845(0.030) 0.837(0.024)
+BBBP 0.908(0.012) 0.942(0.015) 0.933(0.021) 0.926(0.016) 0.873(0.013)
+ClinTox 0.946(0.026) 0.935(0.017) 0.918(0.004) 0.921(0.021) 0.885(0.023)
+Estrogen 0.906(0.034) 0.917(0.050) 0.920(0.046) 0.912(0.056) 0.904(0.069)
+Metstab 0.889(0.057) 0.891(0.059) 0.888(0.063) 0.875(0.047) 0.869(0.072)
+SIDER 0.649(0.011) 0.655(0.016) 0.647(0.023) 0.646(0.017) 0.654(0.015)
+ToxCast 0.745(0.003) 0.724(0.008) 0.723(0.019) 0.717(0.016) 0.706(0.015)
+Tox21 0.848(0.017) 0.827(0.028) 0.823(0.032) 0.821(0.024) 0.815(0.016)
+AVG (Cls) 0.843 0.845 0.838 0.833 0.818
+Regression (RMSE $\downarrow$)
+ESOL 0.804(0.102) 0.761(0.046) 0.740(0.054) 0.803(0.045) 0.956(0.045)
+FreeSolv 2.121(1.025) 2.110(0.959) 2.192(0.832) 2.188(0.820) 2.328(0.919)
+Lipo 0.600(0.012) 0.645(0.023) 0.652(0.022) 0.653(0.027) 0.706(0.014)
+AVG (Reg) 1.175 1.172 1.195 1.215 1.329
+Table 12: Full Ablation Results on MoleculeACE (RMSE $\downarrow$). The column order is consistent with Table 11. The Full model consistently yields the best stability (lowest RMSE).
+Task KPGT CamS (Full) w/o FP 1K Only 67K Only
+CHEMBL1862 0.633 0.600 0.635 0.624 0.609
+CHEMBL1871 0.605 0.604 0.604 0.606 0.623
+CHEMBL2034 0.679 0.619 0.679 0.615 0.655
+CHEMBL204 0.666 0.709 0.749 0.730 0.726
+CHEMBL2047 0.578 0.519 0.578 0.518 0.563
+CHEMBL214 0.652 0.635 0.662 0.660 0.663
+CHEMBL2147 0.587 0.577 0.632 0.605 0.596
+CHEMBL218 0.625 0.632 0.687 0.655 0.646
+CHEMBL219 0.718 0.729 0.736 0.723 0.738
+CHEMBL228 0.669 0.669 0.679 0.676 0.713
+CHEMBL231 0.610 0.630 0.642 0.642 0.638
+CHEMBL233 0.691 0.692 0.719 0.724 0.712
+CHEMBL234 0.606 0.624 0.630 0.643 0.645
+CHEMBL235 0.624 0.612 0.629 0.604 0.629
+CHEMBL236 0.655 0.669 0.709 0.708 0.704
+CHEMBL237E 0.716 0.684 0.695 0.712 0.737
+CHEMBL237K 0.660 0.659 0.659 0.721 0.712
+CHEMBL238 0.537 0.537 0.585 0.564 0.572
+CHEMBL239 0.644 0.647 0.672 0.676 0.655
+CHEMBL244 0.698 0.696 0.726 0.723 0.725
+CHEMBL262 0.627 0.629 0.662 0.637 0.662
+CHEMBL264 0.574 0.562 0.570 0.583 0.583
+CHEMBL2835 0.373 0.384 0.392 0.385 0.384
+CHEMBL287 0.706 0.685 0.685 0.683 0.742
+CHEMBL2971 0.571 0.574 0.596 0.584 0.584
+CHEMBL3979 0.669 0.639 0.672 0.652 0.655
+CHEMBL4005 0.559 0.543 0.581 0.553 0.561
+CHEMBL4203 0.820 0.787 0.820 0.800 0.811
+CHEMBL4616 0.587 0.538 0.565 0.553 0.565
+CHEMBL4792 0.619 0.651 0.659 0.668 0.658
+AVG 0.632 0.624 0.650 0.641 0.649